### v1

# DPBridge: Latent Diffusion Bridge for Dense Prediction

Abstract
Diffusion models have demonstrated remarkable success in dense prediction problems, which aims to model per-pixel relationship between RGB images and dense signal maps, thanks to their ability to effectively capture complex data distributions. However, initiating the reverse sampling trajectory from uninformative noise prior introduces limitations such as degraded performance and slow inference speed. In this work, we propose DPBridge, a generative framework that formulates dense prediction tasks as image-conditioned generation problems and establishes a direct mapping between input image and its corresponding dense map based on fully-tractable diffusion bridge process. This approach addresses aforementioned limitations in conventional diffusion-based solutions. In addition, we introduce finetuning strategies to adapt our model from pretrained image diffusion backbone, leveraging its rich visual prior knowledge to facilitate both efficient training and robust generalization ability. Experimental results shows that our DPBridge can achieve competitive performance compared to both feed-forward and diffusion-based approaches across various benchmarks, highlighting its effectiveness and adaptability.
Figure 1: We present DPBridge to tackle general dense prediction tasks by establishing a stochastic transport between an RGB image and its corresponding dense map in the latent space based on a tractable diffusion bridge process. In addition, we introduce finetuning strategies to adapt large pretrained diffusion backbone to construct our bridge, leveraging its rich prior knowledge to enable both efficient training and robust generalization.
1 Introduction
Diffusion models recently emerge as a family of powerful generative models that are able to capture complex data distributions through a sequence of forward and reverse processes [26, 46]. They accelerate the development of modern generative systems, and have consistently obtained state-of-the-art performance across various domains, such as image synthesis and video generation [17, 46, 27], controlled editing [61, 5, 50], human pose estimation [29, 30, 28], inverse problem solving [11, 53, 40].
Despite such success, diffusion models still face limitations that hinder broader applications, especially in discriminative tasks like dense prediction. One major limitation stems from the fact that standard diffusion reverse process always starts from Gaussian noise, which contains little information about the target distribution. This will lead to degraded performance or prolonged convergence times. Current diffusion-based approaches formulate this type of tasks as image-conditioned generation problems and adopt measures like incorporating external conditions to the model [47, 61, 25], modifying the generation pipeline [41, 54], or adding extra guidance terms during sampling [11, 53] to deal with this issue. In this way they hope to better model the conditional generation process.
Instead of starting from noise, an feasible alternative is to leverage the inherent problem structure. Since the goal of dense prediction tasks is to model the probability transport between image and dense signal map distributions, it’s a rather straightforward idea to directly start the generative process from input RGB images, as they are structurally correlated to target signal maps and can naturally serve as more informative prior compared to noise. This motivates the exploration of a new family of computational frameworks, i.e. diffusion bridge models [39, 63, 36, 37]. They are based on a special type of diffusion process conditioned on given endpoints, thus enabling optimal transport between two distributions and generalizing beyond Gaussian priors.
Inspired by these intuition, we propose Dense Prediction Bridge (DPBridge), a latent diffusion bridge model adapted from pretrained image diffusion backbone to handle general dense prediction tasks though the image-conditioned generation pipeline. In essence, we construct diffusion bridge process between the latent distributions of images and dense signal maps, with all intermediate samples along the trajectory being convex combinations of the paired data. The generative process thus becomes a series of transformations that iteratively synthesize desired signal maps starting from input images rather than pure noise, enabling faster inference speed as well as better generation quality.
In addition, to better adapt pretrained diffusion backbone to our DPBridge construction, we also propose novel finetuning strategies to allow for more effective knowledge transfer. In this way our model is augmented with rich visual prior knowledge, which not only helps boosting the perception performance as well as generalization capability, but also significantly speeds up training convergence. Fig.(1) is a high-level overview of DPBridge, while also demonstrating its effectiveness on three representative dense prediction tasks, i.e. depth estimation, surface normal prediction, and semantic segmentation, proving that our framework is theoretically universal and can be instantiated to different scenarios.
To summarize, our contributions are as follows:
  • We propose DPBridge framework to deal with dense prediction tasks with generative pipeline by formulating them as instances of latent diffusion bridges.
  • We propose finetuning strategies to effectively adapt large pretrained backbone to diffusion bridges, allowing for exploitation of visual prior knowledge and training efficiency improvements.
  • Our proposed method exhibits competitive performance across different tasks in terms of both quantitative metrics and visual inspection, highlighting its effectiveness and generalization capability.
2 Related Works
2.1 Dense Prediction
Dense prediction refers to the type of problems whose outputs match the spatial dimensions of input data, often with the goal of assigning a value to each individual pixel. Typical examples include depth estimation, surface normal prediction, semantic segmentation, optical flow, etc. Common solutions can be categorized in discriminative and generative manner. Discriminative approaches focus on architecture design and introduce techniques like skip connections, dilated convolutions, multi-scale feature aggregation, and attention mechanisms [9, 44, 34] to better capture both local and global context for improved prediction accuracy. Generative methods, on the other hand, offer an alternative paradigm to tackle these problems through distribution modeling. Some works reuse or merge prior features from pretrained vision foundation models [4, 23, 62]. Others formulate these tasks as visual-condition guided generation process which refines the desired dense map iteratively [18, 32, 24].
2.2 Diffusion Bridge Models
Stochastic bridge models are a common tool in probability theory which describe the evolution of stochastic processes with given starting/ending points. When combined with diffusion models, they provide a new perspective to translate between different distributions and model conditional probabilities without the need for prior knowledge, therefore have been gaining increasing attention in the context of generative modeling in recent year [39, 38, 16]. They demonstrate the advantages of data-to-data generation paradigm over noise-to-data one on various tasks, including image-to-image translation [63, 37, 36], image restoration [59, 13], text-to-speech synthesis [10], etc. However, to the best of our knowledge, no prior work has explored leveraging diffusion bridge models on dense prediction tasks to facilitate the distribution transport between images and dense signal maps, despite their intuitive structural closeness. Our work aims to fill in this research gap.
3 Preliminary
3.1 Diffusion Process
A diffusion process =$\mathbb{Q}=\{\mathbf{z}_{t}\in\mathbb{R}^{d}:t\in[0,T]\}$ on space $\mathbb{R}^{d}$ follows the stochastic differential equation of form
$\mathrm{d}\mathbf{z}_{t}=f(\mathbf{z}_{t},t)\mathrm{d}t+g(t)\mathrm{d}\mathbf{w}_{t}$ (1)
where $\mathbf{w}_{t}$ is a standard Wiener process, $f:\mathbb{R}^{d}\times[0,T]\rightarrow\mathbb{R}$ and $g:[0,T]\rightarrow\mathbb{R}$ are the drift and diffusion coefficients that describe the evolution of this stochastic process. The transition kernel is specified as $q(\mathbf{z}_{t}|\mathbf{z}_{0})=\mathcal{N}(\alpha_{t}\mathbf{z}_{0},\sigma_{t}^{2}\mathbf{I})$ and the noise schedule comprising differentiable functions $\alpha_{t},\sigma_{t}$ depends on the design choice of drift and diffusion coefficients. It is also common to assess the diffusion process in terms of the signal-to-noise ratio (SNR), defined as $\text{SNR}_{t}=\alpha_{t}^{2}/\sigma_{t}^{2}$
3.2 Diffusion Bridge Process
Diffusion bridge process is a special type of diffusion process that guarantee to arrive at a prescribed final state $T$. One natural way of its construction is to derive the conditioned process from a reference diffusion process given terminal constraints. We can use Doob’s h-transform technique [42, 48] to modify a reference process $\mathbb{Q}$ defined in Eqn.(1) to ensure that the event $\mathbf{z}_{T}=\mathbf{x}$ happens. Therefore the conditioned process $\mathbb{Q}^{\mathbf{x}}(\cdot):=\mathbb{Q}\left(\cdot|\mathbf{z}_{T}=\mathbf{x}\right)$ follows the law of
$\mathrm{d}\mathbf{z}_{t}=\left[f(\mathbf{z}_{t},t)+g^{2}(t)h(\mathbf{z}_{t},t,\mathbf{x},T)\right]\mathrm{d}t+g(t)\mathrm{d}\mathbf{w}_{t}$ (2)
where $h(\mathbf{z}_{t},t,\mathbf{x},T)=\nabla_{\mathbf{z}_{t}}\log q(\mathbf{z}_{T}=\mathbf{x}|\mathbf{z}_{t})$. Intuitively, the additional drift term $h(\mathbf{z}_{t},t,\mathbf{x},T)$ plays the role of steering $\mathbf{z}_{t}$ towards the specified target point $\mathbf{z}_{T}=\mathbf{x}$, and it is ensured that $\mathbb{Q}^{\mathbf{x}}$ invariably passes $\mathbf{x}$ at timestep $T$ with probability 1 for any initial state $\mathbf{z}_{0}$.
The transition kernel of $\mathbb{Q}^{\mathbf{x}}$, given by $q(\mathbf{z}_{t}|\mathbf{z}_{0},\mathbf{z}_{T})$, can be subsequently derived through Kolmogorov equation:
$q(\mathbf{z}_{t}|\mathbf{z}_{0},\mathbf{z}_{T})$ $=\mathcal{N}(m_{t}\mathbf{z}_{0}+n_{t}\mathbf{z}_{T},\sigma_{t}^{2}\mathbf{I})$ (3)
$m_{t}$ $=\alpha_{t}\left(1-\frac{\text{SNR}_{T}}{\text{SNR}_{t}}\right)$
$n_{t}$ $=\frac{\text{SNR}_{T}}{\text{SNR}_{t}}\frac{\alpha_{t}}{\alpha_{T}}$
$\sigma_{t}$ $=\sigma_{t}\sqrt{1-\frac{\text{SNR}_{T}}{\text{SNR}_{t}}},$
where $\text{SNR}_{t},\alpha_{t},\sigma_{t}^{2}$ are all ingredients defined in the reference process $\mathbb{Q}$.
4 Method
4.1 Problem Formulation
We formulate different dense prediction tasks like depth estimation, surface normal prediction, semantic segmentation etc., as image-conditioned generation problems and propose DPBridge to model the conditional distributions $p(\mathbf{y}|\mathbf{x})$, where $\mathbf{y}\in\mathbb{R}^{H\times W\times C}$ indicates the desired dense signal map (e.g. segmentation, depth, surface normal), and $\mathbf{x}\in\mathbb{R}^{H\times W\times 3}$ is an RGB image. In the following sections, we’ll first introduce the motivation of utilizing diffusion bridge models to handle dense prediction problems. Then, we’ll elaborate the details of our proposed DPBridge framework, which is constructed based on pretrained diffusion backbone through proper finetuning.
Figure 2: Overview of our training pipeline, which is conducted in the latent space based on diffusion bridge theory. To overcome the discrepancy with standard diffusion process and facilitate finetuning, we derive the maximum likelihood training objective of diffusion bridge process, propose a distribution aligned normalization strategy, and assist the training with an auxiliary image consistency loss.
4.2 Motivation: Bridge VS Diffusion
As introduced previously, the incompatibility that employ diffusion models to dense prediction task, which is essentially a image-conditioned generation problem, lies in its underlying iterative denoising process starting from Gaussian noise. In such cases, the distance between prior Gaussian distribution and data distribution remains large, requiring inevitably more sampling steps as well as ad-hoc conditioning mechanisms to converge to target distributions, at the risk of degraded performance and slow inference speed.
Bridge models, on the other hand, directly formulate generative processes as conditional probability paths. In this scenario, the boundary distribution is not restricted to Gaussian noise, but instead can be chosen as more informative priors, such as the condition data distribution. By establishing direct connections between different distributions, this formulation provides a more natural and efficient way to model probability transport compared with previous efforts. For example, it reduce the dependence on manually designed post-hoc guidance [11, 53, 12] or customized generation pipeline [41]. In addition, through implicitly integrates condition as part of the learned process, it provides a more flexible way to model intricacies between data and conditions without relying on explicit mechanisms like concatenation, attention [47, 61, 25], whose inductive bias may harm the overall performance when dealing with complex circumstances. Considering these advantages, we advocate that adopting diffusion bridge model will be a better fit to dense prediction tasks compared with conventional diffusion-based counterparts.
4.3 Dense Prediction Bridge
Given paired training data, a naive implementation is to construct the diffusion bridge model from scratch according to Sec.3.2. However, this approach is extremely resource-demanding, making it hard to implement with limited computation power. What’s more severe, it lacks the ability to harness rich scene understanding knowledge inherent in the vision foundation models, which have been proven significant in terms of boosting both perception quality and generalization performance in dense prediction tasks. Therefore, our DPBridge is build upon pretrained diffusion backbones through finetuning and we explore several strategies to handle the misalignment between original diffusion bridge process and standard diffusion process, aiming for more efficient knowledge transfer. Fig.2 gives an overview of the proposed finetuning pipeline, and we’ll explain the design of each component next.
4.4 Latent Space Implementation
Instead of directly learning in image pixel space, we follow common practices to perform diffusion steps in a compressed but perceptually aligned latent space, provided by an independently trained variational autoencoder (VAE). Specifically, both the target dense maps and RGB images are converted by the encoder to obtain $\mathbf{z}_{0}=\mathcal{E}(\mathbf{y})$ and $\mathbf{z}_{T}=\mathcal{E}(\mathbf{x})$, serving as the boundary states. In this way the diffusion bridge process is henceforth learned in the latent space. To retrieve original signal map we can use the decoder to decode the predicted latents $\mathbf{y}=\mathcal{D}(\hat{\mathbf{z}}_{0})$. In later part of this paper, with little abuse of notation, all derivations are conducted in the latent space.
4.5 Training Objective
4.5.1 Maximum Likelihood Loss
In previous diffusion bridge models, widely-used techniques for learning the reverse process that corresponds to Eqn.(2) is to follow the theory of [1], use a neural network to approximate the score function and apply score matching technique. We notice that analytical form of diffusion bridge’s reverse transition kernel is directly obtainable, therefore the usage of maximum likelihood training can also be applied, which is performed by optimizing the Evidence Lower Bound (ELBO). It yields a more stable loss function compared with score matching and better aligns with the pretrained backbone. Given the fixed terminal state $\mathbf{z}_{T}$, the expectation of data log-likelihood possesses the following ELBO formulation:
$\text{ELBO}=&\mathbb{E}_{p(\mathbf{z}_{0})}\Big{[}\mathbb{E}_{p(\mathbf{z}_{1}\mid\mathbf{z}_{0})}\big{[}\log p_{\theta}(\mathbf{z}_{0}\mid\mathbf{z}_{1},\mathbf{z}_{T})\big{]}-\\ \sum_{t=2}^{T}\mathbb{E}_{p(\mathbf{z}_{t}|\mathbf{z}_{0})}&\big{[}\mathbb{D}_{KL}\big{(}q(\mathbf{z}_{t-1}\mid\mathbf{z}_{0},\mathbf{z}_{t},\mathbf{z}_{T})\,||\,p_{\theta}(\mathbf{z}_{t-1}\mid\mathbf{z}_{t},\mathbf{z}_{T})\big{)}\big{]}\Big{]}\end{aligned}$
where $q(\mathbf{z}_{t-1}|\mathbf{z}_{t},\mathbf{z}_{0},\mathbf{z}_{T})$ and $p_{\theta}(\mathbf{z}_{t-1}|\mathbf{z}_{t},\mathbf{z}_{T})$ are the transition kernel of true reverse process and our modeled generative process and $\mathbb{D}_{KL}$ is the KL divergence. Utilize Bayes’ theorem and Markov chain property, we obtain:
$q(\mathbf{z}_{t-1}|\mathbf{z}_{t},\mathbf{z}_{0},\mathbf{z}_{T})$ $=\frac{q(\mathbf{z}_{t}|\mathbf{z}_{t-1},\mathbf{z}_{T})q(\mathbf{z}_{t-1}|\mathbf{z}_{0},\mathbf{z}_{T})}{q(\mathbf{z}_{t}|\mathbf{z}_{0},\mathbf{z}_{T})}$ (5)
$=\mathcal{N}(\tilde{\boldsymbol{\mu}}_{t}(\mathbf{z}_{t},\mathbf{z}_{0},\mathbf{z}_{T}),\tilde{\sigma}_{t}^{2}\mathbf{I})$
For simplicity, we define the following notations: $a_{t}=m_{t}/m_{t-1},b_{t}=n_{t}-a_{t}\cdot n_{t-1},\delta_{t}^{2}=\sigma_{t}^{2}-a_{t}^{2}\cdot\sigma_{t-1}^{2}$. The mean value term is:
$\tilde{\boldsymbol{\mu}}_{t}(\mathbf{z}_{t},\mathbf{z}_{0},\mathbf{z}_{T})$ $=k_{1}\mathbf{z}_{t}+k_{2}\mathbf{z}_{0}+k_{3}\mathbf{z}_{T}$ (6)
$k_{1}$ $=\frac{\sigma_{t-1}^{2}}{\sigma_{t}^{2}}\cdot a_{t}$
$k_{2}$ $=\frac{\delta_{t}^{2}}{\sigma_{t}^{2}}\cdot m_{t-1}$
$k_{3}$ $=\frac{\delta_{t}^{2}}{\sigma_{t}^{2}}\cdot n_{t-1}-\frac{\sigma_{t-1}^{2}}{\sigma_{t}^{2}}\cdot a_{t}\cdot b_{t},$
and the variance is:
$\tilde{\sigma}_{t}^{2}=\frac{\sigma_{t-1}^{2}}{\sigma_{t}^{2}}\cdot\delta_{t}^{2}$ (7)
Assume $p_{\theta}(\mathbf{z}_{t-1}|\mathbf{z}_{t},\mathbf{z}_{T})$ is also a Gaussian distribution parameterized as $\mathcal{N}(\boldsymbol{\mu}_{\theta}(\mathbf{z}_{t},\mathbf{z}_{T},t),\boldsymbol{\Sigma}_{\theta}(\mathbf{z}_{t},\mathbf{z}_{T},t))$ and let $\boldsymbol{\Sigma}_{\theta}(\mathbf{z}_{t},\mathbf{z}_{T},t))=\tilde{\sigma}_{t}^{2}$ for derivation simplicity, maximizing the ELBO is equivalent to minimizing:
$\mathcal{L}_{elbo}=\mathbb{E}_{t,\mathbf{z}_{0},\mathbf{z}_{t},\mathbf{z}_{T}}\left[\frac{1}{2\tilde{\sigma}_{t}^{2}}\left\|\tilde{\boldsymbol{\mu}}_{t}-\boldsymbol{\mu}_{\theta}(\mathbf{z}_{t},\mathbf{z}_{T},t)\right\|^{2}\right]$ (8)
Similar to the strategy in DDPM [26], we use a network $\boldsymbol{\epsilon}_{\theta}(\mathbf{z}_{t},\mathbf{z}_{T},t)$ to parameterize noise instead of posterior mean and further simplified the loss as:
$\mathcal{L}_{elbo}=\mathbb{E}_{t,\mathbf{z}_{0},\mathbf{z}_{t},\mathbf{z}_{T}}\left[\left\|\boldsymbol{\epsilon}-\boldsymbol{\epsilon}_{\theta}(\mathbf{z}_{t},\mathbf{z}_{T},t)\right\|^{2}\right],$ (9)
4.5.2 Image Consistency Loss
In addition to the common likelihood loss evaluated in the latent space, we also incorporate an image-level consistency loss, where we first convert the latent prediction back to image space, and then evaluate its consistency with the ground truth dense map using a simple MSE loss.
$\mathcal{L}_{ic}=\left\|\mathcal{D}(\hat{\mathbf{z}}_{0})-\mathbf{y})\right\|^{2}$ (10)
The reason is that the compression of input data in the latent space will lead to a loss of high-frequency information, harming the model’s ability of preserving fine-grained spatial details. By introducing an auxiliary loss at image level, we ensure that the model is capable of retaining spatial information which is essential for precise per-pixel prediction. This dual loss strategy leverages the compactness of latent space for efficiency while still anchoring the training process to the fidelity of full-resolution dense map prediction, thus leading to performance improvement. Combined these two losses together, our final training objective becomes:
$\mathcal{L}=\omega_{1}\mathcal{L}_{elbo}+\omega_{2}\mathcal{L}_{ic}$ (11)
where $\omega_{1},\omega_{2}$ are adjustable hyperparameters to balance the loss. An overview of DPBridge’s training pipeline shown in Algorithm.1.
Algorithm 1 Training
1:Encoder $\mathcal{E}$; Decoder $\mathcal{D}$; Loss weights $\omega_{1},\omega_{2}$
2:repeat
3:     Retrieve paired data $(\mathbf{x},\mathbf{y})$
4:     $\mathbf{z}_{0}=\mathcal{E}(\mathbf{y})$, $\mathbf{z}_{T}=\mathcal{E}(\mathbf{x})$
5:     $t\sim\text{Uniform}(\{1,\dots,T\})$, $\boldsymbol{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$
6:     $\mathbf{z}_{t}=n_{t}\mathbf{z}_{T}+m_{t}\mathbf{z}_{0}+\sigma_{t}\boldsymbol{\epsilon}$
7:     $\mathbf{z}_{t}^{\prime}=\frac{1}{\sqrt{n_{t}^{2}+\sigma_{t}^{2}}}(\mathbf{z}_{t}-n_{t}\mathbf{z}_{T})$
8:     $\mathcal{L}_{elbo}=\left\|\boldsymbol{\epsilon}-\boldsymbol{\epsilon}_{\theta}(\mathbf{z}_{t}^{\prime},\mathbf{z}_{T},t)\right\|^{2},$
9:     $\hat{\mathbf{z}}_{0}=\frac{1}{m_{t}}(\mathbf{z}_{t}-n_{t}\mathbf{z}_{T}-\sigma_{t}\boldsymbol{\epsilon}_{\theta}(\mathbf{z}_{t}^{\prime},\mathbf{z}_{T},t))$
10:     $\mathcal{L}_{ic}=\left\|\mathcal{D}(\hat{\mathbf{z}}_{0})-\mathbf{y}\right\|^{2}$
11:     Take gradient step on $\nabla_{\theta}\left(\omega_{1}\mathcal{L}_{elbo}+\omega_{2}\mathcal{L}_{ic}\right)$
12:until converged
Algorithm 2 Inference
1:Encoder $\mathcal{E}$; Decoder $\mathcal{D}$; Trained $\boldsymbol{\epsilon}_{\theta}(\cdot,\cdot,\cdot)$
2:$\mathbf{z}_{T}=\mathcal{E}(\mathbf{x})$
3:for $t=T,\dots,1$ do
4:     $\boldsymbol{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ if $t>1$, else $\boldsymbol{\epsilon}=\mathbf{0}$
5:     $\mathbf{z}_{t}^{\prime}=\frac{1}{\sqrt{n_{t}^{2}+\sigma_{t}^{2}}}(\mathbf{z}_{t}-n_{t}\mathbf{z}_{T})$
6:     $\hat{\mathbf{z}}_{0}=\frac{1}{m_{t}}(\mathbf{z}_{t}-n_{t}\mathbf{z}_{T}-\sigma_{t}\boldsymbol{\epsilon}_{\theta}(\mathbf{z}_{t}^{\prime},\mathbf{z}_{T},t))$
7:     $\mathbf{z}_{t-1}=k_{1}\mathbf{z}_{t}+k_{2}\hat{\mathbf{z}}_{0}+k_{3}\mathbf{z}_{T}+\tilde{\sigma}_{t}\boldsymbol{\epsilon}$
8:end for
9:$\mathbf{y}=\mathcal{D}(\hat{\mathbf{z}}_{0})$
10:return $\mathbf{y}$
4.6 Distribution Alignment Normalization
A notable challenge during finetuning is the misalignment of intermediate sample’s marginal distribution $q(\mathbf{z}_{t})$ between DPBridge and diffusion backbone, due to the difference in the forward transition kernel. This misalignment will weaken the model’s inherent denoising ability, thus limiting the effectiveness of knowledge transfer. To address this issue, we propose a distribution alignment normalization technique to fill in the gap between the two distributions. Specifically, we normalize the intermediate latents computed from Eqn.(3) to match the form of a standard diffusion process:
$\mathbf{z}_{t}^{\prime}=\frac{\mathbf{z}_{t}-m_{t}\mathbf{z}_{T}}{\sqrt{n_{t}^{2}+\sigma_{t}^{2}}}=\frac{n_{t}}{\sqrt{n_{t}^{2}+\sigma_{t}^{2}}}\mathbf{z}_{0}+\frac{\sigma_{t}}{\sqrt{n_{t}^{2}+\sigma_{t}^{2}}}\boldsymbol{\epsilon}_{t}$ (12)
Since the square sum of coefficients is equal to 1, the normalized bridge process $\{\mathbf{z}_{t}^{\prime}:t\in[0,T]\}$ is the same as a standard VP process, making the marginal distribution of these two aligned. Therefore by feeding the normalized sample $\mathbf{z}_{t}^{\prime}$ in the network $\boldsymbol{\epsilon}_{\theta}$, we can fully exploit the denoising capability from pretrained diffusion backbone.
4.7 Network Architecture
To better utilize visual information ($\mathbf{z}_{T}$ in our case), we concatenating both intermediate and terminal state samples along the feature dimension and send in the noise prediction network as a whole. To address the dimension incompatibility with pretrained backbone, the input channels of the network are thus doubled to accommodate the concatenated input. In addition, as suggested by [32], we duplicate the weight tensor of these accommodated layers and divide the values by two to preserve the input magnitude to subsequent layers. As DPBridge is an image-conditioned generative framework, we employ null-text embedding for all cross-attention layers to eliminate the impact of text prompts.
4.8 Accelerate Inference
Similar to the idea of DDIM [52], the inference process of DPBridge can be accelerated by utilizing a non-Markovian process while maintaining the same marginal distributions for all intermediate samples along the sampling path. Specifically, we can sample on a subset of timesteps $\{t_{i}\}_{i=1}^{N}$, where $0=t_{0}<t_{1}<\cdots<t_{N-1}<t_{N}=T$ and the number of sampling steps $N$ can be set independently from the original timesteps $T$ on which the bridge model is trained. The accelerated inference process can be thus described as:
$\begin{split}\mathbf{z}_{t_{n}}&=n_{t_{n}}\mathbf{z}_{T}+m_{t_{n}}\hat{\mathbf{z}}_{0}+\\ &\sqrt{\sigma_{t_{n}}^{2}-g_{n}^{2}}\frac{\mathbf{z}_{t_{n+1}}-n_{t_{n+1}}\mathbf{z}_{T}-m_{t_{n+1}}\hat{\mathbf{z}}_{0}}{\sigma_{t_{n+1}}}+g_{n}\boldsymbol{\epsilon}\end{split}$ (13)
where $g_{n}$ is the variance parameter that controls the stochasticity behavior of the reverse sampling process. We note that when $g_{n}=\tilde{\sigma}_{t}^{2}$ which is calculated from Eqn.(5), the process becomes a Markovian bridge. The basic inference process of DPBridge is summarized in Algorithm.2. If adopt acceleration, one can utilize Eqn.(13) during implementation.
5 Experiments
We evaluate the proposed DPBridge approach mainly on two representative dense prediction tasks, i.e., depth estimation and surface normal prediction. We’ll start by a brief introduction of experiment settings of these two tasks, followed by quantitative and qualitative analysis. Lastly, we conduct detailed ablation studies and investigate the influence each design component respectively.
5.1 Experiments Setup
Datasets. The basic setting for depth estimation and surface normal prediction experiments are similar. We use two synthetic datasets, HyperSim [45] and Virtual KITTI [8], which cover both indoor and outdoor scenes, for training. To comprehensively evaluate the model performance and demonstrate its generalization capability, we conduct zero-shot evaluations on real-world datasets. For depth estimation, this includes NYUv2 [51], KITTI [22], ScanNet [15], ETH3D [49] and DIODE [56]. For surface normal prediction, we additionally incorporate iBims [33] and Sintel [7] datasets in the evaluation pipeline. For fair comparison, we use the same validation split as in previous methods [32].
Evaluation Metrics. For depth estimation, we follow the affine-invariant evaluation protocol, where we first align the prediction results to the ground truth with least squares fitting, and then use Absolute Mean Relative Error (AbsRel) and $\delta_{1}$ accuracy to assess the performance. Denote the ground truth depth $\mathbf{d}$, predicted depth $\hat{\mathbf{d}}$, and the number of valid pixels $N$, the AbsRel is calculated as $\frac{1}{N}\sum_{i=1}^{N}\|\mathbf{d}_{i}-\hat{\mathbf{d}}_{i}\|/\mathbf{d}_{i}$, and $\delta_{1}$ accuracy is measured using the ratio of pixels satisfying $\text{max}(\mathbf{d}_{i}/\hat{\mathbf{d}}_{i},\hat{\mathbf{d}}_{i}/\mathbf{d}_{i})<1.25$. For the metrics of surface normal task, we report the commonly used mean angular error between the ground-truth normal vectors and the prediction results, as well as the percentage of pixels with an angular error below 11.25 degrees.
5.2 Implementation Details
We implement DPBridge based on Stable Diffusion (SD) v1.5 [46] as backbone. During training, the number of timesteps of is set to be 1000, and we use 50 sampling steps during inference with considerations of both quality and efficiency. Training our method takes 30K iterations using the Adam optimizer with a base learning rate of $3\times 10^{-5}$ and an exponential decay strategy after 100-step warm-up. The batch size is 2, but with 8 steps gradient accumulation, the effective batch size is equivalent to 16. We also apply random horizontal flipping augmentation during training. The whole process takes approximately 2 days to complete on a single Nvidia RTX 3090Ti GPU.
Method Model Type NYUv2 KITTI ETH3D ScanNet DIODE
[…]
Marigold [32] 8.6 93.6 13.8 82.6 9.3 93.2 10.4 89.7 33.6 73.9
Geowizard [20] 8.5 93.7 16.2 75.6 9.1 93.1 9.0 93.0 32.0 72.2
E2E-FT [21] 8.2 94.0 15.4 77.9 10.6 89.8 9.2 92.3 32.9 73.7
ControlNet [61] 45.5 54.6 55.2 40.2 30.5 58.8 25.7 68.3 52.1 50.1
InstructPix2Pix [6] 35.2 60.0 47.7 51.2 26.1 66.7 21.4 75.3 44.8 58.0
I2SB [37] 24.9 73.4 37.7 57.9 40.6 52.0 21.2 66.1 40.2 59.1
BBDM [36] 27.9 65.5 47.2 45.1 46.4 48.8 22.7 67.0 34.6 64.7
DPBridge (Ours) 6.9 95.7 11.2 87.3 8.1 93.2 7.5 94.7 29.7 78.6
Table 1: Quantitative comparison of DPBridge with state-of-the-art affine-invariant depth estimation methods on zero-shot benchmarks. All metrics are presented in percentage terms with best and second-best results marked by bold and underline. DM and BM correspond to diffusion models and bridge models respectively. All models are trained using SDv1.5 as backbone.
5.3 Quantitative and Qualitative Results
Tab.1 and Tab.2 summarize the quantitative metrics of DPBridge compared with diffusion-based and bridge-based counterparts, who are most closely related to our work. The reason we use SDv1.5 instead of more advanced SDv2.1 as backbone is that our current pipeline does not support v-prediction parameterization as in SDv2.1, and this incompatibility will undermine the model’s knowledge transfer ability. We’ll present more detail analysis in the ablation study. For a fair comparison, we re-implement all baseline methods using the same SDv1.5 backbone as we do, and report the results. We also conduct comparison with feed-forward and diffusion-based approaches based on SDv2.1 backbone, and the complete results is presented in the Appendix. DPBridge outperforms most methods in our evaluation setting and even be comparable with state-of-the-art that is trained with SDv2.1. Specifically, when compared with bridge models like I2SB and BBDM that are trained from scratch, we exceed their performance by a large margin, justifying our proposal that adapting prior knowledge to bridge process indeed benefits the perception performance. In terms of diffusion-based methods, DPBridge can also achieve the best result, showcasing that adopting bridge process to handle dense prediction tasks is indeed advantageous over standard diffusion framework.
Figure 3: Qualitative comparison of monocular depth estimation methods on KITTI and ETH3D dataset. DPBridge is able to generalize in both indoor and outdoor scenes, successively capturing fine-grained details such as leaves and structures on the fence (KITTI).
Figure 4: Qualitative comparison between our (above) and Marigold trained on SDv1.5 (below) on normal estimation results, DPBridge is able to capture more textures from the photos in-the-wild.
The qualitative results for depth estimation is conducted on KITTI and ETH3D datasets, as shown in Fig. 3. We can see that E2E-FT tends to produce blurry and smooth predictions, while Geowizard fails to provide accurate estimation in relatively long distance range. In contrast, DPBridge perform well in both indoor and outdoor scenes, capturing fine-level details in both near and faraway regions, like the contour of leaves, fences, cars, computers or windows. We also present visualizations on in-the-wild images for surface normal prediction task. DPBridge is able to predict high-quality surface normals equipped with fine-grained textures. This strong generalization capability can be attribute to that we effectively harness extensive visual prior knowledge inherent in the pretrained diffusion backbone.
Method NYUv2 ScanNet iBims-1 Sintel
[…]
Marigold [32] 22.3 40.0 23.3 37.6 28.0 38.5 52.1 8.1
Geowizard [20] 21.7 44.2 22.0 42.1 28.2 37.9 45.8 11.0
E2E-FT [21] 17.7 57.8 17.4 57.4 18.0 64.6 43.9 14.1
DPBridge (Ours) 18.1 54.2 18.0 54.5 24.3 51.2 41.2 19.4
Table 2: Quantitative comparison of DPBridge with state-of-the-art surface normal prediction methods on zero-shot benchmarks. All metrics are presented in percentage terms with best and second-best results marked by bold and underline. DM and BM correspond to diffusion models and bridge models respectively. All models are trained using SDv1.5 as backbone.
Design Components NYUv2 KITTI
[…]
26.8 85.7 33.6 57.7
8.8 93.0 14.0 83.1
8.5 93.3 13.4 83.9
7.5 94.9 13.1 84.6
6.9 95.7 11.2 87.3
Table 3: The ablation study regarding the design components of DPBridge. FT stands for utilizing pretrained diffusion backbone during training. DAN denotes distribution aligned normalization technique. IC indicates image consistency loss.
Parameterization Backbone Steps NYUv2 KITTI
[…]
50 8.2 94.5 12.7 85.0
50 23.7 59.1 30.5 56.7
50 7.6 95.6 12.0 85.8
50 7.7 94.6 12.4 85.2
50 22.8 62.4 28.5 60.1
5 12.1 86.5 19.4 66.5
10 7.7 95.1 12.5 84.1
50 6.9 95.7 11.2 87.3
100 6.8 95.1 11.4 87.5
200 6.6 95.6 12.3 84.8
Table 4: The ablation study regarding the hyperparameter choice of diffusion bridge process. Options include different network parameterizations, the version of SD to use as backbone, and the number of denoising steps during inference.
5.4 Ablation Studies
To evaluate the effectiveness of each design component and training setting, we conduct ablation studies using depth estimation tasks on the NYUv2 and KITTI datasets. The results are summarized in Tab.3 and Tab.4, with detailed discussions provided below.
5.4.1 Design Components
Finetuning VS Training From Scratch. To assess whether adapting pretrained diffusion backbones to our DPBridge improves perception performance, we compare the difference between finetuning and training from scratch paradigm. As shown in the first row of Tab.3, fine-tuning significantly outperforms training from scratch under identical experimental settings. This difference is due to two factors: first, foundation models trained on large, diverse datasets capture universal visual priors, aiding knowledge transfer to dense prediction tasks. Second, training from scratch requires more iterations for convergence. These observations suggest that building DPBridge on pretrained backbones is indeed advantageous.
Distribution Alignment Normalization. Next, we evaluate whether aligning the marginal distributions between diffusion bridge process and standard diffusion process is beneficial. As shown in the second and fourth rows of Tab.3, incorporating a distribution alignment normalization step before inputting data into the noise prediction network noticeably improves performance.
Image Consistency Loss. Regarding the auxiliary image consistency loss, the comparison between the third and fourth rows in Tab.3 shows that while effective, its impact is less significant than that of distribution alignment normalization. This is probably because it primarily influences regions where masked and valid areas overlap, which constitute only a relatively small proportion of the data. However, when integrated with other design components, the overall performance is significantly improved.
5.4.2 Bridge Process Hyperparameters
Parameterization and Backbone Selection. Apart from the noise prediction scheme used in our main experiments, we also consider how the choice of parameterization combined with the backbone selection influences performance. Results in Tab.4 reveal two key findings: (1) v-parameterization performs significantly worse compared with other two options, highlighting the need for further research on deriving objectives specific to the diffusion bridge process, and (2) ”noise prediction + SDv1.5” achieves the best results, indicating that the parameterization during fine-tuning should align with that of the backbone to maximize knowledge transfer efficiency.
Number of Sampling Steps. We evaluate performance with varying numbers of sampling steps during inference. As shown in the lower part of Tab.4, increasing sampling steps initially improves perception performance rapidly before leveling off. Notably, compared to the results in Tab.1, where all other diffusion-based methods are evaluated with 50 steps setting, DPBridge achieves comparable performance with fewer steps, demonstrating the efficiency of our proposed method.
6 Conclusion
In this paper, we present DPBridge, a latent diffusion bridge model to tackle general dense prediction tasks. We establish a direct mapping between input RGB images and their corresponding dense maps in the latent space based on fully-tractable diffusion bridge process. In this way, the prior distribution is no longer constrained to uninformative noise, but instead can be chosen as more structurally correlated ones. In addition, we introduce effective finetuning strategies to adapt large pretrained diffusion backbone to our bridge model construction, leveraging its rich prior knowledge to boost training speed and achieve remarkable perception performance. Future research directions include deriving v-prediction parameterization of diffusion bridge process, and extending to other dense prediction tasks.
Supplementary Material
In the supplementary material, we’ll provide contents that are not included in the main paper due to page limitations. We first fill in the derivation details of the loss function. Afterwards we will provide the additional quantitative comparisons with more baseline methods. Finally, we will present qualitative examples that apply DPBridge on more dense prediction tasks.
Appendix A Derivation Details
In this section, we’ll provide derivation details of the maximum likelihood loss presented in Sec.4.5.1 . Its core is the posterior distribution $q(\mathbf{z}_{t-1}|\mathbf{z}_{t},\mathbf{z}_{0},\mathbf{z}_{T})$ as in Eqn.(5) , whose calculation involves three terms: $q(\mathbf{z}_{t}|\mathbf{z}_{t-1},\mathbf{z}_{T})$, $q(\mathbf{z}_{t-1}|\mathbf{z}_{0},\mathbf{z}_{T})$, and $q(\mathbf{z}_{t}|\mathbf{z}_{0},z\mathbf{z}_{T})$. The last two term of the posterior can be retrieved directly through the marginal distribution of diffusion bridge’s forward process, but further derivation is required for the first term. Specifically, as specified in Eqn.(3)
$\begin{split}q(\mathbf{z}_{t}|\mathbf{z}_{0},\mathbf{z}_{T})&=\mathcal{N}(m_{t}\mathbf{z}_{0}+n_{t}\mathbf{z}_{T},\sigma_{t}^{2}\mathbf{I})\\ \mathbf{z}_{t}&=m_{t}\mathbf{z}_{0}+n_{t}\mathbf{z}_{T}+\sigma_{t}\boldsymbol{\epsilon}_{t}\end{split}$ (14)
$\begin{split}q(\mathbf{z}_{t-1}|\mathbf{z}_{0},\mathbf{z}_{T})&=\mathcal{N}(m_{t-1}\mathbf{z}_{0}+n_{t-1}\mathbf{z}_{T},\sigma_{t-1}^{2}\mathbf{I})\\ \mathbf{z}_{t-1}&=m_{t-1}\mathbf{z}_{0}+n_{t-1}\mathbf{z}_{T}+\sigma_{t-1}\boldsymbol{\epsilon}_{t-1}\end{split}$ (15)
To derive the transition probability $q(\mathbf{z}_{t}|\mathbf{z}_{t-1},\mathbf{z}_{T})$, we substitute $\mathbf{z}_{0}$ in Eqn.(14) with derivation from Eqn.(15), obtain
$\mathbf{z}_{t}$ $=a_{t}\mathbf{z}_{t-1}+b_{t}\mathbf{z}_{T}+\delta_{t}\boldsymbol{\epsilon}$ (16)
$a_{t}$ $=\frac{m_{t}}{m_{t-1}}$
$b_{t}$ $=n_{t}-\frac{m_{t}}{m_{t-1}}n_{t-1}=n_{t}-a_{t}n_{t-1}$
$\delta_{t}^{2}$ $=\sigma_{t}^{2}-\frac{m_{t}}{m_{t-1}}n_{t-1}^{2}\sigma_{t-1}^{2}=\sigma_{t}^{2}-a_{t}^{2}\sigma_{t-1}^{2}$
Thus, we can get $q(\mathbf{z}_{t}|\mathbf{z}_{t-1},\mathbf{z}_{T})=\mathcal{N}(a_{t}\mathbf{z}_{t-1}+b_{t}\mathbf{z}_{T},\delta_{t}^{2}\mathbf{I})$, which is also a Gaussian distribution.
Putting all these terms back to Eqn.(5) , we’ll have:
$q(\mathbf{z}_{t-1}|\mathbf{z}_{t},\mathbf{z}_{0},\mathbf{z}_{T})\sim\mathcal{N}(\tilde{\boldsymbol{\mu}}_{t}(\mathbf{z}_{t},\mathbf{z}_{0},\mathbf{z}_{T}),\tilde{\sigma}_{t}^{2}\mathbf{I})$ (17)
$=$ $\frac{q(\mathbf{z}_{t}|\mathbf{z}_{t-1},\mathbf{z}_{T})q(\mathbf{z}_{t-1}|\mathbf{z}_{0},\mathbf{z}_{T})}{q(\mathbf{z}_{t}|\mathbf{z}_{0},\mathbf{z}_{T})}$
$\propto$ $\exp\bigg{[}-\frac{1}{2}\bigg{(}\frac{(\mathbf{z}_{t}-a_{t}\mathbf{z}_{t-1}-b_{t}\mathbf{z}_{T})^{2}}{\delta_{t}^{2}}+\frac{(\mathbf{z}_{t-1}-m_{t-1}\mathbf{z}_{0}-n_{t-1}\mathbf{z}_{T}^{2}}{\sigma_{t-1}^{2}}-\frac{(\mathbf{z}_{t}-m_{t}\mathbf{z}_{0}-n_{t}\mathbf{z}_{T}^{2}}{\sigma_{t}^{2}}\bigg{)}\bigg{]}$
$=$ $\exp\bigg{[}-\frac{1}{2}\bigg{(}\bigg{(}\frac{a_{t}^{2}}{\delta_{t}^{2}}+\frac{1}{\sigma_{t-1}^{2}}\bigg{)}\mathbf{z}_{t-1}^{2}-2\bigg{(}\frac{a_{t}\mathbf{z}_{t}-a_{t}b_{t}\mathbf{z}_{T}}{\delta_{t}^{2}}+\frac{m_{t-1}\mathbf{z}_{0}+n_{t-1}\mathbf{z}_{T}}{\sigma_{t-1}^{2}}\bigg{)}\mathbf{z}_{t-1}+C(\mathbf{z}_{t},\mathbf{z}_{0},\mathbf{z}_{T})\bigg{)}\bigg{]}.$
where $C(\mathbf{z}_{t},\mathbf{z}_{0},\mathbf{z}_{T})$ indicates the constant term that is independent of $\mathbf{z}_{t-1}$. Since the Gaussian distribution probability density $\mathbf{z}_{t-1}\sim\mathcal{N}(\tilde{\boldsymbol{\mu}}_{t},\tilde{\sigma}_{t}^{2}\mathbf{I})$ can be expressed in the following form:
$\exp\left(-\frac{(\mathbf{z}_{t-1}-\tilde{\boldsymbol{\mu}}_{t})^{2}}{2\tilde{\sigma}_{t}^{2}}\right)=\exp\left[-\frac{1}{2}\left(\frac{1}{\tilde{\sigma}_{t}^{2}}\mathbf{z}_{t-1}^{2}-\frac{2\tilde{\boldsymbol{\mu}}_{t}}{\tilde{\sigma}_{t}^{2}}\mathbf{z}_{t-1}+\frac{\tilde{\boldsymbol{\mu}}_{t}^{2}}{\tilde{\sigma}_{t}^{2}}\right)\right].$ (18)
We can thus derive:
$\frac{1}{\tilde{\sigma}_{t}^{2}}$ $=\frac{a_{t}^{2}}{\delta_{t}^{2}}+\frac{1}{\sigma_{t-1}^{2}}$
$\Rightarrow\tilde{\sigma}_{t}^{2}$ $=\frac{\sigma_{t-1}^{2}(\sigma_{t}^{2}-a_{t}^{2}\sigma_{t-1}^{2})}{\sigma_{t}^{2}}=\frac{\sigma_{t-1}^{2}}{\sigma_{t}^{2}}\delta_{t}^{2}$ (19)
$\frac{\tilde{\boldsymbol{\mu}}_{t}}{\tilde{\sigma}_{t}^{2}}$ $=\frac{a_{t}\mathbf{z}_{t}-a_{t}b_{t}\mathbf{z}_{T}}{\delta_{t}^{2}}+\frac{m_{t-1}\mathbf{z}_{0}+n_{t-1}\mathbf{z}_{T}}{\sigma_{t-1}^{2}}$
$\Rightarrow\tilde{\boldsymbol{\mu}}_{t}$ $=\bigg{[}\frac{\sigma_{t-1}^{2}\big{(}a_{t}\mathbf{z}_{t}-a_{t}b_{t}\mathbf{z}_{T}\big{)}+\delta_{t}^{2}\big{(}m_{t-1}\mathbf{z}_{0}+n_{t-1}\mathbf{z}_{T}\big{)}}{\delta_{t}^{2}\sigma_{t-1}^{2}}\bigg{]}\frac{\sigma_{t-1}^{2}}{\sigma_{t}^{2}}\delta_{t}^{2}$
$=\frac{\sigma_{t-1}^{2}}{\sigma_{t}^{2}}a_{t}\mathbf{z}_{t}+\frac{\delta_{t}^{2}}{\sigma_{t}^{2}}m_{t-1}\mathbf{z}_{0}+\bigg{(}\frac{\delta_{t}^{2}}{\sigma_{t}^{2}}n_{t-1}-\frac{\sigma_{t-1}^{2}}{\sigma_{t}^{2}}a_{t}b_{t}\bigg{)}\mathbf{z}_{T}$ (20)
whose results are equivalent to Eqn.(6) and Eqn.(7).
Accordingly, maximizing the Evidence Lower Bound (ELBO) is equivalent to minimizing the KL divergence between ground truth and predicted posterior:
$KL\left(q(\mathbf{z}_{t-1}|\mathbf{z}_{0},\mathbf{z}_{t},\mathbf{z}_{T})\,||\,p_{\theta}(\mathbf{z}_{t-1}\mid\mathbf{z}_{t},\mathbf{z}_{T})\right)$ (21)
$=$ $\mathbb{E}_{q(\mathbf{z}_{t-1}|\mathbf{z}_{0},\mathbf{z}_{t},\mathbf{z}_{T})}\left[\log\frac{\frac{1}{\sqrt{2\pi}\tilde{\sigma}_{t-1}}e^{-\frac{(\mathbf{z}_{t-1}-\tilde{\boldsymbol{\mu}}_{t-1})^{2}}{2\tilde{\sigma}_{t-1}^{2}}}}{\frac{1}{\sqrt{2\pi}\tilde{\sigma}_{\theta,t-1}}e^{-\frac{(\mathbf{z}_{t-1}-\tilde{\boldsymbol{\mu}}_{\theta,t-1})^{2}}{2\tilde{\sigma}_{\theta,t-1}^{2}}}}\right]$
$=$ $\mathbb{E}_{q(\mathbf{z}_{t-1}|\mathbf{z}_{0},\mathbf{z}_{t},\mathbf{z}_{T})}\left[\log\tilde{\sigma}_{\theta,t-1}-\log\tilde{\sigma}_{t-1}-\frac{(x_{t-1}-\tilde{\boldsymbol{\mu}}_{t-1})^{2}}{2\tilde{\sigma}_{t-1}^{2}}+\frac{(x_{t-1}-\tilde{\boldsymbol{\mu}}_{\theta,t-1})^{2}}{2\tilde{\sigma}_{\theta,t-1}^{2}}\right]$
$=$ $\log\frac{\tilde{\sigma}_{\theta,t-1}}{\tilde{\sigma}_{t-1}}-\frac{1}{2}+\frac{\tilde{\sigma}_{t-1}^{2}}{2\tilde{\sigma}_{\theta,t-1}^{2}}+\frac{(\tilde{\boldsymbol{\mu}}_{t-1}-\tilde{\boldsymbol{\mu}}_{\theta,t-1})^{2}}{2\tilde{\sigma}_{\theta,t-1}^{2}}$
If we assume that only the mean term is learnable during training, and $\tilde{\sigma}_{\theta,t-1}=\tilde{\sigma}_{t-1}$, we can ignore all the unlearnable constants and arrive at the final training objective:
$\mathcal{L}$ $=\mathbb{E}_{t,\mathbf{z}_{0},\mathbf{z}_{t},\mathbf{z}_{T}}\left[\frac{1}{2\tilde{\sigma}_{t}^{2}}\left\|\tilde{\boldsymbol{\mu}}_{t-1}-\tilde{\boldsymbol{\mu}}_{\theta,t-1}(\mathbf{z}_{t},\mathbf{z}_{T},t)\right\|^{2}\right]$ (22)
$=\mathbb{E}_{t,\mathbf{z}_{0},\mathbf{z}_{t},\mathbf{z}_{T}}\left[\left\|\boldsymbol{\epsilon}-\boldsymbol{\epsilon}_{\theta}(\mathbf{z}_{t},\mathbf{z}_{T},t)\right\|^{2}\right]$
which is the same as in Eqn.(8) and Eqn.(9) . This concludes our derivation.
Appendix B Additional Quantitative Results
Next, we provide additional quantitative results on depth estimation (Tab.5) and surface normal prediction (Tab.6) tasks. Compared with Tab.1 and Tab.2 presented in the main paper, we conduct comparisons with more baseline methods, including feed-forward approaches and diffusion-based methods that is trained on more advanced Stable Diffusion (SD) v2.1 backbone. In depth estimation, DPBridge consistently matches state-of-the-art methods, while demonstrating robust generalization across datasets. For example, it achieves 6.9 AbsRel, 95.7 % $\delta_{1}$ on NYUv2, and 7.5 AbsRel, 94.7% $\delta_{1}$ on ScanNet. Similarly, for surface normal prediction, DPBridge also delivers competitive results, achieving mean error of 18.1 and $11.25^{\circ}$ accuracy of $54.2\%$ on NYUv2, as well as maintaining strong performance on other datasets like iBims-1 (mean: 24.3, $11.25^{\circ}:51.2\%$) and Sintel (mean: 41.2, $11.25^{\circ}:19.4\%$) and. These results highlight DPBridge’s versatility to generalize well across tasks and benchmarks, being comparable to both feed-forward and diffusion-based methods while offering a unified, efficient paradigm to handle dense prediction tasks.
Method Model Type NYUv2 KITTI ETH3D ScanNet DIODE
[…]
DiverseDepth [57] 11.7 87.5 19.0 70.4 22.8 69.4 10.9 88.2 37.6 63.1
MiDaS [43] 11.1 88.5 23.6 63.0 18.4 75.2 12.1 84.6 33.2 71.5
LeReS [58] 9.0 91.6 14.9 78.4 17.1 77.7 9.1 91.1 27.1 76.6
Omnidata [19] 7.4 94.5 14.9 83.5 16.6 77.8 7.5 93.2 33.9 74.2
HDN [60] 6.9 94.8 11.5 86.7 12.1 83.3 8.0 93.6 24.6 78.0
DPT [44] 9.8 90.3 10.0 90.1 7.8 94.6 8.2 93.8 18.2 75.8
Marigold* [32] 6.0 95.9 10.5 90.4 7.1 95.1 6.9 94.5 31.0 71.2
DepthFM* [24] 6.5 95.6 8.3 93.4 - - - - 22.5 80.0
Geowizard* [20] 5.7 96.2 14.4 82.0 7.5 94.3 6.1 95.8 31.4 77.1
E2E-FT* [21] 5.4 96.4 9.6 92.1 6.4 95.9 5.8 96.5 30.3 77.6
Marigold [32] 8.6 93.6 13.8 82.6 9.3 93.2 10.4 89.7 33.6 73.9
Geowizard [20] 8.5 93.7 16.2 75.6 9.1 93.1 9.0 93.0 32.0 72.2
E2E-FT [21] 8.2 94.0 15.4 77.9 10.6 89.8 9.2 92.3 32.9 73.7
ControlNet [61] 45.5 54.6 55.2 40.2 30.5 58.8 25.7 68.3 52.1 50.1
InstructPix2Pix [6] 35.2 60.0 47.7 51.2 26.1 66.7 21.4 75.3 44.8 58.0
I2SB [37] 24.9 73.4 37.7 57.9 40.6 52.0 21.2 66.1 40.2 59.1
BBDM [36] 27.9 65.5 47.2 45.1 46.4 48.8 22.7 67.0 34.6 64.7
DPBridge (Ours) 6.9 95.7 11.2 87.3 8.1 93.2 7.5 94.7 29.7 78.6
Table 5: Quantitative comparison of DPBridge with state-of-the-art affine-invariant depth estimation methods on zero-shot benchmarks. All metrics are presented in percentage terms. FF, DM, and BM correspond to feed-forward, diffusion models, and bridge models. The superscript * indicates that this model is trained based on Stable Diffusion v2.1, otherwise is based on Stable Diffusion v1.5.
Method Model Type NYUv2 ScanNet iBims-1 Sintel
[…]
EESNU [3] 16.2 58.6 - - 20.0 58.5 42.1 11.5
Omnidata v1 [19] 23.1 45.8 22.9 47.4 19.0 62.1 41.5 11.4
Omnidata v2 [31] 17.2 55.5 16.2 60.2 18.2 63.9 40.5 14.7
DSINE [2] 16.4 59.6 16.2 61.0 17.1 67.4 34.9 21.5
Marigold* [32] 18.8 55.9 17.7 58.8 18.4 64.3 39.1 14.9
Geowizard* [20] 19.1 49.5 17.3 53.7 19.5 61.6 40.4 13.2
E2E-FT* [21] 16.5 60.4 14.7 66.1 16.1 69.7 33.5 22.3
Marigold [32] 22.3 40.0 23.3 37.6 28.0 38.5 52.1 8.1
Geowizard [20] 21.7 44.2 22.0 42.1 28.2 37.9 45.8 11.0
E2E-FT [21] 17.7 57.8 17.4 57.4 18.0 64.6 43.9 14.1
DPBridge 18.1 54.2 18.0 54.5 24.3 51.2 41.2 19.4
Table 6: Quantitative comparison of DPBridge with state-of-the-art surface normal prediction methods on zero-shot benchmarks. All metrics are presented in percentage terms. FF, DM, and BM correspond to feed-forward, diffusion models, and bridge models. The superscript * indicates that this model is trained based on Stable Diffusion v2.1, otherwise is based on Stable Diffusion v1.5.
Furthermore, we also evaluate the robustness of DPBridge system by perturbing the input image with different types and different level intensities of noise. The evaluation is conducted under both depth estimation and surface normal prediction tasks. We use NYUv2 and KITTI to report the depth estimation results, as for surface normal prediction, we report the results on NYUv2 and iBims-1. Tab. 7 reveals that the model’s performance degrades as noise intensity increases, but the model performance remains acceptable under moderate-level-intensity noise. Among noise types, uniform noise shows the mildest effect on degradation, while salt & pepper and high-intensity gaussian noise cause the most significant performance decline. In addition, the noise intensity has more impact on surface normal prediction than depth estimation, showcasing the inherent characteristics between different tasks. These results emphasize the model’s robustness under moderate noise but still expose its vulnerability to extreme or specific types of noise.
Noise Type Noise Level Depth Estimation Surface Normal Prediction
NYUv2 KITTI NYUv2 iBims-1
[…]
Original Version - 6.9 95.7 11.2 87.3 18.1 54.2 24.3 51.2
Gaussian $\mathcal{N}(0,0.05)$ 7.2 94.4 11.8 86.0 18.6 52.2 25.6 45.1
$\mathcal{N}(0,0.1)$ 8.4 92.7 12.1 85.3 20.3 49.4 27.4 40.4
$\mathcal{N}(0,0.2)$ 10.3 90.0 15.7 78.6 24.6 43.7 37.0 25.4
$\mathcal{N}(0,0.5)$ 15.9 77.2 22.5 67.1 41.3 22.5 50.0 13.2
Uniform $\mathcal{U}(-0.05,0.05)$ 7.0 95.0 11.4 86.5 18.3 53.6 24.5 50.7
$\mathcal{U}(-0.1,0.1)$ 7.2 94.7 11.5 85.4 18.7 53.0 24.8 49.4
$\mathcal{U}(-0.2,0.2)$ 7.6 94.0 12.0 85.0 19.5 50.5 26.0 46.1
$\mathcal{U}(-0.5,0.5)$ 9.3 91.1 16.3 76.1 22.4 42.2 30.2 31.6
Poisson Poisson($\lambda=0.05$) 10.2 90.2 13.1 82.5 21.3 45.6 28.1 38.9
Poisson($\lambda=0.1$) 11.8 88.9 15.4 79.3 23.2 40.3 34.9 24.8
Salt & Pepper Probability=5% 12.0 85.4 13.3 81.7 26.3 28.8 31.8 29.6
Probability=10% 14.3 82.2 16.0 76.2 28.6 26.3 35.3 21.4
Table 7: Robustness evaluation on depth estimation and surface normal prediction tasks with different types and different intensities of noise. During all experiments, the input images are normalized between -1 to 1.
Appendix C Additional Qualitative Results
To further validate the effectiveness and generalization of DPBridge, we provide qualitative results for various dense prediction tasks, including semantic segmentation, optical flow prediction, and style transfer, by framing them as continuous-valued image-to-image translation problems. For semantic segmentation, we convert discrete labels into color maps and use our bridge model to predict colors in RGB space instead of categorical labels. For optical flow, the bridge process is defined between the flow image and the first image in the input RGB pair. To accommodate this, we adapt the UNet architecture to incorporate both RGB images as conditions during inference and convert the flow map into RGB space, predicting corresponding colors rather than uv coordinates. The following figures (Fig.5, Fig.6, Fig.7, Fig.8) illustrate that DPBridge achieves visually satisfactory results across different tasks. Note that only qualitative assessments are presented here; detailed implementations and quantitative evaluations are left as directions for future research.
Figure 5: Qualitative examples of applying DPBridge to semantic segmentation task on CityScape [14] dataset. We transform discrete labels into color maps and us our bridge model to predict colors in RGB space instead of categorical labels.
Figure 6: Qualitative examples of applying DPBridge to semantic segmentation task on CelebAMask-HQ [35] dataset. We transform discrete labels into color maps and us our bridge model to predict colors in RGB space instead of categorical labels.
Figure 7: Qualitative examples of applying DPBridge to optical flow prediction task on MPI-Sintel [7] dataset. The input image is the overlayed RGB pair, and we transform the flow map into RGB space so that the diffusion bridge process can be constructed accordingly.
Figure 8: Qualitative examples of applying DPBridge to style transfer task on Face-to-Comics [55] dataset.

### v2

@@
 Abstract
-Diffusion models have demonstrated remarkable success in dense prediction problems, which aims to model per-pixel relationship between RGB images and dense signal maps, thanks to their ability to effectively capture complex data distributions.
-However, initiating the reverse sampling trajectory from uninformative noise prior introduces limitations such as degraded performance and slow inference speed.
-In this work, we propose DPBridge, a generative framework that formulates dense prediction tasks as image-conditioned generation problems and establishes a direct mapping between input image and its corresponding dense map based on fully-tractable diffusion bridge process.
-This approach addresses aforementioned limitations in conventional diffusion-based solutions.
-In addition, we introduce finetuning strategies to adapt our model from pretrained image diffusion backbone, leveraging its rich visual prior knowledge to facilitate both efficient training and robust generalization ability.
-Experimental results shows that our DPBridge can achieve competitive performance compared to both feed-forward and diffusion-based approaches across various benchmarks, highlighting its effectiveness and adaptability.
-Figure 1: We present DPBridge to tackle general dense prediction tasks by establishing a stochastic transport between an RGB image and its corresponding dense map in the latent space based on a tractable diffusion bridge process.
+Diffusion models have shown remarkable capabilities in modeling complex data distributions by transforming noise into structured data through stochastic processes.
+However, when applied to dense prediction tasks whose goal is to capture per-pixel relationships between RGB images and dense signal maps, starting the sampling process from an uninformative Gaussian noise often leads to inefficient sampling and long latency.
+To overcome these challenges, we propose DPBridge, a generative framework that establishes direct mapping between input RGB images and dense signal maps based on a tractable bridge process.
+Furthermore, we introduce finetuning strategies to leverage a pretrained large-scale image diffusion backbone, enjoying its rich visual prior knowledge to enable both efficient training and robust generalization.
+Experiments show that DPBridge achieves competitive performance compared to both feed-forward and diffusion-based approaches across various benchmarks, validating its effectiveness and adaptability.
+Machine Learning, ICML
+Figure 1: We present DPBridge to tackle general dense prediction tasks by establishing conditional probability transport between RGB images and corresponding dense maps in the latent space based on a tractable diffusion bridge process.
 In addition, we introduce finetuning strategies to adapt large pretrained diffusion backbone to construct our bridge, leveraging its rich prior knowledge to enable both efficient training and robust generalization.
 1 Introduction
-Diffusion models recently emerge as a family of powerful generative models that are able to capture complex data distributions through a sequence of forward and reverse processes [26, 46].
-They accelerate the development of modern generative systems, and have consistently obtained state-of-the-art performance across various domains, such as image synthesis and video generation [17, 46, 27], controlled editing [61, 5, 50], human pose estimation [29, 30, 28], inverse problem solving [11, 53, 40].
-Despite such success, diffusion models still face limitations that hinder broader applications, especially in discriminative tasks like dense prediction.
-One major limitation stems from the fact that standard diffusion reverse process always starts from Gaussian noise, which contains little information about the target distribution.
-This will lead to degraded performance or prolonged convergence times.
-Current diffusion-based approaches formulate this type of tasks as image-conditioned generation problems and adopt measures like incorporating external conditions to the model [47, 61, 25], modifying the generation pipeline [41, 54], or adding extra guidance terms during sampling [11, 53] to deal with this issue.
-In this way they hope to better model the conditional generation process.
-Instead of starting from noise, an feasible alternative is to leverage the inherent problem structure.
-Since the goal of dense prediction tasks is to model the probability transport between image and dense signal map distributions, it’s a rather straightforward idea to directly start the generative process from input RGB images, as they are structurally correlated to target signal maps and can naturally serve as more informative prior compared to noise.
-This motivates the exploration of a new family of computational frameworks, i.e.
-diffusion bridge models [39, 63, 36, 37].
-They are based on a special type of diffusion process conditioned on given endpoints, thus enabling optimal transport between two distributions and generalizing beyond Gaussian priors.
-Inspired by these intuition, we propose Dense Prediction Bridge (DPBridge), a latent diffusion bridge model adapted from pretrained image diffusion backbone to handle general dense prediction tasks though the image-conditioned generation pipeline.
-In essence, we construct diffusion bridge process between the latent distributions of images and dense signal maps, with all intermediate samples along the trajectory being convex combinations of the paired data.
-The generative process thus becomes a series of transformations that iteratively synthesize desired signal maps starting from input images rather than pure noise, enabling faster inference speed as well as better generation quality.
-In addition, to better adapt pretrained diffusion backbone to our DPBridge construction, we also propose novel finetuning strategies to allow for more effective knowledge transfer.
-In this way our model is augmented with rich visual prior knowledge, which not only helps boosting the perception performance as well as generalization capability, but also significantly speeds up training convergence.
-Fig.(1) is a high-level overview of DPBridge, while also demonstrating its effectiveness on three representative dense prediction tasks, i.e.
+Dense prediction refers to the type of problems whose primary goal is to assign a value at every spatial location in the input data, generating output signal maps that possess same structure and size as inputs.
+It incorporates various typical tasks, including depth estimation (Yang et al., 2024; Ranftl et al., 2021; Eftekhar et al., 2021), surface normal prediction (Bae & Davison, 2024; Hu et al., 2024; Bae et al., 2021), semantic segmentation (Chen, 2017; Cheng et al., 2022; Wang et al., 2023; Weber et al., 2021), optical flow (Teed & Deng, 2020; Huang et al., 2022; Weinzaepfel et al., 2023), etc.
+Recently, image diffusion models have been applied to dense prediction tasks by formulating them as conditional generation problems.
+Their ability to capture complex data distributions makes them effective priors, leading to highly favorable results (Ke et al., 2024; Garcia et al., 2024; Fu et al., 2025).
+Following a common practice, these methods diffuse signal maps into pure noise and subsequently leverage input RGB images as conditions to guide the generation process.
+Nonetheless, there still exists limitations in diffusion-based solutions, mostly stem from the underlying modeling mechanism.
+The denoising process of a standard diffusion model always starts from Gaussian noise, which contains little information about the target signal distribution.
+In such cases, the distance between prior Gaussian distribution and target distribution remains large, requiring inevitably more sampling steps as well as ad-hoc conditioning mechanisms, at the risk of degraded performance and prolonged convergence time.
+Instead of starting from noise, one feasible alternative is to directly initiate the generative process from input RGB images, as they are structurally correlated to target signal maps and can naturally serve as more informative prior.
+This leads to the exploration of a new family of computational frameworks, i.e.
+diffusion bridge models (Liu et al., 2022; Zhou et al., 2023; Li et al., 2023; Liu et al., 2023).
+They are based on stochastic process with fixed starting and ending points and are adept at modeling conditional probability paths.
+By establishing direct connections between different distributions, this line of framework provides a more natural and efficient way to model probability transport compared with previous efforts by reducing the dependence on manually designed post-hoc guidance (Chung et al., 2022a; Song et al., 2023; Chung et al., 2022b) or customized generation pipeline (Meng et al., 2021).
+We therefore advocate that adopting diffusion bridge model will be a better fit to dense prediction tasks compared with conventional diffusion-based counterparts.
+Motivated by this, we propose Dense Prediction Bridge (DPBridge), a latent diffusion bridge model to handle general dense prediction tasks through generative pipeline.
+In essence, we modify the standard diffusion process to directly bridge the latents of RGB images and corresponding dense signal maps, and enforce the model to learn this process through maximum likelihood estimation.
+In this way the generative process thus becomes a series of transformations that iteratively synthesize desired signal maps starting from RGB images rather than pure noise, enabling faster inference speed as well as better generation quality.
+Furthermore, we introduce novel finetuning strategies to better adapt pretrained diffusion backbone to DPBridge construction.
+The equipped rich visual prior not only helps boosting the perception performance as well as generalization capability, but also significantly speeds up training convergence.
+Figure 1 is a high-level overview of DPBridge, while also demonstrating its effectiveness on three representative dense prediction tasks, i.e.
 depth estimation, surface normal prediction, and semantic segmentation, proving that our framework is theoretically universal and can be instantiated to different scenarios.
 To summarize, our contributions are as follows:
-  • We propose DPBridge framework to deal with dense prediction tasks with generative pipeline by formulating them as instances of latent diffusion bridges.
+  • We propose the DPBridge framework to address dense prediction tasks within a generative pipeline by formulating them as instances of latent diffusion bridges.
   • We propose finetuning strategies to effectively adapt large pretrained backbone to diffusion bridges, allowing for exploitation of visual prior knowledge and training efficiency improvements.
@@
 2.1 Dense Prediction
-Dense prediction refers to the type of problems whose outputs match the spatial dimensions of input data, often with the goal of assigning a value to each individual pixel.
-Typical examples include depth estimation, surface normal prediction, semantic segmentation, optical flow, etc.
-Common solutions can be categorized in discriminative and generative manner.
-Discriminative approaches focus on architecture design and introduce techniques like skip connections, dilated convolutions, multi-scale feature aggregation, and attention mechanisms [9, 44, 34] to better capture both local and global context for improved prediction accuracy.
-Generative methods, on the other hand, offer an alternative paradigm to tackle these problems through distribution modeling.
-Some works reuse or merge prior features from pretrained vision foundation models [4, 23, 62].
-Others formulate these tasks as visual-condition guided generation process which refines the desired dense map iteratively [18, 32, 24].
-2.2 Diffusion Bridge Models
-Stochastic bridge models are a common tool in probability theory which describe the evolution of stochastic processes with given starting/ending points.
-When combined with diffusion models, they provide a new perspective to translate between different distributions and model conditional probabilities without the need for prior knowledge, therefore have been gaining increasing attention in the context of generative modeling in recent year [39, 38, 16].
-They demonstrate the advantages of data-to-data generation paradigm over noise-to-data one on various tasks, including image-to-image translation [63, 37, 36], image restoration [59, 13], text-to-speech synthesis [10], etc.
-However, to the best of our knowledge, no prior work has explored leveraging diffusion bridge models on dense prediction tasks to facilitate the distribution transport between images and dense signal maps, despite their intuitive structural closeness.
-Our work aims to fill in this research gap.
+Dense predictions can be characterized as a conditional distribution modeling task and common solutions can be categorized into discriminative and generative ones.
+Discriminative methods directly map input data to structured outputs in a feed-forward manner.
+(Ranftl et al., 2021) introduces attention mechanism and transformer architecture to better capture both local and global context feature.
+(Yang et al., 2024) explore the usage of large-scale synthetic and pseudo-labeled images to enhance model’s capacity and robustness.
+(Bae & Davison, 2024) incorporates per-pixel ray directions and models the relative rotations between neighboring pixels.
+Recently, diffusion-based generative models have been applied in this field and showed remarkable performance.
+(Ke et al., 2024) finetunes pretrained diffusion model, leveraging the rich visual knowledge within to enhance depth estimation.
+(Fu et al., 2025) takes advantage of diffusion priors to jointly estimate depth and surface normals from single image, enhancing detail preservation by decoupling complex data distributions into distinct sub-distributions.
+(Gui et al., 2024) exploits flow matching to facilitate the distribution shift between images and depth maps.
+Our work follow conventional generative pipeline, but instead of starting from pure noise, we establish direct connections between input image and target signal map distributions, and model conditional probability transport in iterative denoising manner.
+2.2 Diffusion and Bridge Models
+Diffusion models (Ho et al., 2020; Song et al., 2020b) are a family of generative models that are able to capture complex data distributions through a sequence of forward and reverse processes.
+They accelerate the development of modern generative systems, and have consistently obtained state-of-the-art performance across various domains, such as image and video synthesis (Rombach et al., 2022; Ho et al., 2022), controlled editing (Zhang et al., 2023; Batzolis et al., 2021), inverse problem solving (Chung et al., 2022a; Song et al., 2023), 3D assets generation (Xiang et al., 2024; Zhang et al., 2024), and human-centered applications (Ji et al., 2024b, a; Ji & Li, 2024).
+Diffusion bridge models have gained increasing attention in generative modeling in recent years (Liu et al., 2022; Zhou et al., 2023; Liu & Wu, 2023; De Bortoli et al., 2021).
+They offer a novel perspective for translating between different distributions and enable a more effective data-to-data generation paradigm.
+Recent works have explored different types of bridge processes and their applications.
+(Liu et al., 2023; Li et al., 2023; Yue et al., 2023) develop Schrödinger bridge, Brownian bridge, and Ornstein-Uhlenbeck bridge frameworks, applying them to image restoration tasks such as inpainting, super-resolution.
+(Chung et al., 2024) proposes a modified inference procedure that enforces data consistency.
+(Chen et al., 2023) introduces bridge models to text-to-speech systems, enhancing synthesis quality and sampling efficiency.
+To the best of our knowledge, no prior work has explored the use of bridge models for dense prediction tasks to facilitate distribution transport between images and dense signal maps, despite their structural closeness.
+Our work aims to address this research gap.
 3 Preliminary
@@
 The transition kernel is specified as $q(\mathbf{z}_{t}|\mathbf{z}_{0})=\mathcal{N}(\alpha_{t}\mathbf{z}_{0},\sigma_{t}^{2}\mathbf{I})$ and the noise schedule comprising differentiable functions $\alpha_{t},\sigma_{t}$ depends on the design choice of drift and diffusion coefficients.
-It is also common to assess the diffusion process in terms of the signal-to-noise ratio (SNR), defined as $\text{SNR}_{t}=\alpha_{t}^{2}/\sigma_{t}^{2}$
+It is also common to assess the diffusion process in terms of the signal-to-noise ratio (SNR), defined as $\text{SNR}_{t}=\alpha_{t}^{2}/\sigma_{t}^{2}$.
 3.2 Diffusion Bridge Process
-Diffusion bridge process is a special type of diffusion process that guarantee to arrive at a prescribed final state $T$.
+Diffusion bridge process is a special type of diffusion process that is conditioned to end at a specified terminal point at timestep $T$.
 One natural way of its construction is to derive the conditioned process from a reference diffusion process given terminal constraints.
-We can use Doob’s h-transform technique [42, 48] to modify a reference process $\mathbb{Q}$ defined in Eqn.(1) to ensure that the event $\mathbf{z}_{T}=\mathbf{x}$ happens.
-Therefore the conditioned process $\mathbb{Q}^{\mathbf{x}}(\cdot):=\mathbb{Q}\left(\cdot|\mathbf{z}_{T}=\mathbf{x}\right)$ follows the law of
+We can use Doob’s h-transform technique (Oksendal, 2013; Särkkä & Solin, 2019) to modify a reference process $\mathbb{Q}$ defined in Equation 1 to ensure that the event $\mathbf{z}_{T}=\mathbf{x}$ happens, where $\mathbf{x}$ is a predefined value.
+The conditioned process $\mathbb{Q}^{\mathbf{x}}(\cdot):=\mathbb{Q}\left(\cdot|\mathbf{z}_{T}=\mathbf{x}\right)$ thus follows the law of
 $\mathrm{d}\mathbf{z}_{t}=\left[f(\mathbf{z}_{t},t)+g^{2}(t)h(\mathbf{z}_{t},t,\mathbf{x},T)\right]\mathrm{d}t+g(t)\mathrm{d}\mathbf{w}_{t}$ (2)
 where $h(\mathbf{z}_{t},t,\mathbf{x},T)=\nabla_{\mathbf{z}_{t}}\log q(\mathbf{z}_{T}=\mathbf{x}|\mathbf{z}_{t})$.
-Intuitively, the additional drift term $h(\mathbf{z}_{t},t,\mathbf{x},T)$ plays the role of steering $\mathbf{z}_{t}$ towards the specified target point $\mathbf{z}_{T}=\mathbf{x}$, and it is ensured that $\mathbb{Q}^{\mathbf{x}}$ invariably passes $\mathbf{x}$ at timestep $T$ with probability 1 for any initial state $\mathbf{z}_{0}$.
+Intuitively, the additional drift term $h(\mathbf{z}_{t},t,\mathbf{x},T)$ plays the role of steering $\mathbf{z}_{t}$ towards the target point $\mathbf{z}_{T}=\mathbf{x}$, and it is ensured that $\mathbb{Q}^{\mathbf{x}}$ invariably passes $\mathbf{x}$ at timestep $T$ with probability 1 for any initial point $\mathbf{z}_{0}$.
 The transition kernel of $\mathbb{Q}^{\mathbf{x}}$, given by $q(\mathbf{z}_{t}|\mathbf{z}_{0},\mathbf{z}_{T})$, can be subsequently derived through Kolmogorov equation:
@@
 4.1 Problem Formulation
-We formulate different dense prediction tasks like depth estimation, surface normal prediction, semantic segmentation etc., as image-conditioned generation problems and propose DPBridge to model the conditional distributions $p(\mathbf{y}|\mathbf{x})$, where $\mathbf{y}\in\mathbb{R}^{H\times W\times C}$ indicates the desired dense signal map (e.g.
-segmentation, depth, surface normal), and $\mathbf{x}\in\mathbb{R}^{H\times W\times 3}$ is an RGB image.
-In the following sections, we’ll first introduce the motivation of utilizing diffusion bridge models to handle dense prediction problems.
-Then, we’ll elaborate the details of our proposed DPBridge framework, which is constructed based on pretrained diffusion backbone through proper finetuning.
-Figure 2: Overview of our training pipeline, which is conducted in the latent space based on diffusion bridge theory.
+We formulate different dense prediction tasks as instances of image-conditioned generation problems and propose DPBridge to model the conditional distributions $p(\mathbf{y}|\mathbf{x})$, where $\mathbf{y}\in\mathbb{R}^{H\times W\times C}$ indicates the desired dense signal map (e.g.
+depth, surface normal, segmentation map), and $\mathbf{x}\in\mathbb{R}^{H\times W\times 3}$ is the input RGB image.
+In the following sections, we’ll elaborate the details of DPBridge framework, which is constructed based on pretrained diffusion backbone through proper finetuning.
+Figure 2: Overview of our training pipeline.
 To overcome the discrepancy with standard diffusion process and facilitate finetuning, we derive the maximum likelihood training objective of diffusion bridge process, propose a distribution aligned normalization strategy, and assist the training with an auxiliary image consistency loss.
-4.2 Motivation: Bridge VS Diffusion
-As introduced previously, the incompatibility that employ diffusion models to dense prediction task, which is essentially a image-conditioned generation problem, lies in its underlying iterative denoising process starting from Gaussian noise.
-In such cases, the distance between prior Gaussian distribution and data distribution remains large, requiring inevitably more sampling steps as well as ad-hoc conditioning mechanisms to converge to target distributions, at the risk of degraded performance and slow inference speed.
-Bridge models, on the other hand, directly formulate generative processes as conditional probability paths.
-In this scenario, the boundary distribution is not restricted to Gaussian noise, but instead can be chosen as more informative priors, such as the condition data distribution.
-By establishing direct connections between different distributions, this formulation provides a more natural and efficient way to model probability transport compared with previous efforts.
-For example, it reduce the dependence on manually designed post-hoc guidance [11, 53, 12] or customized generation pipeline [41].
-In addition, through implicitly integrates condition as part of the learned process, it provides a more flexible way to model intricacies between data and conditions without relying on explicit mechanisms like concatenation, attention [47, 61, 25], whose inductive bias may harm the overall performance when dealing with complex circumstances.
-Considering these advantages, we advocate that adopting diffusion bridge model will be a better fit to dense prediction tasks compared with conventional diffusion-based counterparts.
 4.2 Dense Prediction Bridge
-Given paired training data, a naive implementation is to construct the diffusion bridge model from scratch according to Sec.3.2.
+Given paired training data, a naive implementation is to construct the diffusion bridge model from scratch according to Section 3.2.
 However, this approach is extremely resource-demanding, making it hard to implement with limited computation power.
-What’s more severe, it lacks the ability to harness rich scene understanding knowledge inherent in the vision foundation models, which have been proven significant in terms of boosting both perception quality and generalization performance in dense prediction tasks.
-Therefore, our DPBridge is build upon pretrained diffusion backbones through finetuning and we explore several strategies to handle the misalignment between original diffusion bridge process and standard diffusion process, aiming for more efficient knowledge transfer.
-Fig.2 gives an overview of the proposed finetuning pipeline, and we’ll explain the design of each component next.
+What’s more, it lacks the ability to harness rich visual prior knowledge inherent in the foundation models, which have been proven significant in terms of boosting both perception quality and generalization performance in dense prediction tasks.
+Therefore, our DPBridge is build upon pretrained diffusion backbones through finetuning and we explore several strategies to handle the misalignment between bridge process and standard diffusion process, aiming for more efficient knowledge transfer.
+Figure 2 gives an overview of the proposed finetuning pipeline, and next we’ll explain the design of each component.
 4.3 Latent Space Implementation
-Instead of directly learning in image pixel space, we follow common practices to perform diffusion steps in a compressed but perceptually aligned latent space, provided by an independently trained variational autoencoder (VAE).
-Specifically, both the target dense maps and RGB images are converted by the encoder to obtain $\mathbf{z}_{0}=\mathcal{E}(\mathbf{y})$ and $\mathbf{z}_{T}=\mathcal{E}(\mathbf{x})$, serving as the boundary states.
+Instead of directly learning in image pixel space, we follow common practice to perform diffusion steps in a compressed but perceptually aligned latent space, provided by an independently trained variational autoencoder (VAE).
+Specifically, both the dense signal maps and RGB images are converted by the encoder to obtain $\mathbf{z}_{0}=\mathcal{E}(\mathbf{y})$ and $\mathbf{z}_{T}=\mathcal{E}(\mathbf{x})$, serving as the initial and terminal points.
 In this way the diffusion bridge process is henceforth learned in the latent space.
@@
 4.4.1 Maximum Likelihood Loss
-In previous diffusion bridge models, widely-used techniques for learning the reverse process that corresponds to Eqn.(2) is to follow the theory of [1], use a neural network to approximate the score function and apply score matching technique.
+Widely-used techniques for learning the reverse process that corresponds to Equation 2 is to follow the theory of (Anderson, 1982), use a neural network to approximate the score function and apply score matching.
 We notice that analytical form of diffusion bridge’s reverse transition kernel is directly obtainable, therefore the usage of maximum likelihood training can also be applied, which is performed by optimizing the Evidence Lower Bound (ELBO).
 It yields a more stable loss function compared with score matching and better aligns with the pretrained backbone.
-Given the fixed terminal state $\mathbf{z}_{T}$, the expectation of data log-likelihood possesses the following ELBO formulation:
-$\text{ELBO}=&\mathbb{E}_{p(\mathbf{z}_{0})}\Big{[}\mathbb{E}_{p(\mathbf{z}_{1}\mid\mathbf{z}_{0})}\big{[}\log p_{\theta}(\mathbf{z}_{0}\mid\mathbf{z}_{1},\mathbf{z}_{T})\big{]}-\\ \sum_{t=2}^{T}\mathbb{E}_{p(\mathbf{z}_{t}|\mathbf{z}_{0})}&\big{[}\mathbb{D}_{KL}\big{(}q(\mathbf{z}_{t-1}\mid\mathbf{z}_{0},\mathbf{z}_{t},\mathbf{z}_{T})\,||\,p_{\theta}(\mathbf{z}_{t-1}\mid\mathbf{z}_{t},\mathbf{z}_{T})\big{)}\big{]}\Big{]}\end{aligned}$
+Given the specified terminal state $\mathbf{z}_{T}$, the expectation of data log-likelihood possesses the following ELBO formulation:
+[…]
 where $q(\mathbf{z}_{t-1}|\mathbf{z}_{t},\mathbf{z}_{0},\mathbf{z}_{T})$ and $p_{\theta}(\mathbf{z}_{t-1}|\mathbf{z}_{t},\mathbf{z}_{T})$ are the transition kernel of true reverse process and our modeled generative process and $\mathbb{D}_{KL}$ is the KL divergence.
-Utilize Bayes’ theorem and Markov chain property, we obtain:
+Utilizing Bayes’ theorem and Markov chain property, we can obtain:
 $q(\mathbf{z}_{t-1}|\mathbf{z}_{t},\mathbf{z}_{0},\mathbf{z}_{T})$ $=\frac{q(\mathbf{z}_{t}|\mathbf{z}_{t-1},\mathbf{z}_{T})q(\mathbf{z}_{t-1}|\mathbf{z}_{0},\mathbf{z}_{T})}{q(\mathbf{z}_{t}|\mathbf{z}_{0},\mathbf{z}_{T})}$ (5)
@@
 $\mathcal{L}_{elbo}=\mathbb{E}_{t,\mathbf{z}_{0},\mathbf{z}_{t},\mathbf{z}_{T}}\left[\frac{1}{2\tilde{\sigma}_{t}^{2}}\left\|\tilde{\boldsymbol{\mu}}_{t}-\boldsymbol{\mu}_{\theta}(\mathbf{z}_{t},\mathbf{z}_{T},t)\right\|^{2}\right]$ (8)
-Similar to the strategy in DDPM [26], we use a network $\boldsymbol{\epsilon}_{\theta}(\mathbf{z}_{t},\mathbf{z}_{T},t)$ to parameterize noise instead of posterior mean and further simplified the loss as:
+Similar to the strategy in DDPM (Ho et al., 2020), we use a network $\boldsymbol{\epsilon}_{\theta}(\mathbf{z}_{t},\mathbf{z}_{T},t)$ to parameterize noise instead of posterior mean and further simplified the loss as:
 $\mathcal{L}_{elbo}=\mathbb{E}_{t,\mathbf{z}_{0},\mathbf{z}_{t},\mathbf{z}_{T}}\left[\left\|\boldsymbol{\epsilon}-\boldsymbol{\epsilon}_{\theta}(\mathbf{z}_{t},\mathbf{z}_{T},t)\right\|^{2}\right],$ (9)
 4.4.2 Image Consistency Loss
-In addition to the common likelihood loss evaluated in the latent space, we also incorporate an image-level consistency loss, where we first convert the latent prediction back to image space, and then evaluate its consistency with the ground truth dense map using a simple MSE loss.
+In addition to the likelihood loss evaluated in latent space, we also incorporate an image-level consistency loss, where we first convert the latent prediction back to image space, and then evaluate its consistency with the ground truth dense map using a simple MSE loss.
 $\mathcal{L}_{ic}=\left\|\mathcal{D}(\hat{\mathbf{z}}_{0})-\mathbf{y})\right\|^{2}$ (10)
@@
 $\mathcal{L}=\omega_{1}\mathcal{L}_{elbo}+\omega_{2}\mathcal{L}_{ic}$ (11)
-where $\omega_{1},\omega_{2}$ are adjustable hyperparameters to balance the loss. An overview of DPBridge’s training pipeline shown in Algorithm.1.
+where $\omega_{1},\omega_{2}$ are hyperparameters to balance the loss.
 Algorithm 1 Training
-1:Encoder $\mathcal{E}$; Decoder $\mathcal{D}$; Loss weights $\omega_{1},\omega_{2}$
-2:repeat
-3:     Retrieve paired data $(\mathbf{x},\mathbf{y})$
-4:     $\mathbf{z}_{0}=\mathcal{E}(\mathbf{y})$, $\mathbf{z}_{T}=\mathcal{E}(\mathbf{x})$
-5:     $t\sim\text{Uniform}(\{1,\dots,T\})$, $\boldsymbol{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$
-6:     $\mathbf{z}_{t}=n_{t}\mathbf{z}_{T}+m_{t}\mathbf{z}_{0}+\sigma_{t}\boldsymbol{\epsilon}$
-7:     $\mathbf{z}_{t}^{\prime}=\frac{1}{\sqrt{n_{t}^{2}+\sigma_{t}^{2}}}(\mathbf{z}_{t}-n_{t}\mathbf{z}_{T})$
-8:     $\mathcal{L}_{elbo}=\left\|\boldsymbol{\epsilon}-\boldsymbol{\epsilon}_{\theta}(\mathbf{z}_{t}^{\prime},\mathbf{z}_{T},t)\right\|^{2},$
-9:     $\hat{\mathbf{z}}_{0}=\frac{1}{m_{t}}(\mathbf{z}_{t}-n_{t}\mathbf{z}_{T}-\sigma_{t}\boldsymbol{\epsilon}_{\theta}(\mathbf{z}_{t}^{\prime},\mathbf{z}_{T},t))$
-10:     $\mathcal{L}_{ic}=\left\|\mathcal{D}(\hat{\mathbf{z}}_{0})-\mathbf{y}\right\|^{2}$
-11:     Take gradient step on $\nabla_{\theta}\left(\omega_{1}\mathcal{L}_{elbo}+\omega_{2}\mathcal{L}_{ic}\right)$
-12:until converged
+0:  Encoder $\mathcal{E}$; Decoder $\mathcal{D}$; Loss weights $\omega_{1},\omega_{2}$
+1:  repeat
+2:     Retrieve paired data $(\mathbf{x},\mathbf{y})$
+3:     $\mathbf{z}_{0}=\mathcal{E}(\mathbf{y})$, $\mathbf{z}_{T}=\mathcal{E}(\mathbf{x})$
+4:     $t\sim\text{Uniform}(\{1,\dots,T\})$, $\boldsymbol{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$
+5:     $\mathbf{z}_{t}=m_{t}\mathbf{z}_{0}+n_{t}\mathbf{z}_{T}+\sigma_{t}\boldsymbol{\epsilon}$
+6:     $\mathbf{z}_{t}^{\prime}=\frac{1}{\sqrt{n_{t}^{2}+\sigma_{t}^{2}}}(\mathbf{z}_{t}-n_{t}\mathbf{z}_{T})$
+7:     $\mathcal{L}_{elbo}=\left\|\boldsymbol{\epsilon}-\boldsymbol{\epsilon}_{\theta}(\mathbf{z}_{t}^{\prime},\mathbf{z}_{T},t)\right\|^{2},$
+8:     $\hat{\mathbf{z}}_{0}=\frac{1}{m_{t}}(\mathbf{z}_{t}-n_{t}\mathbf{z}_{T}-\sigma_{t}\boldsymbol{\epsilon}_{\theta}(\mathbf{z}_{t}^{\prime},\mathbf{z}_{T},t))$
+9:     $\mathcal{L}_{ic}=\left\|\mathcal{D}(\hat{\mathbf{z}}_{0})-\mathbf{y}\right\|^{2}$
+10:     Take gradient step on $\nabla_{\theta}\left(\omega_{1}\mathcal{L}_{elbo}+\omega_{2}\mathcal{L}_{ic}\right)$
+11:  until Converged
 Algorithm 2 Inference
-1:Encoder $\mathcal{E}$; Decoder $\mathcal{D}$; Trained $\boldsymbol{\epsilon}_{\theta}(\cdot,\cdot,\cdot)$
-2:$\mathbf{z}_{T}=\mathcal{E}(\mathbf{x})$
-3:for $t=T,\dots,1$ do
-4:     $\boldsymbol{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ if $t>1$, else $\boldsymbol{\epsilon}=\mathbf{0}$
-5:     $\mathbf{z}_{t}^{\prime}=\frac{1}{\sqrt{n_{t}^{2}+\sigma_{t}^{2}}}(\mathbf{z}_{t}-n_{t}\mathbf{z}_{T})$
-6:     $\hat{\mathbf{z}}_{0}=\frac{1}{m_{t}}(\mathbf{z}_{t}-n_{t}\mathbf{z}_{T}-\sigma_{t}\boldsymbol{\epsilon}_{\theta}(\mathbf{z}_{t}^{\prime},\mathbf{z}_{T},t))$
-7:     $\mathbf{z}_{t-1}=k_{1}\mathbf{z}_{t}+k_{2}\hat{\mathbf{z}}_{0}+k_{3}\mathbf{z}_{T}+\tilde{\sigma}_{t}\boldsymbol{\epsilon}$
-8:end for
-9:$\mathbf{y}=\mathcal{D}(\hat{\mathbf{z}}_{0})$
-10:return $\mathbf{y}$
+0:  Encoder $\mathcal{E}$; Decoder $\mathcal{D}$; Trained $\boldsymbol{\epsilon}_{\theta}(\cdot,\cdot,\cdot)$
+1:  $\mathbf{z}_{T}=\mathcal{E}(\mathbf{x})$
+2:  for $t=T,\dots,1$ do
+3:     $\boldsymbol{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ if $t>1$, else $\boldsymbol{\epsilon}=\mathbf{0}$
+4:     $\mathbf{z}_{t}^{\prime}=\frac{1}{\sqrt{n_{t}^{2}+\sigma_{t}^{2}}}(\mathbf{z}_{t}-n_{t}\mathbf{z}_{T})$
+5:     $\hat{\mathbf{z}}_{0}=\frac{1}{m_{t}}(\mathbf{z}_{t}-n_{t}\mathbf{z}_{T}-\sigma_{t}\boldsymbol{\epsilon}_{\theta}(\mathbf{z}_{t}^{\prime},\mathbf{z}_{T},t))$
+6:     $\mathbf{z}_{t-1}=k_{1}\mathbf{z}_{t}+k_{2}\hat{\mathbf{z}}_{0}+k_{3}\mathbf{z}_{T}+\tilde{\sigma}_{t}\boldsymbol{\epsilon}$
+7:  end for
+8:  $\mathbf{y}=\mathcal{D}(\hat{\mathbf{z}}_{0})$
+9:  return $\mathbf{y}$
 4.5 Distribution Alignment Normalization
@@
 To address this issue, we propose a distribution alignment normalization technique to fill in the gap between the two distributions.
-Specifically, we normalize the intermediate latents computed from Eqn.(3) to match the form of a standard diffusion process:
+Specifically, we normalize the intermediate latents computed from Equation 3 to match the form of a standard diffusion process:
 $\mathbf{z}_{t}^{\prime}=\frac{\mathbf{z}_{t}-m_{t}\mathbf{z}_{T}}{\sqrt{n_{t}^{2}+\sigma_{t}^{2}}}=\frac{n_{t}}{\sqrt{n_{t}^{2}+\sigma_{t}^{2}}}\mathbf{z}_{0}+\frac{\sigma_{t}}{\sqrt{n_{t}^{2}+\sigma_{t}^{2}}}\boldsymbol{\epsilon}_{t}$ (12)
@@
 Therefore by feeding the normalized sample $\mathbf{z}_{t}^{\prime}$ in the network $\boldsymbol{\epsilon}_{\theta}$, we can fully exploit the denoising capability from pretrained diffusion backbone.
-4.7 Network Architecture
-To better utilize visual information ($\mathbf{z}_{T}$ in our case), we concatenating both intermediate and terminal state samples along the feature dimension and send in the noise prediction network as a whole.
-To address the dimension incompatibility with pretrained backbone, the input channels of the network are thus doubled to accommodate the concatenated input.
-In addition, as suggested by [32], we duplicate the weight tensor of these accommodated layers and divide the values by two to preserve the input magnitude to subsequent layers.
-As DPBridge is an image-conditioned generative framework, we employ null-text embedding for all cross-attention layers to eliminate the impact of text prompts.
+An overview of DPBridge’s training pipeline shown in Algorithm 1.
 4.6 Accelerate Inference
-Similar to the idea of DDIM [52], the inference process of DPBridge can be accelerated by utilizing a non-Markovian process while maintaining the same marginal distributions for all intermediate samples along the sampling path.
+Similar to the idea of DDIM (Song et al., 2020a), the inference process of DPBridge can be accelerated by utilizing a non-Markovian process while maintaining the same marginal distributions for all intermediate samples along the sampling path.
 Specifically, we can sample on a subset of timesteps $\{t_{i}\}_{i=1}^{N}$, where $0=t_{0}<t_{1}<\cdots<t_{N-1}<t_{N}=T$ and the number of sampling steps $N$ can be set independently from the original timesteps $T$ on which the bridge model is trained.
 The accelerated inference process can be thus described as:
-$\begin{split}\mathbf{z}_{t_{n}}&=n_{t_{n}}\mathbf{z}_{T}+m_{t_{n}}\hat{\mathbf{z}}_{0}+\\ &\sqrt{\sigma_{t_{n}}^{2}-g_{n}^{2}}\frac{\mathbf{z}_{t_{n+1}}-n_{t_{n+1}}\mathbf{z}_{T}-m_{t_{n+1}}\hat{\mathbf{z}}_{0}}{\sigma_{t_{n+1}}}+g_{n}\boldsymbol{\epsilon}\end{split}$ (13)
+$\begin{split}\mathbf{z}_{t_{n}}&=m_{t_{n}}\hat{\mathbf{z}}_{0}+n_{t_{n}}\mathbf{z}_{T}+\\ &\sqrt{\sigma_{t_{n}}^{2}-g_{n}^{2}}\frac{\mathbf{z}_{t_{n+1}}-m_{t_{n+1}}\hat{\mathbf{z}}_{0}-n_{t_{n+1}}\mathbf{z}_{T}}{\sigma_{t_{n+1}}}+g_{n}\boldsymbol{\epsilon}\end{split}$ (13)
 where $g_{n}$ is the variance parameter that controls the stochasticity behavior of the reverse sampling process.
-We note that when $g_{n}=\tilde{\sigma}_{t}^{2}$ which is calculated from Eqn.(5), the process becomes a Markovian bridge.
-The basic inference process of DPBridge is summarized in Algorithm.2.
-If adopt acceleration, one can utilize Eqn.(13) during implementation.
+We note that when $g_{n}=\tilde{\sigma}_{t}^{2}$ which is calculated from Equation 5, the process becomes a Markovian bridge.
+The basic inference process of DPBridge is summarized in Algorithm 2.
+If adopt acceleration, one can utilize Equation 13 during implementation.
 5 Experiments
-We evaluate the proposed DPBridge approach mainly on two representative dense prediction tasks, i.e., depth estimation and surface normal prediction.
-We’ll start by a brief introduction of experiment settings of these two tasks, followed by quantitative and qualitative analysis.
+We conduct evaluation mainly on two representative dense prediction tasks, i.e., depth estimation and surface normal prediction.
+We’ll start by a brief introduction of experiment settings as well as implementation details, followed by quantitative and qualitative analysis.
 Lastly, we conduct detailed ablation studies and investigate the influence each design component respectively.
@@
 The basic setting for depth estimation and surface normal prediction experiments are similar.
-We use two synthetic datasets, HyperSim [45] and Virtual KITTI [8], which cover both indoor and outdoor scenes, for training.
+We use two synthetic datasets, HyperSim (Roberts et al., 2021) and Virtual KITTI (Cabon et al., 2020), which cover both indoor and outdoor scenes, for training.
 To comprehensively evaluate the model performance and demonstrate its generalization capability, we conduct zero-shot evaluations on real-world datasets.
-For depth estimation, this includes NYUv2 [51], KITTI [22], ScanNet [15], ETH3D [49] and DIODE [56].
-For surface normal prediction, we additionally incorporate iBims [33] and Sintel [7] datasets in the evaluation pipeline.
-For fair comparison, we use the same validation split as in previous methods [32].
+For depth estimation, this includes NYUv2 (Silberman et al., 2012), KITTI (Geiger et al., 2012), ScanNet (Dai et al., 2017), ETH3D (Schops et al., 2017) and DIODE (Vasiljevic et al., 2019).
+For surface normal prediction, we additionally incorporate iBims (Koch et al., 2018) and Sintel (Butler et al., 2012) datasets.
+For fair comparison, we use the same validation split as in previous methods (Ke et al., 2024).
 Evaluation Metrics.
@@
 5.2 Implementation Details
-We implement DPBridge based on Stable Diffusion (SD) v1.5 [46] as backbone.
+We implement DPBridge based on Stable Diffusion v1.5 (SD1.5) and v2.1-base (SD2.1) (Rombach et al., 2022) and report their results respectively to shown compatibility with different backbones.
 During training, the number of timesteps of is set to be 1000, and we use 50 sampling steps during inference with considerations of both quality and efficiency.
@@
 We also apply random horizontal flipping augmentation during training.
-The whole process takes approximately 2 days to complete on a single Nvidia RTX 3090Ti GPU.
-Method Model Type NYUv2 KITTI ETH3D ScanNet DIODE
+The whole process takes approximately 3 days to complete on a single Nvidia RTX 3090Ti GPU.
+Method Model Type NYUv2 KITTI ETH3D ScanNet DIODE Avg.
 […]
-Marigold [32] 8.6 93.6 13.8 82.6 9.3 93.2 10.4 89.7 33.6 73.9
-Geowizard [20] 8.5 93.7 16.2 75.6 9.1 93.1 9.0 93.0 32.0 72.2
-E2E-FT [21] 8.2 94.0 15.4 77.9 10.6 89.8 9.2 92.3 32.9 73.7
-ControlNet [61] 45.5 54.6 55.2 40.2 30.5 58.8 25.7 68.3 52.1 50.1
-InstructPix2Pix [6] 35.2 60.0 47.7 51.2 26.1 66.7 21.4 75.3 44.8 58.0
-I2SB [37] 24.9 73.4 37.7 57.9 40.6 52.0 21.2 66.1 40.2 59.1
-BBDM [36] 27.9 65.5 47.2 45.1 46.4 48.8 22.7 67.0 34.6 64.7
-DPBridge (Ours) 6.9 95.7 11.2 87.3 8.1 93.2 7.5 94.7 29.7 78.6
+DiverseDepth (Yin et al., 2020) 11.7 87.5 19.0 70.4 22.8 69.4 10.9 88.2 37.6 63.1 12.4
+MiDaS (Ranftl et al., 2020) 11.1 88.5 23.6 63.0 18.4 75.2 12.1 84.6 33.2 71.5 11.8
+LeReS (Yin et al., 2021) 9.0 91.6 14.9 78.4 17.1 77.7 9.1 91.1 27.1 76.6 9.5
+Omnidata (Eftekhar et al., 2021) 7.4 94.5 14.9 83.5 16.6 77.8 7.5 93.2 33.9 74.2 9.2
+HDN (Zhang et al., 2022) 6.9 94.8 11.5 86.7 12.1 83.3 8.0 93.6 24.6 78.0 7.2
+DPT (Ranftl et al., 2021) 9.8 90.3 10.0 90.1 7.8 94.6 8.2 93.8 18.2 75.8 6.9
+Depth Anything V2 (Yang et al., 2024) 4.4 97.9 7.5 94.8 13.2 84.2 4.4 98.0 6.5 95.4 2.3
+Marigold (Ke et al., 2024) 6.0 95.9 10.5 90.4 7.1 95.1 6.9 94.5 31.0 71.2 5.8
+DepthFM (Gui et al., 2024) 6.5 95.6 8.3 93.4 - - - - 22.5 80.0 3.8
+Geowizard (Fu et al., 2025) 5.7 96.2 14.4 82.0 7.5 94.3 6.1 95.8 31.4 77.1 6.1
+E2E-FT (Garcia et al., 2024) 5.4 96.4 9.6 92.1 6.4 95.9 5.8 96.5 30.3 77.6 3.4
+I2SB (Liu et al., 2023) 24.9 73.4 37.7 57.9 40.6 52.0 21.2 66.1 40.2 59.1 13.9
+BBDM (Li et al., 2023) 27.9 65.5 47.2 45.1 46.4 48.8 22.7 67.0 34.6 64.7 14.1
+DPBridge (SD1.5) 6.9 95.7 11.2 87.3 8.1 93.2 7.5 94.7 29.7 78.6 6.1
+DPBridge (SD2.1) 4.4 97.6 9.8 91.6 6.9 94.9 4.5 97.4 23.8 84.0 2.6
 Table 1: Quantitative comparison of DPBridge with state-of-the-art affine-invariant depth estimation methods on zero-shot benchmarks.
-All metrics are presented in percentage terms with best and second-best results marked by bold and underline.
-DM and BM correspond to diffusion models and bridge models respectively.
-All models are trained using SDv1.5 as backbone.
+All metrics are presented in percentage terms.
+FF, DM, and BM correspond to feed-forward, diffusion models, and bridge models.
+The best and second-best results are highlighted in bold and underline formats.
+Method Model Type NYUv2 ScanNet iBims-1 Sintel Avg.
+[…]
+EESNU (Bae et al., 2021) 16.2 58.6 - - 20.0 58.5 42.1 11.5 6.5
+Omnidata v1 (Eftekhar et al., 2021) 23.1 45.8 22.9 47.4 19.0 62.1 41.5 11.4 8.1
+Omnidata v2 (Kar et al., 2022) 17.2 55.5 16.2 60.2 18.2 63.9 40.5 14.7 5.6
+DSINE (Bae & Davison, 2024) 16.4 59.6 16.2 61.0 17.1 67.4 34.9 21.5 3.2
+Marigold (Ke et al., 2024) 18.8 55.9 17.7 58.8 18.4 64.3 39.1 14.9 5.8
+Geowizard (Fu et al., 2025) 17.0 56.5 15.4 61.6 13.0 65.3 40.4 13.2 4.1
+E2E-FT (Garcia et al., 2024) 16.5 60.4 14.7 66.1 16.1 69.7 33.5 22.3 2.1
+DPBridge (SD1.5) 18.1 54.2 18.0 54.5 24.3 51.2 41.2 19.4 7.3
+DPBridge (SD2.1) 14.8 65.3 13.4 68.5 17.1 68.0 25.0 46.3 1.4
+Table 2: Quantitative comparison of DPBridge with state-of-the-art surface normal prediction methods on zero-shot benchmarks.
+All metrics are presented in percentage terms.
+FF, DM, and BM correspond to feed-forward, diffusion models, and bridge models.
+The best and second-best results are highlighted in bold and underline formats.
 5.3 Quantitative and Qualitative Results
-Tab.1 and Tab.2 summarize the quantitative metrics of DPBridge compared with diffusion-based and bridge-based counterparts, who are most closely related to our work.
-The reason we use SDv1.5 instead of more advanced SDv2.1 as backbone is that our current pipeline does not support v-prediction parameterization as in SDv2.1, and this incompatibility will undermine the model’s knowledge transfer ability.
-We’ll present more detail analysis in the ablation study.
-For a fair comparison, we re-implement all baseline methods using the same SDv1.5 backbone as we do, and report the results.
-We also conduct comparison with feed-forward and diffusion-based approaches based on SDv2.1 backbone, and the complete results is presented in the Appendix.
-DPBridge outperforms most methods in our evaluation setting and even be comparable with state-of-the-art that is trained with SDv2.1.
-Specifically, when compared with bridge models like I2SB and BBDM that are trained from scratch, we exceed their performance by a large margin, justifying our proposal that adapting prior knowledge to bridge process indeed benefits the perception performance.
+Table 1 and Table 2 summarize the quantitative metrics of DPBridge compared with state-of-the-art (SOTA) methods from both discriminative and generative fields, all in zero-shot manner.
+In depth estimation, DPBridge consistently matches SOTA results, while demonstrating robust generalization across datasets.
+It achieves 4.4 AbsRel, 97.2 % $\delta_{1}$ on NYUv2, and 4.5 AbsRel, 97.4% $\delta_{1}$ on ScanNet.
+Similarly, for surface normal prediction, DPBridge also delivers competitive results, achieving mean error of 14.8 and $11.25^{\circ}$ accuracy of $65.3\%$ on NYUv2, as well as maintaining strong performance on other datasets like iBims-1 (mean: 17.1, $11.25^{\circ}:68.0\%$) and Sintel (mean: 25.0, $11.25^{\circ}:46.3\%$).
+These results highlight DPBridge’s versatility to generalize well across tasks and benchmarks, being comparable to both feed-forward and diffusion-based methods while offering a unified, efficient paradigm to handle dense prediction tasks.
+Notably, when compared with bridge models like I2SB and BBDM that are trained from scratch, we exceed their performance by a large margin, justifying our proposal that adapting prior knowledge to bridge construction indeed benefits the perception performance.
 In terms of diffusion-based methods, DPBridge can also achieve the best result, showcasing that adopting bridge process to handle dense prediction tasks is indeed advantageous over standard diffusion framework.
-Figure 3: Qualitative comparison of monocular depth estimation methods on KITTI and ETH3D dataset.
+We further provide qualitative examples for both tasks in comparison with SOTA diffusion-based methods (Ke et al., 2024; Fu et al., 2025; Garcia et al., 2024), which are most closely related to our work.
+Figure 3 shows that E2E-FT tends to produce blurry and smooth predictions, while Geowizard fails to provide accurate estimation in relatively long distance range.
+In contrast, DPBridge perform well in both indoor and outdoor scenes, capturing fine-level details in both near and faraway regions, like the contour of leaves, fences, cars, computers or windows.
+Figure 4 presents visualizations on in-the-wild images for surface normal prediction task, where we can see that DPBridge is able to predict high-quality surface normals equipped with fine-grained textures.
+Figure 3: Qualitative comparison of diffusion-based depth estimation methods on KITTI and ETH3D dataset.
 DPBridge is able to generalize in both indoor and outdoor scenes, successively capturing fine-grained details such as leaves and structures on the fence (KITTI).
-Figure 4: Qualitative comparison between our (above) and Marigold trained on SDv1.5 (below) on normal estimation results, DPBridge is able to capture more textures from the photos in-the-wild.
-The qualitative results for depth estimation is conducted on KITTI and ETH3D datasets, as shown in Fig.
-3.
-We can see that E2E-FT tends to produce blurry and smooth predictions, while Geowizard fails to provide accurate estimation in relatively long distance range.
-In contrast, DPBridge perform well in both indoor and outdoor scenes, capturing fine-level details in both near and faraway regions, like the contour of leaves, fences, cars, computers or windows.
-We also present visualizations on in-the-wild images for surface normal prediction task.
-DPBridge is able to predict high-quality surface normals equipped with fine-grained textures.
-This strong generalization capability can be attribute to that we effectively harness extensive visual prior knowledge inherent in the pretrained diffusion backbone.
-Method NYUv2 ScanNet iBims-1 Sintel
-[…]
-Marigold [32] 22.3 40.0 23.3 37.6 28.0 38.5 52.1 8.1
-Geowizard [20] 21.7 44.2 22.0 42.1 28.2 37.9 45.8 11.0
-E2E-FT [21] 17.7 57.8 17.4 57.4 18.0 64.6 43.9 14.1
-DPBridge (Ours) 18.1 54.2 18.0 54.5 24.3 51.2 41.2 19.4
-Table 2: Quantitative comparison of DPBridge with state-of-the-art surface normal prediction methods on zero-shot benchmarks.
-All metrics are presented in percentage terms with best and second-best results marked by bold and underline.
-DM and BM correspond to diffusion models and bridge models respectively.
-All models are trained using SDv1.5 as backbone.
+Figure 4: Qualitative comparison between DPBridge (above) and Marigold (below) on normal estimation results, DPBridge is able to capture more textures from the photos in-the-wild.
 Design Components NYUv2 KITTI
 […]
-26.8 85.7 33.6 57.7
-8.8 93.0 14.0 83.1
-8.5 93.3 13.4 83.9
-7.5 94.9 13.1 84.6
-6.9 95.7 11.2 87.3
-Table 3: The ablation study regarding the design components of DPBridge.
-FT stands for utilizing pretrained diffusion backbone during training.
+21.8 83.7 30.6 62.7
+6.4 94.0 11.2 87.0
+6.2 95.8 11.0 87.5
+5.0 96.3 10.3 90.1
+4.4 97.6 9.8 91.6
+Table 3: Ablation studies regarding different design components.
+FT stands for finetuning pretrained backbone.
 DAN denotes distribution aligned normalization technique.
 IC indicates image consistency loss.
-Parameterization Backbone Steps NYUv2 KITTI
-[…]
-50 8.2 94.5 12.7 85.0
-50 23.7 59.1 30.5 56.7
-50 7.6 95.6 12.0 85.8
-50 7.7 94.6 12.4 85.2
-50 22.8 62.4 28.5 60.1
-5 12.1 86.5 19.4 66.5
-10 7.7 95.1 12.5 84.1
-50 6.9 95.7 11.2 87.3
-100 6.8 95.1 11.4 87.5
-200 6.6 95.6 12.3 84.8
-Table 4: The ablation study regarding the hyperparameter choice of diffusion bridge process.
-Options include different network parameterizations, the version of SD to use as backbone, and the number of denoising steps during inference.
+Figure 5: Ablation of sampling steps. The performance improves as the number of sampling steps increases, while we observe saturation after 10 steps.
 5.4 Ablation Studies
 To evaluate the effectiveness of each design component and training setting, we conduct ablation studies using depth estimation tasks on the NYUv2 and KITTI datasets.
-The results are summarized in Tab.3 and Tab.4, with detailed discussions provided below.
-5.4.1 Design Components
+The results are summarized in Table 3 and Figure 5, with detailed discussions provided below.
 Finetuning VS Training From Scratch.
-To assess whether adapting pretrained diffusion backbones to our DPBridge improves perception performance, we compare the difference between finetuning and training from scratch paradigm.
-As shown in the first row of Tab.3, fine-tuning significantly outperforms training from scratch under identical experimental settings.
-This difference is due to two factors: first, foundation models trained on large, diverse datasets capture universal visual priors, aiding knowledge transfer to dense prediction tasks.
+To assess whether adapting pretrained diffusion backbones improves perception performance, we conduct comparison between finetuning and training from scratch paradigm.
+As shown in the first row of Table 3, fine-tuning significantly outperforms training from scratch under identical experimental settings.
+This difference is due to two factors: first, foundation models trained on large, diverse datasets.
+They capture universal visual priors, aiding knowledge transfer to dense prediction tasks.
 Second, training from scratch requires more iterations for convergence.
-These observations suggest that building DPBridge on pretrained backbones is indeed advantageous.
+In addition, from the last two rows of Table 1 and Table 2, we can see that finetuning on SD2.1 exceeds that on SD1.5, suggesting that the quality of prior knowledge will also influence the performance.
+These observations indicate that building DPBridge on superior pretrained backbones is indeed advantageous.
 Distribution Alignment Normalization.
 Next, we evaluate whether aligning the marginal distributions between diffusion bridge process and standard diffusion process is beneficial.
-As shown in the second and fourth rows of Tab.3, incorporating a distribution alignment normalization step before inputting data into the noise prediction network noticeably improves performance.
+As shown in the second and fourth rows of Table 3, incorporating a distribution alignment normalization step before inputting data into the noise prediction network noticeably improves performance.
 Image Consistency Loss.
-Regarding the auxiliary image consistency loss, the comparison between the third and fourth rows in Tab.3 shows that while effective, its impact is less significant than that of distribution alignment normalization.
+Regarding the auxiliary image consistency loss, the comparison between the third and fourth rows in Table 3 shows that while effective, its impact is less significant than that of distribution alignment normalization.
 This is probably because it primarily influences regions where masked and valid areas overlap, which constitute only a relatively small proportion of the data.
 However, when integrated with other design components, the overall performance is significantly improved.
-5.4.2 Bridge Process Hyperparameters
-Parameterization and Backbone Selection.
-Apart from the noise prediction scheme used in our main experiments, we also consider how the choice of parameterization combined with the backbone selection influences performance.
-Results in Tab.4 reveal two key findings: (1) v-parameterization performs significantly worse compared with other two options, highlighting the need for further research on deriving objectives specific to the diffusion bridge process, and (2) ”noise prediction + SDv1.5” achieves the best results, indicating that the parameterization during fine-tuning should align with that of the backbone to maximize knowledge transfer efficiency.
 Number of Sampling Steps.
 We evaluate performance with varying numbers of sampling steps during inference.
-As shown in the lower part of Tab.4, increasing sampling steps initially improves perception performance rapidly before leveling off.
-Notably, compared to the results in Tab.1, where all other diffusion-based methods are evaluated with 50 steps setting, DPBridge achieves comparable performance with fewer steps, demonstrating the efficiency of our proposed method.
+As shown in Figure 5, increasing sampling steps initially improves perception performance rapidly and reach saturation after 10 steps.
+Notably, compared to the results in Table 1, where all other diffusion-based methods are evaluated with 50 steps setting, DPBridge achieves comparable performance with fewer steps, demonstrating the efficiency of our proposed method.
 6 Conclusion
-In this paper, we present DPBridge, a latent diffusion bridge model to tackle general dense prediction tasks.
-We establish a direct mapping between input RGB images and their corresponding dense maps in the latent space based on fully-tractable diffusion bridge process.
+In this paper, we present DPBridge, a latent diffusion bridge model to tackle general dense prediction tasks through generative pipeline.
+We establish direct mapping between input images and corresponding signal maps in the latent space based on fully-tractable diffusion bridge process.
 In this way, the prior distribution is no longer constrained to uninformative noise, but instead can be chosen as more structurally correlated ones.
 In addition, we introduce effective finetuning strategies to adapt large pretrained diffusion backbone to our bridge model construction, leveraging its rich prior knowledge to boost training speed and achieve remarkable perception performance.
-Future research directions include deriving v-prediction parameterization of diffusion bridge process, and extending to other dense prediction tasks.
-Supplementary Material
-In the supplementary material, we’ll provide contents that are not included in the main paper due to page limitations.
-We first fill in the derivation details of the loss function.
-Afterwards we will provide the additional quantitative comparisons with more baseline methods.
-Finally, we will present qualitative examples that apply DPBridge on more dense prediction tasks.
+Future research directions include deriving other parameterization of diffusion bridge process, and extending to more dense prediction tasks.
+Impact Statement
+This paper presents work whose goal is to advance the field of Machine Learning and we wish to spark further research on image generation and perception.
+There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
 Appendix A Derivation Details
-In this section, we’ll provide derivation details of the maximum likelihood loss presented in Sec.4.5.1 .
-Its core is the posterior distribution $q(\mathbf{z}_{t-1}|\mathbf{z}_{t},\mathbf{z}_{0},\mathbf{z}_{T})$ as in Eqn.(5) , whose calculation involves three terms: $q(\mathbf{z}_{t}|\mathbf{z}_{t-1},\mathbf{z}_{T})$, $q(\mathbf{z}_{t-1}|\mathbf{z}_{0},\mathbf{z}_{T})$, and $q(\mathbf{z}_{t}|\mathbf{z}_{0},z\mathbf{z}_{T})$.
+In this section, we’ll provide derivation details of the maximum likelihood loss presented in Section 4.4.1.
+Its core is the posterior distribution $q(\mathbf{z}_{t-1}|\mathbf{z}_{t},\mathbf{z}_{0},\mathbf{z}_{T})$ as in Equation 5, whose calculation involves three terms: $q(\mathbf{z}_{t}|\mathbf{z}_{t-1},\mathbf{z}_{T})$, $q(\mathbf{z}_{t-1}|\mathbf{z}_{0},\mathbf{z}_{T})$, and $q(\mathbf{z}_{t}|\mathbf{z}_{0},z\mathbf{z}_{T})$.
 The last two term of the posterior can be retrieved directly through the marginal distribution of diffusion bridge’s forward process, but further derivation is required for the first term.
-Specifically, as specified in Eqn.(3)
+Specifically, as specified in Equation 3:
 $\begin{split}q(\mathbf{z}_{t}|\mathbf{z}_{0},\mathbf{z}_{T})&=\mathcal{N}(m_{t}\mathbf{z}_{0}+n_{t}\mathbf{z}_{T},\sigma_{t}^{2}\mathbf{I})\\ \mathbf{z}_{t}&=m_{t}\mathbf{z}_{0}+n_{t}\mathbf{z}_{T}+\sigma_{t}\boldsymbol{\epsilon}_{t}\end{split}$ (14)
 $\begin{split}q(\mathbf{z}_{t-1}|\mathbf{z}_{0},\mathbf{z}_{T})&=\mathcal{N}(m_{t-1}\mathbf{z}_{0}+n_{t-1}\mathbf{z}_{T},\sigma_{t-1}^{2}\mathbf{I})\\ \mathbf{z}_{t-1}&=m_{t-1}\mathbf{z}_{0}+n_{t-1}\mathbf{z}_{T}+\sigma_{t-1}\boldsymbol{\epsilon}_{t-1}\end{split}$ (15)
-To derive the transition probability $q(\mathbf{z}_{t}|\mathbf{z}_{t-1},\mathbf{z}_{T})$, we substitute $\mathbf{z}_{0}$ in Eqn.(14) with derivation from Eqn.(15), obtain
+To derive the transition probability $q(\mathbf{z}_{t}|\mathbf{z}_{t-1},\mathbf{z}_{T})$, we substitute $\mathbf{z}_{0}$ in Equation 14 with derivation from Equation 15, obtain
 $\mathbf{z}_{t}$ $=a_{t}\mathbf{z}_{t-1}+b_{t}\mathbf{z}_{T}+\delta_{t}\boldsymbol{\epsilon}$ (16)
@@
 Thus, we can get $q(\mathbf{z}_{t}|\mathbf{z}_{t-1},\mathbf{z}_{T})=\mathcal{N}(a_{t}\mathbf{z}_{t-1}+b_{t}\mathbf{z}_{T},\delta_{t}^{2}\mathbf{I})$, which is also a Gaussian distribution.
-Putting all these terms back to Eqn.(5) , we’ll have:
+Putting all these terms back to Equation 5, we’ll have:
 $q(\mathbf{z}_{t-1}|\mathbf{z}_{t},\mathbf{z}_{0},\mathbf{z}_{T})\sim\mathcal{N}(\tilde{\boldsymbol{\mu}}_{t}(\mathbf{z}_{t},\mathbf{z}_{0},\mathbf{z}_{T}),\tilde{\sigma}_{t}^{2}\mathbf{I})$ (17)
@@
 $=\frac{\sigma_{t-1}^{2}}{\sigma_{t}^{2}}a_{t}\mathbf{z}_{t}+\frac{\delta_{t}^{2}}{\sigma_{t}^{2}}m_{t-1}\mathbf{z}_{0}+\bigg{(}\frac{\delta_{t}^{2}}{\sigma_{t}^{2}}n_{t-1}-\frac{\sigma_{t-1}^{2}}{\sigma_{t}^{2}}a_{t}b_{t}\bigg{)}\mathbf{z}_{T}$ (20)
-whose results are equivalent to Eqn.(6) and Eqn.(7).
+whose results are equivalent to Equation 6 and Equation 7.
 Accordingly, maximizing the Evidence Lower Bound (ELBO) is equivalent to minimizing the KL divergence between ground truth and predicted posterior:
@@
 $=\mathbb{E}_{t,\mathbf{z}_{0},\mathbf{z}_{t},\mathbf{z}_{T}}\left[\left\|\boldsymbol{\epsilon}-\boldsymbol{\epsilon}_{\theta}(\mathbf{z}_{t},\mathbf{z}_{T},t)\right\|^{2}\right]$
-which is the same as in Eqn.(8) and Eqn.(9) . This concludes our derivation.
-Appendix B Additional Quantitative Results
-Next, we provide additional quantitative results on depth estimation (Tab.5) and surface normal prediction (Tab.6) tasks.
-Compared with Tab.1 and Tab.2 presented in the main paper, we conduct comparisons with more baseline methods, including feed-forward approaches and diffusion-based methods that is trained on more advanced Stable Diffusion (SD) v2.1 backbone.
-In depth estimation, DPBridge consistently matches state-of-the-art methods, while demonstrating robust generalization across datasets.
-For example, it achieves 6.9 AbsRel, 95.7 % $\delta_{1}$ on NYUv2, and 7.5 AbsRel, 94.7% $\delta_{1}$ on ScanNet.
-Similarly, for surface normal prediction, DPBridge also delivers competitive results, achieving mean error of 18.1 and $11.25^{\circ}$ accuracy of $54.2\%$ on NYUv2, as well as maintaining strong performance on other datasets like iBims-1 (mean: 24.3, $11.25^{\circ}:51.2\%$) and Sintel (mean: 41.2, $11.25^{\circ}:19.4\%$) and.
-These results highlight DPBridge’s versatility to generalize well across tasks and benchmarks, being comparable to both feed-forward and diffusion-based methods while offering a unified, efficient paradigm to handle dense prediction tasks.
-Method Model Type NYUv2 KITTI ETH3D ScanNet DIODE
-[…]
-DiverseDepth [57] 11.7 87.5 19.0 70.4 22.8 69.4 10.9 88.2 37.6 63.1
-MiDaS [43] 11.1 88.5 23.6 63.0 18.4 75.2 12.1 84.6 33.2 71.5
-LeReS [58] 9.0 91.6 14.9 78.4 17.1 77.7 9.1 91.1 27.1 76.6
-Omnidata [19] 7.4 94.5 14.9 83.5 16.6 77.8 7.5 93.2 33.9 74.2
-HDN [60] 6.9 94.8 11.5 86.7 12.1 83.3 8.0 93.6 24.6 78.0
-DPT [44] 9.8 90.3 10.0 90.1 7.8 94.6 8.2 93.8 18.2 75.8
-Marigold* [32] 6.0 95.9 10.5 90.4 7.1 95.1 6.9 94.5 31.0 71.2
-DepthFM* [24] 6.5 95.6 8.3 93.4 - - - - 22.5 80.0
-Geowizard* [20] 5.7 96.2 14.4 82.0 7.5 94.3 6.1 95.8 31.4 77.1
-E2E-FT* [21] 5.4 96.4 9.6 92.1 6.4 95.9 5.8 96.5 30.3 77.6
-Marigold [32] 8.6 93.6 13.8 82.6 9.3 93.2 10.4 89.7 33.6 73.9
-Geowizard [20] 8.5 93.7 16.2 75.6 9.1 93.1 9.0 93.0 32.0 72.2
-E2E-FT [21] 8.2 94.0 15.4 77.9 10.6 89.8 9.2 92.3 32.9 73.7
-ControlNet [61] 45.5 54.6 55.2 40.2 30.5 58.8 25.7 68.3 52.1 50.1
-InstructPix2Pix [6] 35.2 60.0 47.7 51.2 26.1 66.7 21.4 75.3 44.8 58.0
-I2SB [37] 24.9 73.4 37.7 57.9 40.6 52.0 21.2 66.1 40.2 59.1
-BBDM [36] 27.9 65.5 47.2 45.1 46.4 48.8 22.7 67.0 34.6 64.7
-DPBridge (Ours) 6.9 95.7 11.2 87.3 8.1 93.2 7.5 94.7 29.7 78.6
-Table 5: Quantitative comparison of DPBridge with state-of-the-art affine-invariant depth estimation methods on zero-shot benchmarks.
-All metrics are presented in percentage terms.
-FF, DM, and BM correspond to feed-forward, diffusion models, and bridge models.
-The superscript * indicates that this model is trained based on Stable Diffusion v2.1, otherwise is based on Stable Diffusion v1.5.
-Method Model Type NYUv2 ScanNet iBims-1 Sintel
-[…]
-EESNU [3] 16.2 58.6 - - 20.0 58.5 42.1 11.5
-Omnidata v1 [19] 23.1 45.8 22.9 47.4 19.0 62.1 41.5 11.4
-Omnidata v2 [31] 17.2 55.5 16.2 60.2 18.2 63.9 40.5 14.7
-DSINE [2] 16.4 59.6 16.2 61.0 17.1 67.4 34.9 21.5
-Marigold* [32] 18.8 55.9 17.7 58.8 18.4 64.3 39.1 14.9
-Geowizard* [20] 19.1 49.5 17.3 53.7 19.5 61.6 40.4 13.2
-E2E-FT* [21] 16.5 60.4 14.7 66.1 16.1 69.7 33.5 22.3
-Marigold [32] 22.3 40.0 23.3 37.6 28.0 38.5 52.1 8.1
-Geowizard [20] 21.7 44.2 22.0 42.1 28.2 37.9 45.8 11.0
-E2E-FT [21] 17.7 57.8 17.4 57.4 18.0 64.6 43.9 14.1
-DPBridge 18.1 54.2 18.0 54.5 24.3 51.2 41.2 19.4
-Table 6: Quantitative comparison of DPBridge with state-of-the-art surface normal prediction methods on zero-shot benchmarks.
-All metrics are presented in percentage terms.
-FF, DM, and BM correspond to feed-forward, diffusion models, and bridge models.
-The superscript * indicates that this model is trained based on Stable Diffusion v2.1, otherwise is based on Stable Diffusion v1.5.
-Furthermore, we also evaluate the robustness of DPBridge system by perturbing the input image with different types and different level intensities of noise.
-The evaluation is conducted under both depth estimation and surface normal prediction tasks.
+which is the same as in Equation 8 and Equation 9. This concludes our derivation.
+Appendix B Experiments
+B.1 Network Architecture
+The basic structure of our network follows the original design of Stable Diffusion.
+To better utilize visual information ($\mathbf{z}_{T}$ in our case), we concatenating both intermediate samples and input image along the feature dimension and send in the noise prediction network as a whole.
+To address the dimension incompatibility with pretrained backbone, the input channels of the network are thus doubled to accommodate the concatenated input.
+In addition, as suggested by (Ke et al., 2024), we duplicate the weight tensor of these accommodated layers and divide the values by two to preserve the input magnitude to subsequent layers.
+As DPBridge is an image-conditioned generative framework, we employ null-text embedding for all cross-attention layers to eliminate the impact of text prompts.
+B.2 Additional Quantitative Results
+In this section, we further evaluate the robustness of DPBridge system by perturbing the input image with different types and different level intensities of noise.
+In this experiment setting, we use Stable Diffusion 1.5 as pretrained backbone.
+The evaluation is conducted using under both depth estimation and surface normal prediction tasks.
 We use NYUv2 and KITTI to report the depth estimation results, as for surface normal prediction, we report the results on NYUv2 and iBims-1.
-Tab.
-7 reveals that the model’s performance degrades as noise intensity increases, but the model performance remains acceptable under moderate-level-intensity noise.
+Table 4 reveals that the model’s performance degrades as noise intensity increases, but the model performance remains acceptable under moderate-level-intensity noise.
 Among noise types, uniform noise shows the mildest effect on degradation, while salt & pepper and high-intensity gaussian noise cause the most significant performance decline.
@@
 During all experiments, the input images are normalized between -1 to 1.
-Appendix C Additional Qualitative Results
+B.3 Additional Qualitative Results
 To further validate the effectiveness and generalization of DPBridge, we provide qualitative results for various dense prediction tasks, including semantic segmentation, optical flow prediction, and style transfer, by framing them as continuous-valued image-to-image translation problems.
@@
 To accommodate this, we adapt the UNet architecture to incorporate both RGB images as conditions during inference and convert the flow map into RGB space, predicting corresponding colors rather than uv coordinates.
-The following figures (Fig.5, Fig.6, Fig.7, Fig.8) illustrate that DPBridge achieves visually satisfactory results across different tasks.
+The following figures (Figure 6, Figure 7, Figure 8, Figure 9) illustrate that DPBridge achieves visually satisfactory results across different tasks.
 Note that only qualitative assessments are presented here; detailed implementations and quantitative evaluations are left as directions for future research.
-Figure 5: Qualitative examples of applying DPBridge to semantic segmentation task on CityScape [14] dataset.
+Figure 6: Qualitative examples of applying DPBridge to semantic segmentation task on CityScape (Cordts et al., 2016) dataset.
 We transform discrete labels into color maps and us our bridge model to predict colors in RGB space instead of categorical labels.
-Figure 6: Qualitative examples of applying DPBridge to semantic segmentation task on CelebAMask-HQ [35] dataset.
+Figure 7: Qualitative examples of applying DPBridge to semantic segmentation task on CelebAMask-HQ (Lee et al., 2020) dataset.
 We transform discrete labels into color maps and us our bridge model to predict colors in RGB space instead of categorical labels.
-Figure 7: Qualitative examples of applying DPBridge to optical flow prediction task on MPI-Sintel [7] dataset.
+Figure 8: Qualitative examples of applying DPBridge to optical flow prediction task on MPI-Sintel (Butler et al., 2012) dataset.
 The input image is the overlayed RGB pair, and we transform the flow map into RGB space so that the diffusion bridge process can be constructed accordingly.
-Figure 8: Qualitative examples of applying DPBridge to style transfer task on Face-to-Comics [55] dataset.
+Figure 9: Qualitative examples of applying DPBridge to style transfer task on Face-to-Comics (Sxela, 2021) dataset.

### v3

@@
 Abstract
-Diffusion models have shown remarkable capabilities in modeling complex data distributions by transforming noise into structured data through stochastic processes.
-However, when applied to dense prediction tasks whose goal is to capture per-pixel relationships between RGB images and dense signal maps, starting the sampling process from an uninformative Gaussian noise often leads to inefficient sampling and long latency.
-To overcome these challenges, we propose DPBridge, a generative framework that establishes direct mapping between input RGB images and dense signal maps based on a tractable bridge process.
-Furthermore, we introduce finetuning strategies to leverage a pretrained large-scale image diffusion backbone, enjoying its rich visual prior knowledge to enable both efficient training and robust generalization.
-Experiments show that DPBridge achieves competitive performance compared to both feed-forward and diffusion-based approaches across various benchmarks, validating its effectiveness and adaptability.
-Machine Learning, ICML
+Diffusion models demonstrate remarkable capabilities in capturing complex data distributions and have achieved compelling results in many generative tasks.
+While they have recently been extended to dense prediction tasks such as depth estimation and surface normal prediction, their full potential in this area remains under-explored.
+In dense prediction settings, target signal maps and input images are pixel-wise aligned.
+This makes conventional noise-to-data generation paradigm inefficient, as input images can serve as more informative prior compared to pure noise.
+Diffusion bridge models, which support data-to-data generation between two general data distributions, offer a promising alternative, but they typically fail to exploit the rich visual priors embedded in large pretrained foundation models.
+To address these limitations, we integrate diffusion bridge formulation with structured visual priors and introduce DPBridge, the first latent diffusion bridge framework for dense prediction tasks.
+Our method presents three key contributions: (1) a tractable reverse transition kernel for diffusion bridge process, enabling maximum likelihood training scheme for better compatibility with pretrained backbones; (2) a distribution-aligned normalization technique to mitigate the discrepancies between the bridge and standard diffusion processes; and (3) an auxiliary image consistency loss to preserve fine-grained details.
+Experiments across extensive benchmarks validate that our method consistently achieves superior performance, demonstrating its effectiveness and generalization capability under different scenarios.
+1 Introduction
+Dense prediction refers to a class of problems in which the goal is to assign a prediction value for every pixel in the input image.
+It encompasses a variety of fundamental tasks, including depth estimation [56, 39, 15], surface normal prediction [2, 22, 3], semantic segmentation [53, 8, 51], optical flow [49, 23, 54], etc.
+Recently, image diffusion models have been applied to dense prediction tasks by framing them as conditional generation problems.
+Leveraging the ability to capture complex data distributions, diffusion models serve as powerful priors and yield promising results [28, 17, 16].
+However, these approaches exhibit inherent limitations stemming from the underlying generative mechanism.
+Specifically, the denoising process always starts from Gaussian noise, which carries minimal information about the target signal distribution.
+In such cases, the distributional gap between initial Gaussian prior and target distribution remains large, necessitating more sampling steps and often reliance on manually designed post-hoc guidance [9, 46, 10] or customized generation pipeline [35].
+This not only increases computational overhead but can also degrade performance and prolong convergence.
+Instead of starting from noise, a more effective alternative is to directly initiate the generative process from the input.
+Since RGB images are pixel-wise aligned with their corresponding dense signal maps, they serve as informative priors.
+This motivates the exploration of diffusion bridge models [34, 37, 14, 33], another line of computational frameworks based on stochastic processes with fixed start and end states, designed to model probability flows between two given distributions.
+We argue that diffusion bridge models are more suitable for dense prediction tasks, as their data-to-data generation paradigm naturally leverage the structural correspondence between inputs and outputs, thereby mitigating inefficiencies inherent in conventional noise-to-data scheme.
+Despite theoretical feasibility, diffusion bridge-based methods [65, 31, 32] have not yet demonstrated competitive performance compared with standard diffusion counterparts.
+This performance gap largely stems from the fact that most bridge models are trained from scratch via score matching objectives, making them incompatible with commonly used pretrained diffusion backbone.
+As a result, they fail to exploit the embedded semantic and spatial prior knowledge, limiting their expressiveness and generalization ability in dense prediction settings.
 Figure 1: We present DPBridge to tackle general dense prediction tasks by establishing conditional probability transport between RGB images and corresponding dense maps in the latent space based on a tractable diffusion bridge process.
-In addition, we introduce finetuning strategies to adapt large pretrained diffusion backbone to construct our bridge, leveraging its rich prior knowledge to enable both efficient training and robust generalization.
-1 Introduction
-Dense prediction refers to the type of problems whose primary goal is to assign a value at every spatial location in the input data, generating output signal maps that possess same structure and size as inputs.
-It incorporates various typical tasks, including depth estimation (Yang et al., 2024; Ranftl et al., 2021; Eftekhar et al., 2021), surface normal prediction (Bae & Davison, 2024; Hu et al., 2024; Bae et al., 2021), semantic segmentation (Chen, 2017; Cheng et al., 2022; Wang et al., 2023; Weber et al., 2021), optical flow (Teed & Deng, 2020; Huang et al., 2022; Weinzaepfel et al., 2023), etc.
-Recently, image diffusion models have been applied to dense prediction tasks by formulating them as conditional generation problems.
-Their ability to capture complex data distributions makes them effective priors, leading to highly favorable results (Ke et al., 2024; Garcia et al., 2024; Fu et al., 2025).
-Following a common practice, these methods diffuse signal maps into pure noise and subsequently leverage input RGB images as conditions to guide the generation process.
-Nonetheless, there still exists limitations in diffusion-based solutions, mostly stem from the underlying modeling mechanism.
-The denoising process of a standard diffusion model always starts from Gaussian noise, which contains little information about the target signal distribution.
-In such cases, the distance between prior Gaussian distribution and target distribution remains large, requiring inevitably more sampling steps as well as ad-hoc conditioning mechanisms, at the risk of degraded performance and prolonged convergence time.
-Instead of starting from noise, one feasible alternative is to directly initiate the generative process from input RGB images, as they are structurally correlated to target signal maps and can naturally serve as more informative prior.
-This leads to the exploration of a new family of computational frameworks, i.e.
-diffusion bridge models (Liu et al., 2022; Zhou et al., 2023; Li et al., 2023; Liu et al., 2023).
-They are based on stochastic process with fixed starting and ending points and are adept at modeling conditional probability paths.
-By establishing direct connections between different distributions, this line of framework provides a more natural and efficient way to model probability transport compared with previous efforts by reducing the dependence on manually designed post-hoc guidance (Chung et al., 2022a; Song et al., 2023; Chung et al., 2022b) or customized generation pipeline (Meng et al., 2021).
-We therefore advocate that adopting diffusion bridge model will be a better fit to dense prediction tasks compared with conventional diffusion-based counterparts.
-Motivated by this, we propose Dense Prediction Bridge (DPBridge), a latent diffusion bridge model to handle general dense prediction tasks through generative pipeline.
-In essence, we modify the standard diffusion process to directly bridge the latents of RGB images and corresponding dense signal maps, and enforce the model to learn this process through maximum likelihood estimation.
-In this way the generative process thus becomes a series of transformations that iteratively synthesize desired signal maps starting from RGB images rather than pure noise, enabling faster inference speed as well as better generation quality.
-Furthermore, we introduce novel finetuning strategies to better adapt pretrained diffusion backbone to DPBridge construction.
-The equipped rich visual prior not only helps boosting the perception performance as well as generalization capability, but also significantly speeds up training convergence.
-Figure 1 is a high-level overview of DPBridge, while also demonstrating its effectiveness on three representative dense prediction tasks, i.e.
-depth estimation, surface normal prediction, and semantic segmentation, proving that our framework is theoretically universal and can be instantiated to different scenarios.
-To summarize, our contributions are as follows:
-  • We propose the DPBridge framework to address dense prediction tasks within a generative pipeline by formulating them as instances of latent diffusion bridges.
-  • We propose finetuning strategies to effectively adapt large pretrained backbone to diffusion bridges, allowing for exploitation of visual prior knowledge and training efficiency improvements.
-  • Our proposed method exhibits competitive performance across different tasks in terms of both quantitative metrics and visual inspection, highlighting its effectiveness and generalization capability.
+In this paper, we propose Dense Prediction Bridge (DPBridge), the first latent diffusion bridge model tailored for general dense prediction tasks by integrating the strengths of diffusion bridge modeling with visual priors from pretrained backbones.
+Specifically, we derive a fully tractable bridge process to connect the latent representations of RGB images and their corresponding dense signal maps, and formulate the generation process as a sequence of transformations that progressively synthesize the target maps from input data.
+Furthermore, we introduce finetuning strategies to better adapt pretrained diffusion backbones for DPBridge construction.
+Leveraging the rich visual priors not only enhances perceptual performance and generalization capability, but also significantly accelerates training convergence.
+Figure 1 provides a high-level overview of DPBridge framework and illustrates its effectiveness across three representative dense prediction tasks: depth estimation, surface normal prediction, and semantic segmentation.
+These results demonstrate that our framework is both theoretically general and practically adaptable to diverse dense prediction scenarios.
+Contributions of this paper can be summarized as follows:
+  • We propose DPBridge, a generative framework to address dense prediction tasks by formulating them as instances of latent diffusion bridges.
+  • We derive a tractable reverse transition kernel for diffusion bridge process, enabling maximum likelihood training scheme.
+This is compatible with pretrained backbones, thus improving both training stability and the exploitation of visual prior knowledge.
+  • We introduce a distribution-aligned normalization technique that mitigates distributional discrepancies between the bridge and standard diffusion processes, allowing for more effective finetuning.
+In addition, by incorporating auxiliary image consistency loss, we can better preserve fine-grained spatial details often lost in latent representations, enhancing per-pixel prediction accuracy.
+  • Our method exhibits competitive performance across different tasks in terms of both quantitative metrics and visual inspection.
 2 Related Works
-2.1 Dense Prediction
-Dense predictions can be characterized as a conditional distribution modeling task and common solutions can be categorized into discriminative and generative ones.
+Dense Prediction.
+Dense predictions can be characterized as conditional distribution modeling and common solutions are categorized into discriminative and generative ones.
 Discriminative methods directly map input data to structured outputs in a feed-forward manner.
-(Ranftl et al., 2021) introduces attention mechanism and transformer architecture to better capture both local and global context feature.
-(Yang et al., 2024) explore the usage of large-scale synthetic and pseudo-labeled images to enhance model’s capacity and robustness.
-(Bae & Davison, 2024) incorporates per-pixel ray directions and models the relative rotations between neighboring pixels.
-Recently, diffusion-based generative models have been applied in this field and showed remarkable performance.
-(Ke et al., 2024) finetunes pretrained diffusion model, leveraging the rich visual knowledge within to enhance depth estimation.
-(Fu et al., 2025) takes advantage of diffusion priors to jointly estimate depth and surface normals from single image, enhancing detail preservation by decoupling complex data distributions into distinct sub-distributions.
-(Gui et al., 2024) exploits flow matching to facilitate the distribution shift between images and depth maps.
+[39] introduces transformer architecture and fusion mechanism to better capture both local and global context feature.
+[56] explore the usage of large-scale synthetic and pseudo-labeled images to enhance model’s capacity and robustness.
+[2] incorporates per-pixel ray directions and models the relative rotations between neighboring pixels.
+Recently, diffusion-based generative methods have been applied in this field and showed remarkable performance.
+[28] finetunes pretrained diffusion backbone, leveraging the rich visual knowledge within to enhance depth estimation.
+[16] takes advantage of diffusion priors to jointly estimate depth and surface normals from single image, enhancing detail preservation by decoupling complex data distributions into distinct sub-distributions.
+[19] exploits flow matching to facilitate the distribution shift between images and depth maps.
 Our work follow conventional generative pipeline, but instead of starting from pure noise, we establish direct connections between input image and target signal map distributions, and model conditional probability transport in iterative denoising manner.
-2.2 Diffusion and Bridge Models
-Diffusion models (Ho et al., 2020; Song et al., 2020b) are a family of generative models that are able to capture complex data distributions through a sequence of forward and reverse processes.
-They accelerate the development of modern generative systems, and have consistently obtained state-of-the-art performance across various domains, such as image and video synthesis (Rombach et al., 2022; Ho et al., 2022), controlled editing (Zhang et al., 2023; Batzolis et al., 2021), inverse problem solving (Chung et al., 2022a; Song et al., 2023), 3D assets generation (Xiang et al., 2024; Zhang et al., 2024), and human-centered applications (Ji et al., 2024b, a; Ji & Li, 2024).
-Diffusion bridge models have gained increasing attention in generative modeling in recent years (Liu et al., 2022; Zhou et al., 2023; Liu & Wu, 2023; De Bortoli et al., 2021).
-They offer a novel perspective for translating between different distributions and enable a more effective data-to-data generation paradigm.
+Diffusion and Bridge Models.
+Diffusion models [20, 47] are a family of generative models that are able to capture complex data distributions through a sequence of forward and reverse processes.
+They accelerate the development of modern generative systems, and have consistently obtained state-of-the-art performance across various domains, such as image and video synthesis [41, 21], controlled editing [62, 4], inverse problem solving [9, 46], 3D assets generation [55, 63], and human-centered applications [26, 25, 24].
+Diffusion bridge models have gained increasing attention in generative modeling in recent years [37, 34, 65, 33, 14].
+They offer a novel perspective for translating between distributions and enable a more effective data-to-data generation paradigm.
 Recent works have explored different types of bridge processes and their applications.
-(Liu et al., 2023; Li et al., 2023; Yue et al., 2023) develop Schrödinger bridge, Brownian bridge, and Ornstein-Uhlenbeck bridge frameworks, applying them to image restoration tasks such as inpainting, super-resolution.
-(Chung et al., 2024) proposes a modified inference procedure that enforces data consistency.
-(Chen et al., 2023) introduces bridge models to text-to-speech systems, enhancing synthesis quality and sampling efficiency.
+[32, 31, 60] develop Schrödinger bridge, Brownian bridge, and Ornstein-Uhlenbeck bridge frameworks, applying them to image restoration tasks such as inpainting, super-resolution.
+[11] proposes a modified inference procedure that enforces data consistency.
+[7] introduces bridge models to text-to-speech systems, enhancing synthesis quality and sampling efficiency.
 To the best of our knowledge, no prior work has explored the use of bridge models for dense prediction tasks to facilitate distribution transport between images and dense signal maps, despite their structural closeness.
 Our work aims to address this research gap.
+Conceptually, DPBridge is most related to DDBM [65], but the two differ in several key aspects: DDBM is trained from scratch using a score matching objective, whereas DPBridge is finetuned from large pretrained diffusion backbones using a maximum likelihood objective.
+In addition, we introduce training strategies tailored for efficient finetuning, enabling better use of visual priors and significantly improving performance on dense prediction settings.
 3 Preliminary
@@
 One natural way of its construction is to derive the conditioned process from a reference diffusion process given terminal constraints.
-We can use Doob’s h-transform technique (Oksendal, 2013; Särkkä & Solin, 2019) to modify a reference process $\mathbb{Q}$ defined in Equation 1 to ensure that the event $\mathbf{z}_{T}=\mathbf{x}$ happens, where $\mathbf{x}$ is a predefined value.
+We can use Doob’s h-transform technique [36, 42] to modify a reference process $\mathbb{Q}$ defined in Equation 1 to ensure that the event $\mathbf{z}_{T}=\mathbf{x}$ happens, where $\mathbf{x}$ is a predefined value.
 The conditioned process $\mathbb{Q}^{\mathbf{x}}(\cdot):=\mathbb{Q}\left(\cdot|\mathbf{z}_{T}=\mathbf{x}\right)$ thus follows the law of
@@
 $q(\mathbf{z}_{t}|\mathbf{z}_{0},\mathbf{z}_{T})$ $=\mathcal{N}(m_{t}\mathbf{z}_{0}+n_{t}\mathbf{z}_{T},\sigma_{t}^{2}\mathbf{I})$ (3)
-$m_{t}$ $=\alpha_{t}\left(1-\frac{\text{SNR}_{T}}{\text{SNR}_{t}}\right)$
-$n_{t}$ $=\frac{\text{SNR}_{T}}{\text{SNR}_{t}}\frac{\alpha_{t}}{\alpha_{T}}$
-$\sigma_{t}$ $=\sigma_{t}\sqrt{1-\frac{\text{SNR}_{T}}{\text{SNR}_{t}}},$
+$m_{t}=\alpha_{t}\left(1-\frac{\text{SNR}_{T}}{\text{SNR}_{t}}\right),\quad n_{t}$ $=\frac{\text{SNR}_{T}}{\text{SNR}_{t}}\frac{\alpha_{t}}{\alpha_{T}},\quad\sigma_{t}=\sigma_{t}\sqrt{1-\frac{\text{SNR}_{T}}{\text{SNR}_{t}}}$
 where $\text{SNR}_{t},\alpha_{t},\sigma_{t}^{2}$ are all ingredients defined in the reference process $\mathbb{Q}$.
@@
 4.1 Problem Formulation
-We formulate different dense prediction tasks as instances of image-conditioned generation problems and propose DPBridge to model the conditional distributions $p(\mathbf{y}|\mathbf{x})$, where $\mathbf{y}\in\mathbb{R}^{H\times W\times C}$ indicates the desired dense signal map (e.g.
+We formulate different dense prediction tasks as instances of image-conditioned generation problems and propose DPBridge to model the conditional distributions $p(\mathbf{y}|\mathbf{x})$, where $\mathbf{y}\in\mathbb{R}^{H\times W\times C}$ denotes the target dense signal map (e.g.
 depth, surface normal, segmentation map), and $\mathbf{x}\in\mathbb{R}^{H\times W\times 3}$ is the input RGB image.
 In the following sections, we’ll elaborate the details of DPBridge framework, which is constructed based on pretrained diffusion backbone through proper finetuning.
+4.2 Dense Prediction Bridge
+Given paired training data, a straightforward approach is to construct the diffusion bridge model from scratch, following the formulation in Section 3.2.
+However, this approach is extremely resource-demanding, making it hard to implement with limited computation power.
+What’s more, it lacks the ability to harness rich visual prior embedded in the foundation models pretrained with large-scale data,which have been shown to significantly enhance both perceptual quality and generalization capability in dense prediction tasks.
+To address these limitations and fully exploit the visual prior, our DPBridge is build upon pretrained diffusion backbones through finetuning.
+We explore several strategies to handle the misalignment between bridge process and standard diffusion process, facilitating more efficient knowledge transfer.
+An overview of our training pipeline is provided in Figure 2, followed by a detailed explanation of each component.
+4.3 Latent Space Implementation
+We follow common practice to perform diffusion steps in a compressed yet perceptually aligned latent space, provided by an independently trained variational autoencoder (VAE).
+Specifically, both the dense signal map and RGB image are encoded using the VAE encoder, yielding latent representations $\mathbf{z}_{0}=\mathcal{E}(\mathbf{y})$ and $\mathbf{z}_{T}=\mathcal{E}(\mathbf{x})$, which serve as the initial and terminal points of the diffusion bridge process.
+In this way the diffusion bridge process is operated entirely in the latent space.
+To recover the original signal map, we use the VAE decoder to decode the predicted latent $\mathbf{y}=\mathcal{D}(\hat{\mathbf{z}}_{0})$.
+For clarity, all subsequent derivations in this paper are conducted in latent space.
 Figure 2: Overview of our training pipeline.
-To overcome the discrepancy with standard diffusion process and facilitate finetuning, we derive the maximum likelihood training objective of diffusion bridge process, propose a distribution aligned normalization strategy, and assist the training with an auxiliary image consistency loss.
-4.2 Dense Prediction Bridge
-Given paired training data, a naive implementation is to construct the diffusion bridge model from scratch according to Section 3.2.
-However, this approach is extremely resource-demanding, making it hard to implement with limited computation power.
-What’s more, it lacks the ability to harness rich visual prior knowledge inherent in the foundation models, which have been proven significant in terms of boosting both perception quality and generalization performance in dense prediction tasks.
-Therefore, our DPBridge is build upon pretrained diffusion backbones through finetuning and we explore several strategies to handle the misalignment between bridge process and standard diffusion process, aiming for more efficient knowledge transfer.
-Figure 2 gives an overview of the proposed finetuning pipeline, and next we’ll explain the design of each component.
-4.3 Latent Space Implementation
-Instead of directly learning in image pixel space, we follow common practice to perform diffusion steps in a compressed but perceptually aligned latent space, provided by an independently trained variational autoencoder (VAE).
-Specifically, both the dense signal maps and RGB images are converted by the encoder to obtain $\mathbf{z}_{0}=\mathcal{E}(\mathbf{y})$ and $\mathbf{z}_{T}=\mathcal{E}(\mathbf{x})$, serving as the initial and terminal points.
-In this way the diffusion bridge process is henceforth learned in the latent space.
-To retrieve original signal map we can use the decoder to decode the predicted latents $\mathbf{y}=\mathcal{D}(\hat{\mathbf{z}}_{0})$.
-In later part of this paper, with little abuse of notation, all derivations are conducted in the latent space.
+We derive the maximum likelihood training objective of diffusion bridge process, propose a distribution-aligned normalization strategy, and assist the training with an auxiliary image consistency loss.
 4.4 Training Objective
-4.4.1 Maximum Likelihood Loss
-Widely-used techniques for learning the reverse process that corresponds to Equation 2 is to follow the theory of (Anderson, 1982), use a neural network to approximate the score function and apply score matching.
-We notice that analytical form of diffusion bridge’s reverse transition kernel is directly obtainable, therefore the usage of maximum likelihood training can also be applied, which is performed by optimizing the Evidence Lower Bound (ELBO).
-It yields a more stable loss function compared with score matching and better aligns with the pretrained backbone.
-Given the specified terminal state $\mathbf{z}_{T}$, the expectation of data log-likelihood possesses the following ELBO formulation:
+Maximum Likelihood Loss.
+A widely adopted approach for learning the reverse process associated with Equation 2 follows the theoretical framework of [1], where a neural network is trained to approximate the score function [65].
+In contrast, we observe that in the diffusion bridge setting, the reverse transition kernel admits a tractable analytical form.
+This insight enables us to adopt a maximum likelihood training paradigm in place of score matching.
+Since widely used latent diffusion models like Stable Diffusion [41] are pretrained under a likelihood-based loss function, finetuning with a consistent objective is more compatible, which helps preserve prior knowledge and improves training stability.
+Specifically, the training process is performed by optimizing the Evidence Lower Bound (ELBO).
+Given the encoded training pair $\big{(}\mathbf{z}_{0},\mathbf{z}_{T}\big{)}$, the expectation of data log-likelihood possesses the following ELBO formulation:
 […]
-where $q(\mathbf{z}_{t-1}|\mathbf{z}_{t},\mathbf{z}_{0},\mathbf{z}_{T})$ and $p_{\theta}(\mathbf{z}_{t-1}|\mathbf{z}_{t},\mathbf{z}_{T})$ are the transition kernel of true reverse process and our modeled generative process and $\mathbb{D}_{KL}$ is the KL divergence.
+$\sum_{t=2}^{T}\mathbb{E}_{p(\mathbf{z}_{t}|\mathbf{z}_{0})}$ $\big{[}\mathbb{D}_{KL}\big{(}q(\mathbf{z}_{t-1}\mid\mathbf{z}_{t},\mathbf{z}_{0},\mathbf{z}_{T})\,||\,p_{\theta}(\mathbf{z}_{t-1}\mid\mathbf{z}_{t},\mathbf{z}_{T})\big{)}\big{]}\Big{]}$
+where $q(\mathbf{z}_{t-1}|\mathbf{z}_{t},\mathbf{z}_{0},\mathbf{z}_{T})$ and $p_{\theta}(\mathbf{z}_{t-1}|\mathbf{z}_{t},\mathbf{z}_{T})$ are the transition kernel of true reverse process and our modeled generative process.
+$\mathbb{D}_{KL}$ is the KL divergence.
 Utilizing Bayes’ theorem and Markov chain property, we can obtain:
@@
 $=\mathcal{N}(\tilde{\boldsymbol{\mu}}_{t}(\mathbf{z}_{t},\mathbf{z}_{0},\mathbf{z}_{T}),\tilde{\sigma}_{t}^{2}\mathbf{I})$
-For simplicity, we define the following notations: $a_{t}=m_{t}/m_{t-1},b_{t}=n_{t}-a_{t}\cdot n_{t-1},\delta_{t}^{2}=\sigma_{t}^{2}-a_{t}^{2}\cdot\sigma_{t-1}^{2}$. The mean value term is:
+For simplicity, we define the following notations: $a_{t}=m_{t}/m_{t-1},b_{t}=n_{t}-a_{t}\cdot n_{t-1},\delta_{t}^{2}=\sigma_{t}^{2}-a_{t}^{2}\cdot\sigma_{t-1}^{2}$.
+Based on Equation 3, we further derive the mean value term:
 $\tilde{\boldsymbol{\mu}}_{t}(\mathbf{z}_{t},\mathbf{z}_{0},\mathbf{z}_{T})$ $=k_{1}\mathbf{z}_{t}+k_{2}\mathbf{z}_{0}+k_{3}\mathbf{z}_{T}$ (6)
-$k_{1}$ $=\frac{\sigma_{t-1}^{2}}{\sigma_{t}^{2}}\cdot a_{t}$
-$k_{2}$ $=\frac{\delta_{t}^{2}}{\sigma_{t}^{2}}\cdot m_{t-1}$
-$k_{3}$ $=\frac{\delta_{t}^{2}}{\sigma_{t}^{2}}\cdot n_{t-1}-\frac{\sigma_{t-1}^{2}}{\sigma_{t}^{2}}\cdot a_{t}\cdot b_{t},$
-and the variance is:
+$k_{1}=\frac{\sigma_{t-1}^{2}}{\sigma_{t}^{2}}\cdot a_{t},\quad k_{2}=\frac{\delta_{t}^{2}}{\sigma_{t}^{2}}\cdot$ $m_{t-1},\quad k_{3}=\frac{\delta_{t}^{2}}{\sigma_{t}^{2}}\cdot n_{t-1}-\frac{\sigma_{t-1}^{2}}{\sigma_{t}^{2}}\cdot a_{t}\cdot b_{t},$
+and the variance term:
 $\tilde{\sigma}_{t}^{2}=\frac{\sigma_{t-1}^{2}}{\sigma_{t}^{2}}\cdot\delta_{t}^{2}$ (7)
-Assume $p_{\theta}(\mathbf{z}_{t-1}|\mathbf{z}_{t},\mathbf{z}_{T})$ is also a Gaussian distribution parameterized as $\mathcal{N}(\boldsymbol{\mu}_{\theta}(\mathbf{z}_{t},\mathbf{z}_{T},t),\boldsymbol{\Sigma}_{\theta}(\mathbf{z}_{t},\mathbf{z}_{T},t))$ and let $\boldsymbol{\Sigma}_{\theta}(\mathbf{z}_{t},\mathbf{z}_{T},t))=\tilde{\sigma}_{t}^{2}$ for derivation simplicity, maximizing the ELBO is equivalent to minimizing:
+Assuming $p_{\theta}(\mathbf{z}_{t-1}|\mathbf{z}_{t},\mathbf{z}_{T})\sim\mathcal{N}(\boldsymbol{\mu}_{\theta}(\mathbf{z}_{t},\mathbf{z}_{T},t),\boldsymbol{\Sigma}_{\theta}(\mathbf{z}_{t},\mathbf{z}_{T},t))$ is also a Gaussian distribution with fixed variance $\boldsymbol{\Sigma}_{\theta}(\mathbf{z}_{t},\mathbf{z}_{T},t)=\tilde{\sigma}_{t}^{2}$ for simplicity of derivation.
+Following the strategy of DDPM [20], we reparameterize the model $\boldsymbol{\epsilon}_{\theta}(\mathbf{z}_{t},\mathbf{z}_{T},t)$ to predict noise instead of posterior mean.
+Under this formulation, the ELBO objective is equivalent to:
 $\mathcal{L}_{elbo}$ $=\mathbb{E}_{t,\mathbf{z}_{0},\mathbf{z}_{t},\mathbf{z}_{T}}\left[\frac{1}{2\tilde{\sigma}_{t}^{2}}\left\|\tilde{\boldsymbol{\mu}}_{t}-\boldsymbol{\mu}_{\theta}(\mathbf{z}_{t},\mathbf{z}_{T},t)\right\|^{2}\right]$ (8)
-Similar to the strategy in DDPM (Ho et al., 2020), we use a network $\boldsymbol{\epsilon}_{\theta}(\mathbf{z}_{t},\mathbf{z}_{T},t)$ to parameterize noise instead of posterior mean and further simplified the loss as:
-$\mathcal{L}_{elbo}=\mathbb{E}_{t,\mathbf{z}_{0},\mathbf{z}_{t},\mathbf{z}_{T}}\left[\left\|\boldsymbol{\epsilon}-\boldsymbol{\epsilon}_{\theta}(\mathbf{z}_{t},\mathbf{z}_{T},t)\right\|^{2}\right],$ (9)
-4.4.2 Image Consistency Loss
-In addition to the likelihood loss evaluated in latent space, we also incorporate an image-level consistency loss, where we first convert the latent prediction back to image space, and then evaluate its consistency with the ground truth dense map using a simple MSE loss.
+$\cong\mathbb{E}_{t,\mathbf{z}_{0},\mathbf{z}_{t},\mathbf{z}_{T}}\left[\left\|\boldsymbol{\epsilon}-\boldsymbol{\epsilon}_{\theta}(\mathbf{z}_{t},\mathbf{z}_{T},t)\right\|^{2}\right]$
+Image Consistency Loss.
+In addition to the latent-space likelihood loss, we also incorporate an image-level consistency loss, where we first decode the latent prediction back to image space, and then evaluate its consistency with the ground truth dense map using a simple MSE.
 $\mathcal{L}_{ic}=\left\|\mathcal{D}(\hat{\mathbf{z}}_{0})-\mathbf{y})\right\|^{2}$ (9)
-The reason is that the compression of input data in the latent space will lead to a loss of high-frequency information, harming the model’s ability of preserving fine-grained spatial details.
-By introducing an auxiliary loss at image level, we ensure that the model is capable of retaining spatial information which is essential for precise per-pixel prediction.
-This dual loss strategy leverages the compactness of latent space for efficiency while still anchoring the training process to the fidelity of full-resolution dense map prediction, thus leading to performance improvement.
-Combined these two losses together, our final training objective becomes:
+This additional supervision addresses the loss of high-frequency structures caused by latent-space compression, which can impair the model’s ability to preserve fine-grained details.
+By introducing this auxiliary image-space loss, we encourage the model to retain spatial information critical for accurate per-pixel prediction.
+This dual-loss strategy combines the efficiency of latent-space modeling with the fidelity of image-space supervision, leading to improved performance.
+The final training objective is given by:
 $\mathcal{L}=\omega_{1}\mathcal{L}_{elbo}+\omega_{2}\mathcal{L}_{ic}$ (10)
 where $\omega_{1},\omega_{2}$ are hyperparameters to balance the loss.
-Algorithm 1 Training
-0:  Encoder $\mathcal{E}$; Decoder $\mathcal{D}$; Loss weights $\omega_{1},\omega_{2}$
-1:  repeat
-2:     Retrieve paired data $(\mathbf{x},\mathbf{y})$
-3:     $\mathbf{z}_{0}=\mathcal{E}(\mathbf{y})$, $\mathbf{z}_{T}=\mathcal{E}(\mathbf{x})$
-4:     $t\sim\text{Uniform}(\{1,\dots,T\})$, $\boldsymbol{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$
-5:     $\mathbf{z}_{t}=m_{t}\mathbf{z}_{0}+n_{t}\mathbf{z}_{T}+\sigma_{t}\boldsymbol{\epsilon}$
-6:     $\mathbf{z}_{t}^{\prime}=\frac{1}{\sqrt{n_{t}^{2}+\sigma_{t}^{2}}}(\mathbf{z}_{t}-n_{t}\mathbf{z}_{T})$
-7:     $\mathcal{L}_{elbo}=\left\|\boldsymbol{\epsilon}-\boldsymbol{\epsilon}_{\theta}(\mathbf{z}_{t}^{\prime},\mathbf{z}_{T},t)\right\|^{2},$
-8:     $\hat{\mathbf{z}}_{0}=\frac{1}{m_{t}}(\mathbf{z}_{t}-n_{t}\mathbf{z}_{T}-\sigma_{t}\boldsymbol{\epsilon}_{\theta}(\mathbf{z}_{t}^{\prime},\mathbf{z}_{T},t))$
-9:     $\mathcal{L}_{ic}=\left\|\mathcal{D}(\hat{\mathbf{z}}_{0})-\mathbf{y}\right\|^{2}$
-10:     Take gradient step on $\nabla_{\theta}\left(\omega_{1}\mathcal{L}_{elbo}+\omega_{2}\mathcal{L}_{ic}\right)$
-11:  until Converged
-Algorithm 2 Inference
-0:  Encoder $\mathcal{E}$; Decoder $\mathcal{D}$; Trained $\boldsymbol{\epsilon}_{\theta}(\cdot,\cdot,\cdot)$
-1:  $\mathbf{z}_{T}=\mathcal{E}(\mathbf{x})$
-2:  for $t=T,\dots,1$ do
-3:     $\boldsymbol{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ if $t>1$, else $\boldsymbol{\epsilon}=\mathbf{0}$
-4:     $\mathbf{z}_{t}^{\prime}=\frac{1}{\sqrt{n_{t}^{2}+\sigma_{t}^{2}}}(\mathbf{z}_{t}-n_{t}\mathbf{z}_{T})$
-5:     $\hat{\mathbf{z}}_{0}=\frac{1}{m_{t}}(\mathbf{z}_{t}-n_{t}\mathbf{z}_{T}-\sigma_{t}\boldsymbol{\epsilon}_{\theta}(\mathbf{z}_{t}^{\prime},\mathbf{z}_{T},t))$
-6:     $\mathbf{z}_{t-1}=k_{1}\mathbf{z}_{t}+k_{2}\hat{\mathbf{z}}_{0}+k_{3}\mathbf{z}_{T}+\tilde{\sigma}_{t}\boldsymbol{\epsilon}$
-7:  end for
-8:  $\mathbf{y}=\mathcal{D}(\hat{\mathbf{z}}_{0})$
-9:  return $\mathbf{y}$
 4.5 Distribution Alignment Normalization
-A notable challenge during finetuning is the misalignment of intermediate sample’s marginal distribution $q(\mathbf{z}_{t})$ between DPBridge and diffusion backbone, due to the difference in the forward transition kernel.
-This misalignment will weaken the model’s inherent denoising ability, thus limiting the effectiveness of knowledge transfer.
-To address this issue, we propose a distribution alignment normalization technique to fill in the gap between the two distributions.
-Specifically, we normalize the intermediate latents computed from Equation 3 to match the form of a standard diffusion process:
+A key challenge during finetuning is the misalignment between the marginal distribution $q(\mathbf{z}_{t})$ of intermediate samples in DPBridge and that of the diffusion backbone, due to the difference in the forward transition kernels.
+This misalignment weakens the model’s denoising ability and hinders effective knowledge transfer.
+To address this issue, we propose a distribution alignment normalization technique that transforms intermediate latents computed from our bridge process Equation 3 to match the distributional form of a standard diffusion process.
 $\mathbf{z}_{t}^{\prime}=\frac{\mathbf{z}_{t}-m_{t}\mathbf{z}_{T}}{\sqrt{n_{t}^{2}+\sigma_{t}^{2}}}=\frac{n_{t}}{\sqrt{n_{t}^{2}+\sigma_{t}^{2}}}\mathbf{z}_{0}+\frac{\sigma_{t}}{\sqrt{n_{t}^{2}+\sigma_{t}^{2}}}\boldsymbol{\epsilon}_{t}$ (11)
@@
 Therefore by feeding the normalized sample $\mathbf{z}_{t}^{\prime}$ in the network $\boldsymbol{\epsilon}_{\theta}$, we can fully exploit the denoising capability from pretrained diffusion backbone.
-An overview of DPBridge’s training pipeline shown in Algorithm 1.
-4.6 Accelerate Inference
-Similar to the idea of DDIM (Song et al., 2020a), the inference process of DPBridge can be accelerated by utilizing a non-Markovian process while maintaining the same marginal distributions for all intermediate samples along the sampling path.
-Specifically, we can sample on a subset of timesteps $\{t_{i}\}_{i=1}^{N}$, where $0=t_{0}<t_{1}<\cdots<t_{N-1}<t_{N}=T$ and the number of sampling steps $N$ can be set independently from the original timesteps $T$ on which the bridge model is trained.
-The accelerated inference process can be thus described as:
-$\begin{split}\mathbf{z}_{t_{n}}&=m_{t_{n}}\hat{\mathbf{z}}_{0}+n_{t_{n}}\mathbf{z}_{T}+\\ &\sqrt{\sigma_{t_{n}}^{2}-g_{n}^{2}}\frac{\mathbf{z}_{t_{n+1}}-m_{t_{n+1}}\hat{\mathbf{z}}_{0}-n_{t_{n+1}}\mathbf{z}_{T}}{\sigma_{t_{n+1}}}+g_{n}\boldsymbol{\epsilon}\end{split}$ (13)
-where $g_{n}$ is the variance parameter that controls the stochasticity behavior of the reverse sampling process.
-We note that when $g_{n}=\tilde{\sigma}_{t}^{2}$ which is calculated from Equation 5, the process becomes a Markovian bridge.
-The basic inference process of DPBridge is summarized in Algorithm 2.
-If adopt acceleration, one can utilize Equation 13 during implementation.
+We note that a similar strategy was independently proposed in FrameBridge [52], where it is applied in the context of image-to-video generation.
+Our method, developed concurrently, is specifically designed to facilitate adaptation in dense prediction tasks under the DPBridge framework.
+An overview of the full training and inference procedures are summarized in Algorithm 1 Algorithm 2.
+For more derivation details, please refer to Appendix A.
 5 Experiments
-We conduct evaluation mainly on two representative dense prediction tasks, i.e., depth estimation and surface normal prediction.
-We’ll start by a brief introduction of experiment settings as well as implementation details, followed by quantitative and qualitative analysis.
-Lastly, we conduct detailed ablation studies and investigate the influence each design component respectively.
+We conduct evaluation primarily on two representative dense prediction tasks: depth estimation and surface normal prediction.
+We begin with a brief overview of experimental settings, followed by both quantitative and qualitative analyses.
+Next, we present ablation studies to investigate the contribution of each design component.
+Implementation details are presented in Appendix D for reference.
 5.1 Experiments Setup
@@
 The basic setting for depth estimation and surface normal prediction experiments are similar.
-We use two synthetic datasets, HyperSim (Roberts et al., 2021) and Virtual KITTI (Cabon et al., 2020), which cover both indoor and outdoor scenes, for training.
+We use two synthetic datasets, HyperSim [40] and Virtual KITTI [6], for training, which cover both indoor and outdoor scenes with high-quality ground-truth annotations.
 To comprehensively evaluate the model performance and demonstrate its generalization capability, we conduct zero-shot evaluations on real-world datasets.
-For depth estimation, this includes NYUv2 (Silberman et al., 2012), KITTI (Geiger et al., 2012), ScanNet (Dai et al., 2017), ETH3D (Schops et al., 2017) and DIODE (Vasiljevic et al., 2019).
-For surface normal prediction, we additionally incorporate iBims (Koch et al., 2018) and Sintel (Butler et al., 2012) datasets.
-For fair comparison, we use the same validation split as in previous methods (Ke et al., 2024).
+For depth estimation, this includes NYUv2 [44], KITTI [18], ScanNet [13], ETH3D [43] and DIODE [50].
+For surface normal prediction, we additionally incorporate iBims [29] and Sintel [5] datasets.
+For fair comparison, we use the same validation split as in previous methods [28].
 Evaluation Metrics.
@@
 For the metrics of surface normal task, we report the commonly used mean angular error between the ground-truth normal vectors and the prediction results, as well as the percentage of pixels with an angular error below 11.25 degrees.
-5.2 Implementation Details
-We implement DPBridge based on Stable Diffusion v1.5 (SD1.5) and v2.1-base (SD2.1) (Rombach et al., 2022) and report their results respectively to shown compatibility with different backbones.
-During training, the number of timesteps of is set to be 1000, and we use 50 sampling steps during inference with considerations of both quality and efficiency.
-Training our method takes 30K iterations using the Adam optimizer with a base learning rate of $3\times 10^{-5}$ and an exponential decay strategy after 100-step warm-up.
-The batch size is 2, but with 8 steps gradient accumulation, the effective batch size is equivalent to 16.
-We also apply random horizontal flipping augmentation during training.
-The whole process takes approximately 3 days to complete on a single Nvidia RTX 3090Ti GPU.
 Method Model Type NYUv2 KITTI ETH3D ScanNet DIODE Avg.
 […]
-DiverseDepth (Yin et al., 2020) 11.7 87.5 19.0 70.4 22.8 69.4 10.9 88.2 37.6 63.1 12.4
-MiDaS (Ranftl et al., 2020) 11.1 88.5 23.6 63.0 18.4 75.2 12.1 84.6 33.2 71.5 11.8
-LeReS (Yin et al., 2021) 9.0 91.6 14.9 78.4 17.1 77.7 9.1 91.1 27.1 76.6 9.5
-Omnidata (Eftekhar et al., 2021) 7.4 94.5 14.9 83.5 16.6 77.8 7.5 93.2 33.9 74.2 9.2
-HDN (Zhang et al., 2022) 6.9 94.8 11.5 86.7 12.1 83.3 8.0 93.6 24.6 78.0 7.2
-DPT (Ranftl et al., 2021) 9.8 90.3 10.0 90.1 7.8 94.6 8.2 93.8 18.2 75.8 6.9
-Depth Anything V2 (Yang et al., 2024) 4.4 97.9 7.5 94.8 13.2 84.2 4.4 98.0 6.5 95.4 2.3
-Marigold (Ke et al., 2024) 6.0 95.9 10.5 90.4 7.1 95.1 6.9 94.5 31.0 71.2 5.8
-DepthFM (Gui et al., 2024) 6.5 95.6 8.3 93.4 - - - - 22.5 80.0 3.8
-Geowizard (Fu et al., 2025) 5.7 96.2 14.4 82.0 7.5 94.3 6.1 95.8 31.4 77.1 6.1
-E2E-FT (Garcia et al., 2024) 5.4 96.4 9.6 92.1 6.4 95.9 5.8 96.5 30.3 77.6 3.4
-I2SB (Liu et al., 2023) 24.9 73.4 37.7 57.9 40.6 52.0 21.2 66.1 40.2 59.1 13.9
-BBDM (Li et al., 2023) 27.9 65.5 47.2 45.1 46.4 48.8 22.7 67.0 34.6 64.7 14.1
-DPBridge (SD1.5) 6.9 95.7 11.2 87.3 8.1 93.2 7.5 94.7 29.7 78.6 6.1
-DPBridge (SD2.1) 4.4 97.6 9.8 91.6 6.9 94.9 4.5 97.4 23.8 84.0 2.6
-Table 1: Quantitative comparison of DPBridge with state-of-the-art affine-invariant depth estimation methods on zero-shot benchmarks.
+DiverseDepth [58] 11.7 87.5 19.0 70.4 22.8 69.4 10.9 88.2 37.6 63.1 12.2
+MiDaS [38] 11.1 88.5 23.6 63.0 18.4 75.2 12.1 84.6 33.2 71.5 11.8
+LeReS [59] 9.0 91.6 14.9 78.4 17.1 77.7 9.1 91.1 27.1 76.6 9.5
+Omnidata [15] 7.4 94.5 14.9 83.5 16.6 77.8 7.5 93.2 33.9 74.2 9.2
+HDN [61] 6.9 94.8 11.5 86.7 12.1 83.3 8.0 93.6 24.6 78.0 7.2
+DPT [39] 9.8 90.3 10.0 90.1 7.8 94.6 8.2 93.8 18.2 75.8 7.0
+Depth Anything V2 [56] 4.4 97.9 7.5 94.8 13.2 84.2 4.4 98.0 6.5 95.4 2.3
+Marigold [28] 6.0 95.9 10.5 90.4 7.1 95.1 6.9 94.5 31.0 71.2 6.0
+DepthFM [19] 6.5 95.6 8.3 93.4 - - - - 22.5 80.0 3.8
+Geowizard [16] 5.7 96.2 14.4 82.0 7.5 94.3 6.1 95.8 31.4 77.1 6.2
+E2E-FT [17] 5.4 96.4 9.6 92.1 6.4 95.9 5.8 96.5 30.3 77.6 3.6
+I2SB [32] 23.5 63.6 35.1 46.4 53.4 40.5 21.1 67.4 55.5 58.2 14.0
+BBDM [31] 33.2 49.6 58.6 25.6 52.0 40.6 25.2 61.2 70.1 46.9 15.2
+DDBM [65] 26.4 62.2 55.1 27.6 69.5 33.8 23.2 65.6 52.3 56.9 14.5
+DPBridge (SD1.5) 6.9 95.7 11.2 87.3 8.1 93.2 7.5 94.7 29.7 78.6 5.4
+DPBridge (SD2.1) 4.4 97.6 9.8 91.6 6.9 94.9 4.5 97.4 23.8 84.0 2.7
+Table 1: Quantitative comparison on affine-invariant depth estimation benchmarks.
 All metrics are presented in percentage terms.
@@
 […]
-EESNU (Bae et al., 2021) 16.2 58.6 - - 20.0 58.5 42.1 11.5 6.5
-Omnidata v1 (Eftekhar et al., 2021) 23.1 45.8 22.9 47.4 19.0 62.1 41.5 11.4 8.1
-Omnidata v2 (Kar et al., 2022) 17.2 55.5 16.2 60.2 18.2 63.9 40.5 14.7 5.6
-DSINE (Bae & Davison, 2024) 16.4 59.6 16.2 61.0 17.1 67.4 34.9 21.5 3.2
-Marigold (Ke et al., 2024) 18.8 55.9 17.7 58.8 18.4 64.3 39.1 14.9 5.8
-Geowizard (Fu et al., 2025) 17.0 56.5 15.4 61.6 13.0 65.3 40.4 13.2 4.1
-E2E-FT (Garcia et al., 2024) 16.5 60.4 14.7 66.1 16.1 69.7 33.5 22.3 2.1
-DPBridge (SD1.5) 18.1 54.2 18.0 54.5 24.3 51.2 41.2 19.4 7.3
+EESNU [3] 16.2 58.6 - - 20.0 58.5 42.1 11.5 6.8
+Omnidata v1 [15] 23.1 45.8 22.9 47.4 19.0 62.1 41.5 11.4 8.8
+Omnidata v2 [27] 17.2 55.5 16.2 60.2 18.2 63.9 40.5 14.7 6.2
+DSINE [2] 16.4 59.6 16.2 61.0 17.1 67.4 34.9 21.5 3.5
+Marigold [28] 18.8 55.9 17.7 58.8 18.4 64.3 39.1 14.9 6.4
+Geowizard [16] 17.0 56.5 15.4 61.6 13.0 65.3 40.4 13.2 4.2
+E2E-FT [17] 16.5 60.4 14.7 66.1 16.1 69.7 33.5 22.3 2.1
+StableNormal [57] 17.7 55.9 15.8 61.2 17.2 66.6 - - 5.0
+DPBridge (SD1.5) 18.1 54.2 18.0 54.5 24.3 51.2 41.2 19.4 8.0
 DPBridge (SD2.1) 14.8 65.3 13.4 68.5 17.1 68.0 25.0 46.3 1.4
-Table 2: Quantitative comparison of DPBridge with state-of-the-art surface normal prediction methods on zero-shot benchmarks.
+Table 2: Quantitative comparison on surface normal prediction benchmarks.
 All metrics are presented in percentage terms.
@@
 The best and second-best results are highlighted in bold and underline formats.
+Input Image
+DPBridge
+Marigold
+Geowizard
+DepthFM
+Figure 4: Qualitative comparison with diffusion-based depth estimation methods on in-the-wild images. DPBridge can generate depth maps with sharper boundaries and more coherent scene geometry.
+Input Image
+DPBridge
+Marigold
+Geowizard
+StableNormal
+Figure 5: Qualitative comparison with diffusion-based surface normal prediction methods on in-the-wild images.
+DPBridge’s results exhibit more fine-grained geometric details, such as the curvature of architectural structures, the contour of coral and reef, and the fur texture of the cat.
 5.2 Quantitative and Qualitative Results
-Table 1 and Table 2 summarize the quantitative metrics of DPBridge compared with state-of-the-art (SOTA) methods from both discriminative and generative fields, all in zero-shot manner.
-In depth estimation, DPBridge consistently matches SOTA results, while demonstrating robust generalization across datasets.
-It achieves 4.4 AbsRel, 97.2 % $\delta_{1}$ on NYUv2, and 4.5 AbsRel, 97.4% $\delta_{1}$ on ScanNet.
-Similarly, for surface normal prediction, DPBridge also delivers competitive results, achieving mean error of 14.8 and $11.25^{\circ}$ accuracy of $65.3\%$ on NYUv2, as well as maintaining strong performance on other datasets like iBims-1 (mean: 17.1, $11.25^{\circ}:68.0\%$) and Sintel (mean: 25.0, $11.25^{\circ}:46.3\%$).
-These results highlight DPBridge’s versatility to generalize well across tasks and benchmarks, being comparable to both feed-forward and diffusion-based methods while offering a unified, efficient paradigm to handle dense prediction tasks.
-Notably, when compared with bridge models like I2SB and BBDM that are trained from scratch, we exceed their performance by a large margin, justifying our proposal that adapting prior knowledge to bridge construction indeed benefits the perception performance.
-In terms of diffusion-based methods, DPBridge can also achieve the best result, showcasing that adopting bridge process to handle dense prediction tasks is indeed advantageous over standard diffusion framework.
-We further provide qualitative examples for both tasks in comparison with SOTA diffusion-based methods (Ke et al., 2024; Fu et al., 2025; Garcia et al., 2024), which are most closely related to our work.
-Figure 3 shows that E2E-FT tends to produce blurry and smooth predictions, while Geowizard fails to provide accurate estimation in relatively long distance range.
-In contrast, DPBridge perform well in both indoor and outdoor scenes, capturing fine-level details in both near and faraway regions, like the contour of leaves, fences, cars, computers or windows.
-Figure 4 presents visualizations on in-the-wild images for surface normal prediction task, where we can see that DPBridge is able to predict high-quality surface normals equipped with fine-grained textures.
-Figure 3: Qualitative comparison of diffusion-based depth estimation methods on KITTI and ETH3D dataset.
-DPBridge is able to generalize in both indoor and outdoor scenes, successively capturing fine-grained details such as leaves and structures on the fence (KITTI).
-Figure 4: Qualitative comparison between DPBridge (above) and Marigold (below) on normal estimation results, DPBridge is able to capture more textures from the photos in-the-wild.
+Table 1 and Table 2 summarize the quantitative results of DPBridge on zero-shot benchmarks across depth estimation and surface normal prediction tasks.
+We evaluate two variants of our model—DPBridge (SD1.5) and DPBridge (SD2.1)—built on top of Stable Diffusion v1.5 and v2.1-base [41], respectively.
+For depth estimation, DPBridge (SD2.1) achieves an AbsRel of 4.4 and $\delta_{1}$ accuracy of 97.6% on NYUv2, and 4.5 AbsRel / 97.4% $\delta_{1}$ on ScanNet, both among the best in class.
+Across all five datasets, it achieves an average ranking of 2.7—second only to Depth Anything V2 (rank 2.3), which is trained on over 1,000 times more data.
+For surface normal prediction, DPBridge also delivers state-of-the-art performance with a mean angular error of 14.8° and $65.3\%$ accuracy under $11.25^{\circ}$ on NYUv2, outperforming both feed-forward and diffusion-based methods.
+It also maintains strong results on iBims-1 ($17.1$, $68.0\%$) and Sintel ($25.0$, $46.3\%$).
+These results highlight DPBridge’s versatility to generalize well across tasks and benchmarks.
+Notably, compared to prior bridge models such as I2SB, BBDM, and DDBM, DPBridge reduces average error by a large margin (e.g., 4.4 vs.
+23.5 AbsRel in NYUv2 depth), underscoring the benefits and importance of incorporating pretrained visual priors to bridge construction.
+Furthermore, among diffusion-based approaches, DPBridge consistently achieves the best or second-best scores, validating our core hypothesis that bridge formulation provides clear advantages over standard diffusion frameworks in dense prediction tasks, and combining bridge modeling with pretrained visual prior offers a more efficient and generalizable framework.
+Figure 5 and Figure 5 present qualitative comparisons of DPBridge with SOTA diffusion-based methods on in-the-wild images for depth estimation and surface normal prediction, respectively.
+As shown in Figure 5, DPBridge generates depth maps with sharper object boundaries and more coherent scene geometry, particularly in challenging cases such as urban scenes or complex foreground objects like animals, grass and flowers.
+In contrast, competing methods like Geowizard and DepthFM often produce over-smoothed outputs or fail to preserve depth discontinuities.
+In Figure 5, DPBridge demonstrates stronger fidelity in fine-grained geometric details, such as the curvature of architectural structures, the contour of coral and reef, and the fur texture of the cat.
+Competing methods like Marigold tends to focus on the center object in the image and ignore the environment, while StableNormal produces unreliable outputs.
+These visual results reinforce the quantitative findings and highlight the effectiveness of DPBridge in modeling structured output distributions with both global coherence and fine details.
+Figure 6: Ablation of sampling steps. The performance improves as the number of steps increases, while observing saturation after 10 steps.
 Design Components NYUv2 KITTI
@@
 6.2 95.8 11.0 87.5
-5.0 96.3 10.3 90.1
+4.7 97.1 10.0 91.0
 4.4 97.6 9.8 91.6
@@
 IC indicates image consistency loss.
-Figure 5: Ablation of sampling steps. The performance improves as the number of sampling steps increases, while we observe saturation after 10 steps.
 5.3 Ablation Studies
@@
 The results are summarized in Table 3 and Figure 6, with detailed discussions provided below.
-Finetuning VS Training From Scratch.
-To assess whether adapting pretrained diffusion backbones improves perception performance, we conduct comparison between finetuning and training from scratch paradigm.
-As shown in the first row of Table 3, fine-tuning significantly outperforms training from scratch under identical experimental settings.
-This difference is due to two factors: first, foundation models trained on large, diverse datasets.
-They capture universal visual priors, aiding knowledge transfer to dense prediction tasks.
-Second, training from scratch requires more iterations for convergence.
-In addition, from the last two rows of Table 1 and Table 2, we can see that finetuning on SD2.1 exceeds that on SD1.5, suggesting that the quality of prior knowledge will also influence the performance.
-These observations indicate that building DPBridge on superior pretrained backbones is indeed advantageous.
+Visual Prior Exploitation.
+To evaluate the usage of visual priors, we compare finetuning with training from scratch.
+As shown in the first row of Table 3, finetuning yields significantly better results under identical settings.
+This is attributed to stronger visual priors from foundation models and faster convergence.
+Additionally, the last rows of Table 1 and Table 2 show that SD2.1 outperforms SD1.5, highlighting that higher-quality priors further boost performance.
+These results confirm the advantage of building DPBridge on strong pretrained backbones.
 Distribution Alignment Normalization.
-Next, we evaluate whether aligning the marginal distributions between diffusion bridge process and standard diffusion process is beneficial.
-As shown in the second and fourth rows of Table 3, incorporating a distribution alignment normalization step before inputting data into the noise prediction network noticeably improves performance.
+We next assess the benefit of aligning marginal distributions between the diffusion bridge and standard diffusion processes.
+As shown in the second and fourth rows of Table 3, adding a distribution alignment normalization step before the noise prediction network consistently improves performance.
 Image Consistency Loss.
-Regarding the auxiliary image consistency loss, the comparison between the third and fourth rows in Table 3 shows that while effective, its impact is less significant than that of distribution alignment normalization.
-This is probably because it primarily influences regions where masked and valid areas overlap, which constitute only a relatively small proportion of the data.
-However, when integrated with other design components, the overall performance is significantly improved.
+As shown in the third and fourth rows of Table 3, the auxiliary image consistency loss improves performance but has a smaller impact than distribution alignment.
+This is likely because it mainly affects the boundaries or overlapping regions between masked and valid areas, which are limited in scope.
+Nonetheless, when combined with other components, it contributes to a notable overall performance gain.
 Number of Sampling Steps.
-We evaluate performance with varying numbers of sampling steps during inference.
-As shown in Figure 5, increasing sampling steps initially improves perception performance rapidly and reach saturation after 10 steps.
-Notably, compared to the results in Table 1, where all other diffusion-based methods are evaluated with 50 steps setting, DPBridge achieves comparable performance with fewer steps, demonstrating the efficiency of our proposed method.
+We evaluate performance under varying sampling steps during inference.
+As shown in Figure 6, perception quality improves rapidly with more steps and saturates after 10.
+Notably, compared to Table 1, where other diffusion-based methods use 50 steps, DPBridge achieves comparable results with significantly fewer steps, highlighting its efficiency.
 6 Conclusion
-In this paper, we present DPBridge, a latent diffusion bridge model to tackle general dense prediction tasks through generative pipeline.
-We establish direct mapping between input images and corresponding signal maps in the latent space based on fully-tractable diffusion bridge process.
-In this way, the prior distribution is no longer constrained to uninformative noise, but instead can be chosen as more structurally correlated ones.
-In addition, we introduce effective finetuning strategies to adapt large pretrained diffusion backbone to our bridge model construction, leveraging its rich prior knowledge to boost training speed and achieve remarkable perception performance.
-Future research directions include deriving other parameterization of diffusion bridge process, and extending to more dense prediction tasks.
-Impact Statement
-This paper presents work whose goal is to advance the field of Machine Learning and we wish to spark further research on image generation and perception.
-There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
+In this paper, we propose DPBridge, a latent diffusion bridge framework for dense prediction tasks through integrating the strengths of diffusion bridge modeling with visual priors from pretrained foundation models.
+We derive a fully tractable bridge process to establish direct probability transport in the latent space, and propose finetuning strategies to better adapt pretrained diffusion backbones in bridge construction.
+While DPBridge achieves competitive performance across different tasks and benchmarks, several limitations remain.
+First, our method still requires iterative sampling, which makes it slower than feed-forward approaches.
+Exploring acceleration strategies such as distillation or single-step bridge solvers could further improve practical deployment, especially in real-time settings.
+Second, our formulation is restricted to certain bridge parameterization, and exploring alternatives to adapt to more advanced pretrained backbone is a promising future direction.
 Appendix A Derivation Details
-In this section, we’ll provide derivation details of the maximum likelihood loss presented in Section 4.4.1.
+In this section, we’ll provide derivation details of the maximum likelihood loss presented in Section 4.4.
 Its core is the posterior distribution $q(\mathbf{z}_{t-1}|\mathbf{z}_{t},\mathbf{z}_{0},\mathbf{z}_{T})$ as in Equation 5, whose calculation involves three terms: $q(\mathbf{z}_{t}|\mathbf{z}_{t-1},\mathbf{z}_{T})$, $q(\mathbf{z}_{t-1}|\mathbf{z}_{0},\mathbf{z}_{T})$, and $q(\mathbf{z}_{t}|\mathbf{z}_{0},z\mathbf{z}_{T})$.
@@
 $=\mathbb{E}_{t,\mathbf{z}_{0},\mathbf{z}_{t},\mathbf{z}_{T}}\left[\left\|\boldsymbol{\epsilon}-\boldsymbol{\epsilon}_{\theta}(\mathbf{z}_{t},\mathbf{z}_{T},t)\right\|^{2}\right]$
-which is the same as in Equation 8 and Equation 9. This concludes our derivation.
-Appendix B Experiments
-B.1 Network Architecture
+which is the same as in Equation 8. This concludes our derivation.
+Appendix B Training and Inference Pipeline
+B.1 Pipeline Pseudo-code
+Algorithm 1 Training
+0:  Encoder $\mathcal{E}$; Decoder $\mathcal{D}$; Loss weights $\omega_{1},\omega_{2}$
+1:  repeat
+2:     Retrieve paired data $(\mathbf{x},\mathbf{y})$
+3:     $\mathbf{z}_{0}=\mathcal{E}(\mathbf{y})$, $\mathbf{z}_{T}=\mathcal{E}(\mathbf{x})$
+4:     $t\sim\text{Uniform}(\{1,\dots,T\})$, $\boldsymbol{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$
+5:     $\mathbf{z}_{t}=m_{t}\mathbf{z}_{0}+n_{t}\mathbf{z}_{T}+\sigma_{t}\boldsymbol{\epsilon}$
+6:     $\mathbf{z}_{t}^{\prime}=\frac{1}{\sqrt{n_{t}^{2}+\sigma_{t}^{2}}}(\mathbf{z}_{t}-n_{t}\mathbf{z}_{T})$
+7:     $\mathcal{L}_{elbo}=\left\|\boldsymbol{\epsilon}-\boldsymbol{\epsilon}_{\theta}(\mathbf{z}_{t}^{\prime},\mathbf{z}_{T},t)\right\|^{2},$
+8:     $\hat{\mathbf{z}}_{0}=\frac{1}{m_{t}}(\mathbf{z}_{t}-n_{t}\mathbf{z}_{T}-\sigma_{t}\boldsymbol{\epsilon}_{\theta}(\mathbf{z}_{t}^{\prime},\mathbf{z}_{T},t))$
+9:     $\mathcal{L}_{ic}=\left\|\mathcal{D}(\hat{\mathbf{z}}_{0})-\mathbf{y}\right\|^{2}$
+10:     Take gradient step on $\nabla_{\theta}\left(\omega_{1}\mathcal{L}_{elbo}+\omega_{2}\mathcal{L}_{ic}\right)$
+11:  until Converged
+Algorithm 2 Inference
+0:  Encoder $\mathcal{E}$; Decoder $\mathcal{D}$; Trained $\boldsymbol{\epsilon}_{\theta}(\cdot,\cdot,\cdot)$
+1:  $\mathbf{z}_{T}=\mathcal{E}(\mathbf{x})$
+2:  for $t=T,\dots,1$ do
+3:     $\boldsymbol{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ if $t>1$, else $\boldsymbol{\epsilon}=\mathbf{0}$
+4:     $\mathbf{z}_{t}^{\prime}=\frac{1}{\sqrt{n_{t}^{2}+\sigma_{t}^{2}}}(\mathbf{z}_{t}-n_{t}\mathbf{z}_{T})$
+5:     $\hat{\mathbf{z}}_{0}=\frac{1}{m_{t}}(\mathbf{z}_{t}-n_{t}\mathbf{z}_{T}-\sigma_{t}\boldsymbol{\epsilon}_{\theta}(\mathbf{z}_{t}^{\prime},\mathbf{z}_{T},t))$
+6:     $\mathbf{z}_{t-1}=k_{1}\mathbf{z}_{t}+k_{2}\hat{\mathbf{z}}_{0}+k_{3}\mathbf{z}_{T}+\tilde{\sigma}_{t}\boldsymbol{\epsilon}$
+7:  end for
+8:  $\mathbf{y}=\mathcal{D}(\hat{\mathbf{z}}_{0})$
+9:  return $\mathbf{y}$
+B.2 Accelerate Inference
+Similar to DDIM [45] and DBIM [64], the inference process of DPBridge can be accelerated using non-Markovian sampling scheme that preserves the marginal distributions of intermediate states.
+Specifically, we perform sampling over a reduced set of timesteps $\{t_{i}\}_{i=1}^{N}$, where $0=t_{0}<t_{1}<\cdots<t_{N-1}<t_{N}=T$, allowing the number of inference steps $N$ to be decoupled from the total number of training steps $T$.
+The accelerated inference process is given by:
+$\begin{split}\mathbf{z}_{t_{n}}&=m_{t_{n}}\hat{\mathbf{z}}_{0}+n_{t_{n}}\mathbf{z}_{T}+\\ &\sqrt{\sigma_{t_{n}}^{2}-g_{n}^{2}}\frac{\mathbf{z}_{t_{n+1}}-m_{t_{n+1}}\hat{\mathbf{z}}_{0}-n_{t_{n+1}}\mathbf{z}_{T}}{\sigma_{t_{n+1}}}+g_{n}\boldsymbol{\epsilon}\end{split}$ (21)
+where $g_{n}$ denotes variance parameter that governs the stochasticity of the reverse sampling process.
+Notably, when $g_{n}=\tilde{\sigma}_{t}^{2}$, as defined Equation 5, the process becomes a Markovian bridge.
+Appendix C Network Architecture
 The basic structure of our network follows the original design of Stable Diffusion.
@@
 To address the dimension incompatibility with pretrained backbone, the input channels of the network are thus doubled to accommodate the concatenated input.
-In addition, as suggested by (Ke et al., 2024), we duplicate the weight tensor of these accommodated layers and divide the values by two to preserve the input magnitude to subsequent layers.
+In addition, as suggested by [28], we duplicate the weight tensor of these accommodated layers and divide the values by two to preserve the input magnitude to subsequent layers.
 As DPBridge is an image-conditioned generative framework, we employ null-text embedding for all cross-attention layers to eliminate the impact of text prompts.
-B.2 Additional Quantitative Results
+Appendix D Implementation Details
+During training, the number of timesteps of is set to be 1000, and we use 50 sampling steps during inference with considerations of both quality and efficiency.
+Training our method takes 30K iterations using the Adam optimizer with a base learning rate of $3\times 10^{-5}$ and an exponential decay strategy after 100-step warm-up.
+The batch size is 2, but with 8 steps gradient accumulation, the effective batch size is equivalent to 16.
+We also apply random horizontal flipping augmentation during training.
+The whole process takes approximately 3 days to complete on a single Nvidia RTX 3090Ti GPU.
+Appendix E Experiments
+E.1 Additional Quantitative Results
 In this section, we further evaluate the robustness of DPBridge system by perturbing the input image with different types and different level intensities of noise.
@@
 During all experiments, the input images are normalized between -1 to 1.
-B.3 Additional Qualitative Results
-To further validate the effectiveness and generalization of DPBridge, we provide qualitative results for various dense prediction tasks, including semantic segmentation, optical flow prediction, and style transfer, by framing them as continuous-valued image-to-image translation problems.
-For semantic segmentation, we convert discrete labels into color maps and use our bridge model to predict colors in RGB space instead of categorical labels.
-For optical flow, the bridge process is defined between the flow image and the first image in the input RGB pair.
-To accommodate this, we adapt the UNet architecture to incorporate both RGB images as conditions during inference and convert the flow map into RGB space, predicting corresponding colors rather than uv coordinates.
-The following figures (Figure 6, Figure 7, Figure 8, Figure 9) illustrate that DPBridge achieves visually satisfactory results across different tasks.
-Note that only qualitative assessments are presented here; detailed implementations and quantitative evaluations are left as directions for future research.
-Figure 6: Qualitative examples of applying DPBridge to semantic segmentation task on CityScape (Cordts et al., 2016) dataset.
+E.2 Additional Qualitative Results
+To further evaluate the effectiveness and generalization capability of DPBridge, we extend the framework to additional dense prediction tasks, including semantic segmentation, optical flow estimation, and style transfer.
+Qualitative results are shown in Figure 7, Figure 8, Figure 9, and Figure 10.
+These extensions are implemented by framing each task as a continuous-valued image-to-image translation problem, and replacing the paired ground truth data accordingly.
+For semantic segmentation, we convert discrete labels into color maps and train the bridge model to predict colors in RGB space rather than class indices.
+For optical flow, we define the bridge process between the flow image and the first frame of the input RGB pair.
+To accommodate this, we modify the UNet architecture to condition on both RGB frames during inference and represent flow as a color-encoded image, enabling prediction in RGB space rather than raw uv coordinates.
+Across all tasks, DPBridge produces visually compelling results, demonstrating its versatility as a unified framework that can be easily adapted to a range of dense prediction scenarios.
+We note that these results are qualitative; quantitative evaluation and task-specific architectural refinements are left for future work.
+Figure 7: Qualitative examples of applying DPBridge to semantic segmentation task on CityScape [12] dataset.
 We transform discrete labels into color maps and us our bridge model to predict colors in RGB space instead of categorical labels.
-Figure 7: Qualitative examples of applying DPBridge to semantic segmentation task on CelebAMask-HQ (Lee et al., 2020) dataset.
+Figure 8: Qualitative examples of applying DPBridge to semantic segmentation task on CelebAMask-HQ [30] dataset.
 We transform discrete labels into color maps and us our bridge model to predict colors in RGB space instead of categorical labels.
-Figure 8: Qualitative examples of applying DPBridge to optical flow prediction task on MPI-Sintel (Butler et al., 2012) dataset.
+Figure 9: Qualitative examples of applying DPBridge to optical flow prediction task on MPI-Sintel [5] dataset.
 The input image is the overlayed RGB pair, and we transform the flow map into RGB space so that the diffusion bridge process can be constructed accordingly.
-Figure 9: Qualitative examples of applying DPBridge to style transfer task on Face-to-Comics (Sxela, 2021) dataset.
+Figure 10: Qualitative examples of applying DPBridge to style transfer task on Face-to-Comics [48] dataset.
+Appendix F Broader Impacts
+DPBridge aims to advance the field of dense prediction through a generative modeling perspective, integrating the strengths of diffusion bridge formulation with visual priors from pretrained backbones.
+Our proposed method is a unified framework that can be easily extend to different tasks, including depth prediction, surface normal estimation, semantic segmentation, optical flow, etc., which will benefit the whole dense prediction research community.
+In addition, improved generalization and precision in dense prediction can also benefit a wide range of real-world applications, such as autonomous driving, augmented reality, and digital content creation.

### v4

@@
 Diffusion models demonstrate remarkable capabilities in capturing complex data distributions and have achieved compelling results in many generative tasks.
-While they have recently been extended to dense prediction tasks such as depth estimation and surface normal prediction, their full potential in this area remains under-explored.
-In dense prediction settings, target signal maps and input images are pixel-wise aligned.
-This makes conventional noise-to-data generation paradigm inefficient, as input images can serve as more informative prior compared to pure noise.
+While they have recently been extended to dense prediction tasks such as depth estimation and surface normal prediction, their full potential in this area remains underexplored.
+As target signal maps and input images are pixel-wise aligned, the conventional noise-to-data generation paradigm is inefficient, and input images can serve as a more informative prior compared to pure noise.
 Diffusion bridge models, which support data-to-data generation between two general data distributions, offer a promising alternative, but they typically fail to exploit the rich visual priors embedded in large pretrained foundation models.
 To address these limitations, we integrate diffusion bridge formulation with structured visual priors and introduce DPBridge, the first latent diffusion bridge framework for dense prediction tasks.
-Our method presents three key contributions: (1) a tractable reverse transition kernel for diffusion bridge process, enabling maximum likelihood training scheme for better compatibility with pretrained backbones; (2) a distribution-aligned normalization technique to mitigate the discrepancies between the bridge and standard diffusion processes; and (3) an auxiliary image consistency loss to preserve fine-grained details.
+To resolve the incompatibility between diffusion bridge models and pretrained diffusion backbones, we propose (1) a tractable reverse transition kernel for the diffusion bridge process, enabling maximum likelihood training scheme; (2) finetuning strategies including distribution-aligned normalization and image consistency loss.
 Experiments across extensive benchmarks validate that our method consistently achieves superior performance, demonstrating its effectiveness and generalization capability under different scenarios.
+Figure 1: We present DPBridge to tackle general dense prediction tasks by establishing conditional probability transport between RGB images and corresponding dense maps in the latent space based on a tractable diffusion bridge process.
 1 Introduction
 Dense prediction refers to a class of problems in which the goal is to assign a prediction value for every pixel in the input image.
-It encompasses a variety of fundamental tasks, including depth estimation [56, 39, 15], surface normal prediction [2, 22, 3], semantic segmentation [53, 8, 51], optical flow [49, 23, 54], etc.
-Recently, image diffusion models have been applied to dense prediction tasks by framing them as conditional generation problems.
-Leveraging the ability to capture complex data distributions, diffusion models serve as powerful priors and yield promising results [28, 17, 16].
-However, these approaches exhibit inherent limitations stemming from the underlying generative mechanism.
-Specifically, the denoising process always starts from Gaussian noise, which carries minimal information about the target signal distribution.
-In such cases, the distributional gap between initial Gaussian prior and target distribution remains large, necessitating more sampling steps and often reliance on manually designed post-hoc guidance [9, 46, 10] or customized generation pipeline [35].
+It encompasses a variety of fundamental tasks, including depth estimation [yang2024depth, ranftl2021vision, eftekhar2021omnidata], surface normal prediction [bae2024rethinking, hu2024metric3d, bae2021estimating], semantic segmentation [weber2021step, cheng2022masked, wang2023images], optical flow [teed2020raft, huang2022flowformer, weinzaepfel2023croco], etc.
+Dense prediction has traditionally relied on discriminative models [ranftl2021vision, zheng2021rethinking, bae2024rethinking], which map RGB inputs to dense signal maps directly via feed-forward manners.
+While these methods offer high inference speed, they usually struggle in ambiguous scenarios-such as occluded regions or dynamic surface normals of glossy objects.
+This limitation is largely due to the lack of uncertainty modeling.
+Recently, generative models such as diffusions have been applied to mitigate these drawbacks by framing dense prediction as a conditional generation task.
+Leveraging the ability of modeling complex data distributions, diffusions serve as robust priors and deliver promising performance as different tasks [ke2024repurposing, garcia2024fine, fu2025geowizard].
+Nevertheless, diffusion-based generative approaches suffer from inherent limitations rooted in their core mechanism: their denoising process initiates from Gaussian noise—an initial prior that contains minimal information about the target dense signal distribution.
+In such cases, the distributional gap between initial Gaussian prior and target distribution remains large, necessitating more sampling steps and often reliance on manually designed post-hoc guidance [chung2022diffusion, song2023pseudoinverse, chung2022improving] or customized generation pipeline [meng2021sdedit].
 This not only increases computational overhead but can also degrade performance and prolong convergence.
 Instead of starting from noise, a more effective alternative is to directly initiate the generative process from the input.
-Since RGB images are pixel-wise aligned with their corresponding dense signal maps, they serve as informative priors.
-This motivates the exploration of diffusion bridge models [34, 37, 14, 33], another line of computational frameworks based on stochastic processes with fixed start and end states, designed to model probability flows between two given distributions.
-We argue that diffusion bridge models are more suitable for dense prediction tasks, as their data-to-data generation paradigm naturally leverage the structural correspondence between inputs and outputs, thereby mitigating inefficiencies inherent in conventional noise-to-data scheme.
-Despite theoretical feasibility, diffusion bridge-based methods [65, 31, 32] have not yet demonstrated competitive performance compared with standard diffusion counterparts.
-This performance gap largely stems from the fact that most bridge models are trained from scratch via score matching objectives, making them incompatible with commonly used pretrained diffusion backbone.
-As a result, they fail to exploit the embedded semantic and spatial prior knowledge, limiting their expressiveness and generalization ability in dense prediction settings.
-Figure 1: We present DPBridge to tackle general dense prediction tasks by establishing conditional probability transport between RGB images and corresponding dense maps in the latent space based on a tractable diffusion bridge process.
+Since RGB images are pixel-wise aligned with their corresponding dense signal maps, they shall act as highly informative priors.
+This motivates the exploration of diffusion bridge models [liu2022let, pavon2021data, de2021diffusion, liu2023learning], another line of computational frameworks based on stochastic processes with fixed start and end states, designed to model probability flows between two given distributions.
+We argue that diffusion bridge models are more suitable for dense prediction tasks, as their data-to-data generation paradigm naturally leverages the structural correspondence between inputs and outputs, thereby mitigating inefficiencies inherent in the conventional noise-to-data scheme.
+Despite theoretical feasibility, diffusion bridge-based methods [zhou2023denoising, li2023bbdm, liu20232] have not yet demonstrated competitive performance compared with standard diffusion counterparts.
+This performance gap primarily arises because most bridge models are trained from scratch via score matching objectives, making them incompatible with the widely adopted pretrained diffusion backbones.
+Consequently, these methods fail to leverage the embedded semantic and spatial prior knowledge, constraining their expressive power and generalization capability in dense prediction scenarios.
 In this paper, we propose Dense Prediction Bridge (DPBridge), the first latent diffusion bridge model tailored for general dense prediction tasks by integrating the strengths of diffusion bridge modeling with visual priors from pretrained backbones.
@@
 Furthermore, we introduce finetuning strategies to better adapt pretrained diffusion backbones for DPBridge construction.
-Leveraging the rich visual priors not only enhances perceptual performance and generalization capability, but also significantly accelerates training convergence.
-Figure 1 provides a high-level overview of DPBridge framework and illustrates its effectiveness across three representative dense prediction tasks: depth estimation, surface normal prediction, and semantic segmentation.
-These results demonstrate that our framework is both theoretically general and practically adaptable to diverse dense prediction scenarios.
+Leveraging rich visual priors not only enhances perceptual performance and generalization ability, but also significantly accelerates training convergence.
+Figure˜1 provides a high-level overview of the DPBridge framework and illustrates its effectiveness across two representative tasks: depth estimation and surface normal prediction.
+These results demonstrate that our method is both theoretically feasible and practically adaptable to diverse scenarios.
 Contributions of this paper can be summarized as follows:
-  • We propose DPBridge, a generative framework to address dense prediction tasks by formulating them as instances of latent diffusion bridges.
+  • We propose DPBridge, a generative framework to address general dense prediction tasks by formulating them as instances of latent diffusion bridges.
   • We derive a tractable reverse transition kernel for diffusion bridge process, enabling maximum likelihood training scheme.
 This is compatible with pretrained backbones, thus improving both training stability and the exploitation of visual prior knowledge.
-  • We introduce a distribution-aligned normalization technique that mitigates distributional discrepancies between the bridge and standard diffusion processes, allowing for more effective finetuning.
-In addition, by incorporating auxiliary image consistency loss, we can better preserve fine-grained spatial details often lost in latent representations, enhancing per-pixel prediction accuracy.
-  • Our method exhibits competitive performance across different tasks in terms of both quantitative metrics and visual inspection.
+  • We introduce finetuning strategies to improve finetuning performance, including a distribution-aligned normalization technique that mitigates distributional discrepancies between the bridge and standard diffusion processes, and an auxiliary image consistency loss to better preserve fine-grained spatial details.
+  • Our method shows competitive performance across different tasks in terms of both quantitative metrics and visual inspection.
 2 Related Works
 Dense Prediction.
-Dense predictions can be characterized as conditional distribution modeling and common solutions are categorized into discriminative and generative ones.
+Common solutions to dense predictions are categorized into discriminative and generative ones.
 Discriminative methods directly map input data to structured outputs in a feed-forward manner.
-[39] introduces transformer architecture and fusion mechanism to better capture both local and global context feature.
-[56] explore the usage of large-scale synthetic and pseudo-labeled images to enhance model’s capacity and robustness.
-[2] incorporates per-pixel ray directions and models the relative rotations between neighboring pixels.
-Recently, diffusion-based generative methods have been applied in this field and showed remarkable performance.
-[28] finetunes pretrained diffusion backbone, leveraging the rich visual knowledge within to enhance depth estimation.
-[16] takes advantage of diffusion priors to jointly estimate depth and surface normals from single image, enhancing detail preservation by decoupling complex data distributions into distinct sub-distributions.
-[19] exploits flow matching to facilitate the distribution shift between images and depth maps.
-Our work follow conventional generative pipeline, but instead of starting from pure noise, we establish direct connections between input image and target signal map distributions, and model conditional probability transport in iterative denoising manner.
+[ranftl2021vision] introduces transformer architecture and fusion mechanism to better capture both local and global context features.
+[bae2024rethinking] incorporates per-pixel ray directions and models the relative rotations between neighboring pixels.
+Recently, diffusion-based generative methods have been applied in this field and have shown remarkable performance.
+[ke2024repurposing] finetunes pretrained diffusion backbone, leveraging the rich visual knowledge within to enhance depth estimation.
+[fu2025geowizard] takes advantage of diffusion priors to jointly estimate depth and surface normals from single image, enhancing detail preservation by decoupling complex data distributions into distinct sub-distributions.
+[gui2024depthfm] exploits flow matching to facilitate the distribution shift between images and depth maps.
+Our work follows the conventional generative pipeline, but instead of starting from pure noise, we establish direct connections between input image and target signal map distributions, and model conditional probability transport in an iterative denoising manner.
 Diffusion and Bridge Models.
-Diffusion models [20, 47] are a family of generative models that are able to capture complex data distributions through a sequence of forward and reverse processes.
-They accelerate the development of modern generative systems, and have consistently obtained state-of-the-art performance across various domains, such as image and video synthesis [41, 21], controlled editing [62, 4], inverse problem solving [9, 46], 3D assets generation [55, 63], and human-centered applications [26, 25, 24].
-Diffusion bridge models have gained increasing attention in generative modeling in recent years [37, 34, 65, 33, 14].
+Diffusion models [ho2020denoising, song2020score] are a family of generative models that excel at capturing complex data distributions through a sequence of forward and reverse processes.
+They have driven the advancement of modern generative systems, and have consistently achieved state-of-the-art performance across various domains, such as image and video synthesis [rombach2022high, ho2022video], controlled editing [zhang2023adding, batzolis2021conditional], inverse problem solving [chung2022diffusion, song2023pseudoinverse], 3D assets generation [xiang2024structured, zhang2024clay], and human-centered applications [ji2024jade, ji2024unsupervised, ji20243d].
+In recent years, diffusion bridge models have gained growing interest in generative modeling [pavon2021data, liu2022let, zhou2023denoising, liu2023learning, de2021diffusion].
 They offer a novel perspective for translating between distributions and enable a more effective data-to-data generation paradigm.
 Recent works have explored different types of bridge processes and their applications.
-[32, 31, 60] develop Schrödinger bridge, Brownian bridge, and Ornstein-Uhlenbeck bridge frameworks, applying them to image restoration tasks such as inpainting, super-resolution.
-[11] proposes a modified inference procedure that enforces data consistency.
-[7] introduces bridge models to text-to-speech systems, enhancing synthesis quality and sampling efficiency.
+[liu20232, li2023bbdm, yue2023image] develop Schrödinger bridge, Brownian bridge, and Ornstein-Uhlenbeck bridge frameworks, applying them to image restoration tasks such as inpainting, super-resolution.
+[chung2024direct] proposes a modified inference procedure that enforces data consistency.
+[chen2023schrodinger] introduces bridge models to text-to-speech systems, enhancing synthesis quality.
 To the best of our knowledge, no prior work has explored the use of bridge models for dense prediction tasks to facilitate distribution transport between images and dense signal maps, despite their structural closeness.
-Our work aims to address this research gap.
-Conceptually, DPBridge is most related to DDBM [65], but the two differ in several key aspects: DDBM is trained from scratch using a score matching objective, whereas DPBridge is finetuned from large pretrained diffusion backbones using a maximum likelihood objective.
-In addition, we introduce training strategies tailored for efficient finetuning, enabling better use of visual priors and significantly improving performance on dense prediction settings.
+Our work aims to fill this research gap.
+Conceptually, DPBridge is most related to DDBM [zhou2023denoising], but the two differ in several key aspects: DDBM is trained from scratch using a score matching objective, whereas DPBridge is finetuned from large pretrained diffusion backbones using a maximum likelihood objective.
+In addition, we introduce strategies for efficient finetuning, enabling better use of visual priors and significantly improving performance on dense prediction settings.
 3 Preliminary
 3.1 Diffusion Process
-A diffusion process =$\mathbb{Q}=\{\mathbf{z}_{t}\in\mathbb{R}^{d}:t\in[0,T]\}$ on space $\mathbb{R}^{d}$ follows the stochastic differential equation of form
+A diffusion process $\mathbb{Q}=\{\mathbf{z}_{t}\in\mathbb{R}^{d}:t\in[0,T]\}$ on space $\mathbb{R}^{d}$ follows the stochastic differential equation of form
 $\mathrm{d}\mathbf{z}_{t}=f(\mathbf{z}_{t},t)\mathrm{d}t+g(t)\mathrm{d}\mathbf{w}_{t}$ (1)
 where $\mathbf{w}_{t}$ is a standard Wiener process, $f:\mathbb{R}^{d}\times[0,T]\rightarrow\mathbb{R}$ and $g:[0,T]\rightarrow\mathbb{R}$ are the drift and diffusion coefficients that describe the evolution of this stochastic process.
-The transition kernel is specified as $q(\mathbf{z}_{t}|\mathbf{z}_{0})=\mathcal{N}(\alpha_{t}\mathbf{z}_{0},\sigma_{t}^{2}\mathbf{I})$ and the noise schedule comprising differentiable functions $\alpha_{t},\sigma_{t}$ depends on the design choice of drift and diffusion coefficients.
-It is also common to assess the diffusion process in terms of the signal-to-noise ratio (SNR), defined as $\text{SNR}_{t}=\alpha_{t}^{2}/\sigma_{t}^{2}$.
+The transition kernel is specified as $q(\mathbf{z}_{t}|\mathbf{z}_{0})=\mathcal{N}(\alpha_{t}\mathbf{z}_{0},\eta_{t}^{2}\mathbf{I})$ and the noise schedule comprising differentiable functions $\alpha_{t},\eta_{t}$ depends on the design choice of drift and diffusion coefficients.
+It is also common to assess the diffusion process in terms of the signal-to-noise ratio (SNR), defined as $\text{SNR}_{t}=\alpha_{t}^{2}/\eta_{t}^{2}$.
 3.2 Diffusion Bridge Process
 Diffusion bridge process is a special type of diffusion process that is conditioned to end at a specified terminal point at timestep $T$.
-One natural way of its construction is to derive the conditioned process from a reference diffusion process given terminal constraints.
-We can use Doob’s h-transform technique [36, 42] to modify a reference process $\mathbb{Q}$ defined in Equation 1 to ensure that the event $\mathbf{z}_{T}=\mathbf{x}$ happens, where $\mathbf{x}$ is a predefined value.
+Its dual fixed-endpoint characteristics enable a bidirectional distribution transport.
+Its construction can be derived from a reference diffusion process conditioned on terminal constraints.
+We can use Doob’s h-transform technique [oksendal2013stochastic, sarkka2019applied] to modify a reference process $\mathbb{Q}$ defined in Equation˜1 to ensure that the event $\mathbf{z}_{T}=\mathbf{x}$ happens, where $\mathbf{x}$ is a predefined value.
 The conditioned process $\mathbb{Q}^{\mathbf{x}}(\cdot):=\mathbb{Q}\left(\cdot|\mathbf{z}_{T}=\mathbf{x}\right)$ thus follows the law of
@@
 $q(\mathbf{z}_{t}|\mathbf{z}_{0},\mathbf{z}_{T})$ $=\mathcal{N}(m_{t}\mathbf{z}_{0}+n_{t}\mathbf{z}_{T},\sigma_{t}^{2}\mathbf{I})$ (3)
-$m_{t}=\alpha_{t}\left(1-\frac{\text{SNR}_{T}}{\text{SNR}_{t}}\right),\quad n_{t}$ $=\frac{\text{SNR}_{T}}{\text{SNR}_{t}}\frac{\alpha_{t}}{\alpha_{T}},\quad\sigma_{t}=\sigma_{t}\sqrt{1-\frac{\text{SNR}_{T}}{\text{SNR}_{t}}}$
-where $\text{SNR}_{t},\alpha_{t},\sigma_{t}^{2}$ are all ingredients defined in the reference process $\mathbb{Q}$.
+$m_{t}$ $=\alpha_{t}\left(1-\frac{\text{SNR}_{T}}{\text{SNR}_{t}}\right),\quad n_{t}=\frac{\text{SNR}_{T}}{\text{SNR}_{t}}\frac{\alpha_{t}}{\alpha_{T}},$
+$\sigma_{t}$ $=\eta_{t}\sqrt{1-\frac{\text{SNR}_{T}}{\text{SNR}_{t}}}$
+where $\text{SNR}_{t},\alpha_{t},\sigma_{t}^{2}$ are all defined in the reference process $\mathbb{Q}$.
+In Equation˜3, $m_{t}$ and $n_{t}$ are coefficients that enforce the bridge’s dual endpoints: at $t=0,m_{0}=1,n_{0}=0\rightarrow z_{t}=z_{0}$; at $t=T,m_{T}=0,n_{T}=1\rightarrow z_{t}=z_{T}$.
 4 Method
@@
 depth, surface normal, segmentation map), and $\mathbf{x}\in\mathbb{R}^{H\times W\times 3}$ is the input RGB image.
-In the following sections, we’ll elaborate the details of DPBridge framework, which is constructed based on pretrained diffusion backbone through proper finetuning.
+In the following sections, we’ll elaborate the details of the DPBridge framework, which is constructed based on pretrained diffusion backbone through proper finetuning.
 4.2 Dense Prediction Bridge
-Given paired training data, a straightforward approach is to construct the diffusion bridge model from scratch, following the formulation in Section 3.2.
+Given paired training data, a straightforward approach is to construct the diffusion bridge model from scratch, following the formulation in Sec.˜3.2.
 However, this approach is extremely resource-demanding, making it hard to implement with limited computation power.
-What’s more, it lacks the ability to harness rich visual prior embedded in the foundation models pretrained with large-scale data,which have been shown to significantly enhance both perceptual quality and generalization capability in dense prediction tasks.
-To address these limitations and fully exploit the visual prior, our DPBridge is build upon pretrained diffusion backbones through finetuning.
+What’s more, it cannot to harness rich visual prior embedded in the foundation models pretrained with large-scale data, which have been shown to significantly enhance both perceptual quality and generalization capability in dense prediction tasks.
+To address these limitations and fully exploit the visual prior, our DPBridge is built upon pretrained diffusion backbones through finetuning.
 We explore several strategies to handle the misalignment between bridge process and standard diffusion process, facilitating more efficient knowledge transfer.
-An overview of our training pipeline is provided in Figure 2, followed by a detailed explanation of each component.
+An overview of our training pipeline is provided in Figure˜2, followed by a detailed explanation of each component.
 4.3 Latent Space Implementation
@@
 In this way the diffusion bridge process is operated entirely in the latent space.
-To recover the original signal map, we use the VAE decoder to decode the predicted latent $\mathbf{y}=\mathcal{D}(\hat{\mathbf{z}}_{0})$.
+To recover the original signal map, we use the VAE decoder to decode the predicted latent $\hat{\mathbf{y}}=\mathcal{D}(\hat{\mathbf{z}}_{0})$.
 For clarity, all subsequent derivations in this paper are conducted in latent space.
@@
 Maximum Likelihood Loss.
-A widely adopted approach for learning the reverse process associated with Equation 2 follows the theoretical framework of [1], where a neural network is trained to approximate the score function [65].
+A widely adopted approach for learning the reverse process associated with Equation˜2 follows the theoretical framework of [anderson1982reverse], where a neural network is trained to approximate the score function [zhou2023denoising].
 In contrast, we observe that in the diffusion bridge setting, the reverse transition kernel admits a tractable analytical form.
 This insight enables us to adopt a maximum likelihood training paradigm in place of score matching.
-Since widely used latent diffusion models like Stable Diffusion [41] are pretrained under a likelihood-based loss function, finetuning with a consistent objective is more compatible, which helps preserve prior knowledge and improves training stability.
+Since widely used latent diffusion models like Stable Diffusion [rombach2022high] are pretrained under a likelihood-based loss function, finetuning with a consistent objective is more compatible, which helps preserve prior knowledge and improves training stability.
 Specifically, the training process is performed by optimizing the Evidence Lower Bound (ELBO).
 Given the encoded training pair $\big(\mathbf{z}_{0},\mathbf{z}_{T}\big)$, the expectation of data log-likelihood possesses the following ELBO formulation:
-[…]
+$\emph{ELBO}=$ $\mathbb{E}_{p(\mathbf{z}_{0})}\Big[\mathbb{E}_{p(\mathbf{z}_{1}\mid\mathbf{z}_{0})}\big[\log p_{\theta}(\mathbf{z}_{0}\mid\mathbf{z}_{1},\mathbf{z}_{T})\big]$- (4)
 $\sum_{t=2}^{T}\mathbb{E}_{p(\mathbf{z}_{t}|\mathbf{z}_{0})}$ $\big[\mathbb{D}_{KL}\big(q(\mathbf{z}_{t-1}\mid\mathbf{z}_{t},\mathbf{z}_{0},\mathbf{z}_{T})\,||\,p_{\theta}(\mathbf{z}_{t-1}\mid\mathbf{z}_{t},\mathbf{z}_{T})\big)\big]\Big]$
@@
 For simplicity, we define the following notations: $a_{t}=m_{t}/m_{t-1},b_{t}=n_{t}-a_{t}\cdot n_{t-1},\delta_{t}^{2}=\sigma_{t}^{2}-a_{t}^{2}\cdot\sigma_{t-1}^{2}$.
-Based on Equation 3, we further derive the mean value term:
+Based on Equation˜3, we further derive the mean value term:
 $\tilde{\boldsymbol{\mu}}_{t}(\mathbf{z}_{t},\mathbf{z}_{0},\mathbf{z}_{T})$ $=k_{1}\mathbf{z}_{t}+k_{2}\mathbf{z}_{0}+k_{3}\mathbf{z}_{T}$ (6)
-$k_{1}=\frac{\sigma_{t-1}^{2}}{\sigma_{t}^{2}}\cdot a_{t},\quad k_{2}=\frac{\delta_{t}^{2}}{\sigma_{t}^{2}}\cdot$ $m_{t-1},\quad k_{3}=\frac{\delta_{t}^{2}}{\sigma_{t}^{2}}\cdot n_{t-1}-\frac{\sigma_{t-1}^{2}}{\sigma_{t}^{2}}\cdot a_{t}\cdot b_{t},$
+$k_{1}$ $=\frac{\sigma_{t-1}^{2}}{\sigma_{t}^{2}}\cdot a_{t},\quad k_{2}=\frac{\delta_{t}^{2}}{\sigma_{t}^{2}}\cdot m_{t-1},$
+$k_{3}$ $=\frac{\delta_{t}^{2}}{\sigma_{t}^{2}}\cdot n_{t-1}-\frac{\sigma_{t-1}^{2}}{\sigma_{t}^{2}}\cdot a_{t}\cdot b_{t},$
 and the variance term:
 $\tilde{\sigma}_{t}^{2}=\frac{\sigma_{t-1}^{2}}{\sigma_{t}^{2}}\cdot\delta_{t}^{2}$ (7)
-Assuming $p_{\theta}(\mathbf{z}_{t-1}|\mathbf{z}_{t},\mathbf{z}_{T})\sim\mathcal{N}(\boldsymbol{\mu}_{\theta}(\mathbf{z}_{t},\mathbf{z}_{T},t),\boldsymbol{\Sigma}_{\theta}(\mathbf{z}_{t},\mathbf{z}_{T},t))$ is also a Gaussian distribution with fixed variance $\boldsymbol{\Sigma}_{\theta}(\mathbf{z}_{t},\mathbf{z}_{T},t)=\tilde{\sigma}_{t}^{2}$ for simplicity of derivation.
-Following the strategy of DDPM [20], we reparameterize the model $\boldsymbol{\epsilon}_{\theta}(\mathbf{z}_{t},\mathbf{z}_{T},t)$ to predict noise instead of posterior mean.
+Assuming $p_{\theta}(\mathbf{z}_{t-1}|\mathbf{z}_{t},\mathbf{z}_{T})$ also follows a Gaussian distribution $\mathcal{N}(\boldsymbol{\mu}_{\theta}(\mathbf{z}_{t},\mathbf{z}_{T},t),\boldsymbol{\Sigma}_{\theta}(\mathbf{z}_{t},\mathbf{z}_{T},t))$ with fixed variance $\boldsymbol{\Sigma}_{\theta}(\mathbf{z}_{t},\mathbf{z}_{T},t)=\tilde{\sigma}_{t}^{2}$ for simplicity of derivation.
+Following the strategy of DDPM [ho2020denoising], we reparameterize the model $\boldsymbol{\epsilon}_{\theta}(\mathbf{z}_{t},\mathbf{z}_{T},t)$ to predict noise instead of posterior mean.
 Under this formulation, the ELBO objective is equivalent to:
@@
 $\cong\mathbb{E}_{t,\mathbf{z}_{0},\mathbf{z}_{t},\mathbf{z}_{T}}\left[\left\|\boldsymbol{\epsilon}-\boldsymbol{\epsilon}_{\theta}(\mathbf{z}_{t},\mathbf{z}_{T},t)\right\|^{2}\right]$
+While Equation˜8’s MSE-like form resembles DDPM’s noise loss, its derivation relies on the analytical reverse kernel of the bridge process.
+The inclusion of $z_{T}$ is not “standard conditional generation” but a requirement for enforcing the bridge’s dual-endpoint constraint.
+For more derivation details, please refer to Appendix˜A
 Image Consistency Loss.
@@
 This misalignment weakens the model’s denoising ability and hinders effective knowledge transfer.
-To address this issue, we propose a distribution alignment normalization technique that transforms intermediate latents computed from our bridge process Equation 3 to match the distributional form of a standard diffusion process.
+To address this issue, we propose a distribution alignment normalization technique that transforms intermediate latents computed from our bridge process Equation˜3 to match the distributional form of a standard diffusion process.
 $\mathbf{z}_{t}^{\prime}=\frac{\mathbf{z}_{t}-m_{t}\mathbf{z}_{T}}{\sqrt{n_{t}^{2}+\sigma_{t}^{2}}}=\frac{n_{t}}{\sqrt{n_{t}^{2}+\sigma_{t}^{2}}}\mathbf{z}_{0}+\frac{\sigma_{t}}{\sqrt{n_{t}^{2}+\sigma_{t}^{2}}}\boldsymbol{\epsilon}_{t}$ (11)
-Since the square sum of coefficients is equal to 1, the normalized bridge process $\{\mathbf{z}_{t}^{\prime}:t\in[0,T]\}$ is the same as a standard VP process, making the marginal distribution of these two aligned.
-Therefore by feeding the normalized sample $\mathbf{z}_{t}^{\prime}$ in the network $\boldsymbol{\epsilon}_{\theta}$, we can fully exploit the denoising capability from pretrained diffusion backbone.
-We note that a similar strategy was independently proposed in FrameBridge [52], where it is applied in the context of image-to-video generation.
+Since the square sum of coefficients is equal to 1, the marginal distribution of the normalized bridge process $\{\mathbf{z}_{t}^{\prime}:t\in[0,T]\}$ can thus align with that of the standard diffusion process, while preserving bridge-specific dynamics.
+Therefore, by feeding the normalized sample $\mathbf{z}_{t}^{\prime}$ in the network $\boldsymbol{\epsilon}_{\theta}$, we can fully exploit the denoising capability from the pretrained diffusion backbone.
+We note that a similar strategy was independently proposed in FrameBridge [wang2024framebridge], where it is applied in the context of image-to-video generation.
 Our method, developed concurrently, is specifically designed to facilitate adaptation in dense prediction tasks under the DPBridge framework.
-An overview of the full training and inference procedures are summarized in Algorithm 1 Algorithm 2.
-For more derivation details, please refer to Appendix A.
+An overview of the full training and inference procedures is summarized in Algorithm˜1 Algorithm˜2.
 5 Experiments
@@
 Next, we present ablation studies to investigate the contribution of each design component.
-Implementation details are presented in Appendix D for reference.
+Implementation details are presented in Appendix˜D for reference.
 5.1 Experiments Setup
@@
 The basic setting for depth estimation and surface normal prediction experiments are similar.
-We use two synthetic datasets, HyperSim [40] and Virtual KITTI [6], for training, which cover both indoor and outdoor scenes with high-quality ground-truth annotations.
+We use two synthetic datasets, HyperSim [roberts2021hypersim] and Virtual KITTI [cabon2020virtual], for training, which cover both indoor and outdoor scenes with high-quality ground-truth annotations.
 To comprehensively evaluate the model performance and demonstrate its generalization capability, we conduct zero-shot evaluations on real-world datasets.
-For depth estimation, this includes NYUv2 [44], KITTI [18], ScanNet [13], ETH3D [43] and DIODE [50].
-For surface normal prediction, we additionally incorporate iBims [29] and Sintel [5] datasets.
-For fair comparison, we use the same validation split as in previous methods [28].
+For depth estimation, this includes NYUv2 [silberman2012indoor], KITTI [geiger2012we], ScanNet [dai2017scannet], ETH3D [schops2017multi] and DIODE [vasiljevic2019diode].
+For surface normal prediction, we additionally incorporate iBims [koch2018evaluation] and Sintel [butler2012naturalistic] datasets.
+For fair comparison, we use the same validation split as in previous methods [ke2024repurposing].
 Evaluation Metrics.
@@
 […]
-DiverseDepth [58] 11.7 87.5 19.0 70.4 22.8 69.4 10.9 88.2 37.6 63.1 12.2
-MiDaS [38] 11.1 88.5 23.6 63.0 18.4 75.2 12.1 84.6 33.2 71.5 11.8
-LeReS [59] 9.0 91.6 14.9 78.4 17.1 77.7 9.1 91.1 27.1 76.6 9.5
-Omnidata [15] 7.4 94.5 14.9 83.5 16.6 77.8 7.5 93.2 33.9 74.2 9.2
-HDN [61] 6.9 94.8 11.5 86.7 12.1 83.3 8.0 93.6 24.6 78.0 7.2
-DPT [39] 9.8 90.3 10.0 90.1 7.8 94.6 8.2 93.8 18.2 75.8 7.0
-Depth Anything V2 [56] 4.4 97.9 7.5 94.8 13.2 84.2 4.4 98.0 6.5 95.4 2.3
-Marigold [28] 6.0 95.9 10.5 90.4 7.1 95.1 6.9 94.5 31.0 71.2 6.0
-DepthFM [19] 6.5 95.6 8.3 93.4 - - - - 22.5 80.0 3.8
-Geowizard [16] 5.7 96.2 14.4 82.0 7.5 94.3 6.1 95.8 31.4 77.1 6.2
-E2E-FT [17] 5.4 96.4 9.6 92.1 6.4 95.9 5.8 96.5 30.3 77.6 3.6
-I2SB [32] 23.5 63.6 35.1 46.4 53.4 40.5 21.1 67.4 55.5 58.2 14.0
-BBDM [31] 33.2 49.6 58.6 25.6 52.0 40.6 25.2 61.2 70.1 46.9 15.2
-DDBM [65] 26.4 62.2 55.1 27.6 69.5 33.8 23.2 65.6 52.3 56.9 14.5
+DiverseDepth [yin2020diversedepth] 11.7 87.5 19.0 70.4 22.8 69.4 10.9 88.2 37.6 63.1 12.2
+MiDaS [ranftl2020towards] 11.1 88.5 23.6 63.0 18.4 75.2 12.1 84.6 33.2 71.5 11.8
+LeReS [yin2021learning] 9.0 91.6 14.9 78.4 17.1 77.7 9.1 91.1 27.1 76.6 9.5
+Omnidata [eftekhar2021omnidata] 7.4 94.5 14.9 83.5 16.6 77.8 7.5 93.2 33.9 74.2 9.2
+HDN [zhang2022hierarchical] 6.9 94.8 11.5 86.7 12.1 83.3 8.0 93.6 24.6 78.0 7.2
+DPT [ranftl2021vision] 9.8 90.3 10.0 90.1 7.8 94.6 8.2 93.8 18.2 75.8 7.0
+Depth Anything V2 [yang2024depth] 4.4 97.9 7.5 94.8 13.2 84.2 4.4 98.0 6.5 95.4 2.3
+Marigold [ke2024repurposing] 6.0 95.9 10.5 90.4 7.1 95.1 6.9 94.5 31.0 71.2 6.0
+DepthFM [gui2024depthfm] 6.5 95.6 8.3 93.4 - - - - 22.5 80.0 3.8
+Geowizard [fu2025geowizard] 5.7 96.2 14.4 82.0 7.5 94.3 6.1 95.8 31.4 77.1 6.2
+E2E-FT [garcia2024fine] 5.4 96.4 9.6 92.1 6.4 95.9 5.8 96.5 30.3 77.6 3.6
+I2SB [liu20232] 23.5 63.6 35.1 46.4 53.4 40.5 21.1 67.4 55.5 58.2 14.0
+BBDM [li2023bbdm] 33.2 49.6 58.6 25.6 52.0 40.6 25.2 61.2 70.1 46.9 15.2
+DDBM [zhou2023denoising] 26.4 62.2 55.1 27.6 69.5 33.8 23.2 65.6 52.3 56.9 14.5
 DPBridge (SD1.5) 6.9 95.7 11.2 87.3 8.1 93.2 7.5 94.7 29.7 78.6 5.4
@@
 All metrics are presented in percentage terms.
-FF, DM, and BM correspond to feed-forward, diffusion models, and bridge models.
+$F$, DM, and BM correspond to feed-forward, diffusion models, and bridge models.
 The best and second-best results are highlighted in bold and underline formats.
@@
 […]
-EESNU [3] 16.2 58.6 - - 20.0 58.5 42.1 11.5 6.8
-Omnidata v1 [15] 23.1 45.8 22.9 47.4 19.0 62.1 41.5 11.4 8.8
-Omnidata v2 [27] 17.2 55.5 16.2 60.2 18.2 63.9 40.5 14.7 6.2
-DSINE [2] 16.4 59.6 16.2 61.0 17.1 67.4 34.9 21.5 3.5
-Marigold [28] 18.8 55.9 17.7 58.8 18.4 64.3 39.1 14.9 6.4
-Geowizard [16] 17.0 56.5 15.4 61.6 13.0 65.3 40.4 13.2 4.2
-E2E-FT [17] 16.5 60.4 14.7 66.1 16.1 69.7 33.5 22.3 2.1
-StableNormal [57] 17.7 55.9 15.8 61.2 17.2 66.6 - - 5.0
+EESNU [bae2021estimating] 16.2 58.6 - - 20.0 58.5 42.1 11.5 6.8
+Omnidata v1 [eftekhar2021omnidata] 23.1 45.8 22.9 47.4 19.0 62.1 41.5 11.4 8.8
+Omnidata v2 [kar20223d] 17.2 55.5 16.2 60.2 18.2 63.9 40.5 14.7 6.2
+DSINE [bae2024rethinking] 16.4 59.6 16.2 61.0 17.1 67.4 34.9 21.5 3.5
+Marigold [ke2024repurposing] 18.8 55.9 17.7 58.8 18.4 64.3 39.1 14.9 6.4
+Geowizard [fu2025geowizard] 17.0 56.5 15.4 61.6 13.0 65.3 40.4 13.2 4.2
+E2E-FT [garcia2024fine] 16.5 60.4 14.7 66.1 16.1 69.7 33.5 22.3 2.1
+StableNormal [ye2024stablenormal] 17.7 55.9 15.8 61.2 17.2 66.6 - - 5.0
 DPBridge (SD1.5) 18.1 54.2 18.0 54.5 24.3 51.2 41.2 19.4 8.0
@@
 All metrics are presented in percentage terms.
-FF, DM, and BM correspond to feed-forward, diffusion models, and bridge models.
+$F$, DM, and BM correspond to feed-forward, diffusion models, and bridge models.
 The best and second-best results are highlighted in bold and underline formats.
@@
 Figure 5: Qualitative comparison with diffusion-based surface normal prediction methods on in-the-wild images.
-DPBridge’s results exhibit more fine-grained geometric details, such as the curvature of architectural structures, the contour of coral and reef, and the fur texture of the cat.
+DPBridge exhibits more fine-grained geometric details, such as the coral and reef’s contour, the kitten’s fur details, and the robot’s body composition curvature.
 5.2 Quantitative and Qualitative Results
-Table 1 and Table 2 summarize the quantitative results of DPBridge on zero-shot benchmarks across depth estimation and surface normal prediction tasks.
-We evaluate two variants of our model—DPBridge (SD1.5) and DPBridge (SD2.1)—built on top of Stable Diffusion v1.5 and v2.1-base [41], respectively.
+Tab.˜1 and Tab.˜2 summarize the quantitative results of DPBridge on zero-shot benchmarks across depth estimation and surface normal prediction tasks.
+We evaluate two variants of our model—DPBridge (SD1.5) and DPBridge (SD2.1)—built on top of Stable Diffusion v1.5 and v2.1-base [rombach2022high], respectively.
 For depth estimation, DPBridge (SD2.1) achieves an AbsRel of 4.4 and $\delta_{1}$ accuracy of 97.6% on NYUv2, and 4.5 AbsRel / 97.4% $\delta_{1}$ on ScanNet, both among the best in class.
@@
 Furthermore, among diffusion-based approaches, DPBridge consistently achieves the best or second-best scores, validating our core hypothesis that bridge formulation provides clear advantages over standard diffusion frameworks in dense prediction tasks, and combining bridge modeling with pretrained visual prior offers a more efficient and generalizable framework.
-Figure 5 and Figure 5 present qualitative comparisons of DPBridge with SOTA diffusion-based methods on in-the-wild images for depth estimation and surface normal prediction, respectively.
-As shown in Figure 5, DPBridge generates depth maps with sharper object boundaries and more coherent scene geometry, particularly in challenging cases such as urban scenes or complex foreground objects like animals, grass and flowers.
+Figure˜5 and Figure˜5 present qualitative comparisons of DPBridge with SOTA diffusion-based methods on in-the-wild images for depth estimation and surface normal prediction, respectively.
+As shown in Figure˜5, DPBridge generates depth maps with sharper object boundaries and more coherent scene geometry, particularly in challenging cases such as urban scenes or complex foreground objects like animals, grass, and flowers.
 In contrast, competing methods like Geowizard and DepthFM often produce over-smoothed outputs or fail to preserve depth discontinuities.
-In Figure 5, DPBridge demonstrates stronger fidelity in fine-grained geometric details, such as the curvature of architectural structures, the contour of coral and reef, and the fur texture of the cat.
-Competing methods like Marigold tends to focus on the center object in the image and ignore the environment, while StableNormal produces unreliable outputs.
+In Figure˜5, DPBridge demonstrates stronger fidelity in fine-grained geometric details, such as the contour of coral and reef, the fur details of the cat, and the body composition curvature of the robot.
+Competing method like Marigold tends to focus on the center object in the image and ignore the environment, while StableNormal produces unreliable outputs.
 These visual results reinforce the quantitative findings and highlight the effectiveness of DPBridge in modeling structured output distributions with both global coherence and fine details.
@@
 To evaluate the effectiveness of each design component and training setting, we conduct ablation studies using depth estimation tasks on the NYUv2 and KITTI datasets.
-The results are summarized in Table 3 and Figure 6, with detailed discussions provided below.
+The results are summarized in Tab.˜3 and Figure˜6, with detailed discussions provided below.
 Visual Prior Exploitation.
 To evaluate the usage of visual priors, we compare finetuning with training from scratch.
-As shown in the first row of Table 3, finetuning yields significantly better results under identical settings.
+As shown in the first row of Tab.˜3, finetuning yields significantly better results under identical settings.
 This is attributed to stronger visual priors from foundation models and faster convergence.
-Additionally, the last rows of Table 1 and Table 2 show that SD2.1 outperforms SD1.5, highlighting that higher-quality priors further boost performance.
+Additionally, the last rows of Tab.˜1 and Tab.˜2 show that SD2.1 outperforms SD1.5, highlighting that higher-quality priors further boost performance.
 These results confirm the advantage of building DPBridge on strong pretrained backbones.
@@
 We next assess the benefit of aligning marginal distributions between the diffusion bridge and standard diffusion processes.
-As shown in the second and fourth rows of Table 3, adding a distribution alignment normalization step before the noise prediction network consistently improves performance.
+As shown in the second and fourth rows of Tab.˜3, adding a distribution alignment normalization step before the noise prediction network consistently improves performance.
 Image Consistency Loss.
-As shown in the third and fourth rows of Table 3, the auxiliary image consistency loss improves performance but has a smaller impact than distribution alignment.
+As shown in the third and fourth rows of Tab.˜3, the auxiliary image consistency loss improves performance but has a smaller impact than distribution alignment.
 This is likely because it mainly affects the boundaries or overlapping regions between masked and valid areas, which are limited in scope.
@@
 We evaluate performance under varying sampling steps during inference.
-As shown in Figure 6, perception quality improves rapidly with more steps and saturates after 10.
-Notably, compared to Table 1, where other diffusion-based methods use 50 steps, DPBridge achieves comparable results with significantly fewer steps, highlighting its efficiency.
+As shown in Figure˜6, perception quality improves rapidly with more steps and saturates after 10.
+Notably, compared to Tab.˜1, where other diffusion-based methods use 50 steps, DPBridge achieves comparable results with significantly fewer steps, highlighting its efficiency.
 6 Conclusion
@@
 Second, our formulation is restricted to certain bridge parameterization, and exploring alternatives to adapt to more advanced pretrained backbone is a promising future direction.
+Acknowledgement
+This research is funded in part by an ARC (Australian Research Council) Discovery Grant of DP220100800.
+Prof Hongdong Li holds concurrent appointments as a Full Professor with the ANU and as an Amazon Scholar with Amazon (parttime).
+This paper describes work performed at ANU and is not associated with Amazon.
 Appendix A Derivation Details
-In this section, we’ll provide derivation details of the maximum likelihood loss presented in Section 4.4.
-Its core is the posterior distribution $q(\mathbf{z}_{t-1}|\mathbf{z}_{t},\mathbf{z}_{0},\mathbf{z}_{T})$ as in Equation 5, whose calculation involves three terms: $q(\mathbf{z}_{t}|\mathbf{z}_{t-1},\mathbf{z}_{T})$, $q(\mathbf{z}_{t-1}|\mathbf{z}_{0},\mathbf{z}_{T})$, and $q(\mathbf{z}_{t}|\mathbf{z}_{0},z\mathbf{z}_{T})$.
-The last two term of the posterior can be retrieved directly through the marginal distribution of diffusion bridge’s forward process, but further derivation is required for the first term.
-Specifically, as specified in Equation 3:
+In this section, we’ll provide derivation details of the maximum likelihood loss presented in Sec.˜4.4.
+Its core is the posterior distribution $q(\mathbf{z}_{t-1}|\mathbf{z}_{t},\mathbf{z}_{0},\mathbf{z}_{T})$ as in Equation˜5, whose calculation involves three terms: $q(\mathbf{z}_{t}|\mathbf{z}_{t-1},\mathbf{z}_{T})$, $q(\mathbf{z}_{t-1}|\mathbf{z}_{0},\mathbf{z}_{T})$, and $q(\mathbf{z}_{t}|\mathbf{z}_{0},\mathbf{z}_{T})$.
+The last two terms of the posterior can be retrieved directly through the marginal distribution of the diffusion bridge’s forward process, but further derivation is required for the first term.
+Specifically, as specified in Equation˜3:
 $\begin{split}q(\mathbf{z}_{t}|\mathbf{z}_{0},\mathbf{z}_{T})&=\mathcal{N}(m_{t}\mathbf{z}_{0}+n_{t}\mathbf{z}_{T},\sigma_{t}^{2}\mathbf{I})\\ \mathbf{z}_{t}&=m_{t}\mathbf{z}_{0}+n_{t}\mathbf{z}_{T}+\sigma_{t}\boldsymbol{\epsilon}_{t}\end{split}$ (12)
 $\begin{split}q(\mathbf{z}_{t-1}|\mathbf{z}_{0},\mathbf{z}_{T})&=\mathcal{N}(m_{t-1}\mathbf{z}_{0}+n_{t-1}\mathbf{z}_{T},\sigma_{t-1}^{2}\mathbf{I})\\ \mathbf{z}_{t-1}&=m_{t-1}\mathbf{z}_{0}+n_{t-1}\mathbf{z}_{T}+\sigma_{t-1}\boldsymbol{\epsilon}_{t-1}\end{split}$ (13)
-To derive the transition probability $q(\mathbf{z}_{t}|\mathbf{z}_{t-1},\mathbf{z}_{T})$, we substitute $\mathbf{z}_{0}$ in Equation 12 with derivation from Equation 13, obtain
+To derive the transition probability $q(\mathbf{z}_{t}|\mathbf{z}_{t-1},\mathbf{z}_{T})$, we substitute $\mathbf{z}_{0}$ in Equation˜12 with derivation from Equation˜13, obtain
 $\mathbf{z}_{t}$ $=a_{t}\mathbf{z}_{t-1}+b_{t}\mathbf{z}_{T}+\delta_{t}\boldsymbol{\epsilon}$ (14)
@@
 Thus, we can get $q(\mathbf{z}_{t}|\mathbf{z}_{t-1},\mathbf{z}_{T})=\mathcal{N}(a_{t}\mathbf{z}_{t-1}+b_{t}\mathbf{z}_{T},\delta_{t}^{2}\mathbf{I})$, which is also a Gaussian distribution.
-Putting all these terms back to Equation 5, we’ll have:
+Putting all these terms back to Equation˜5, we’ll have:
 $q(\mathbf{z}_{t-1}|\mathbf{z}_{t},\mathbf{z}_{0},\mathbf{z}_{T})\sim\mathcal{N}(\tilde{\boldsymbol{\mu}}_{t}(\mathbf{z}_{t},\mathbf{z}_{0},\mathbf{z}_{T}),\tilde{\sigma}_{t}^{2}\mathbf{I})$ (15)
 $=$ $\frac{q(\mathbf{z}_{t}|\mathbf{z}_{t-1},\mathbf{z}_{T})q(\mathbf{z}_{t-1}|\mathbf{z}_{0},\mathbf{z}_{T})}{q(\mathbf{z}_{t}|\mathbf{z}_{0},\mathbf{z}_{T})}$
-$\propto$ $\exp\bigg{[}-\frac{1}{2}\bigg{(}\frac{(\mathbf{z}_{t}-a_{t}\mathbf{z}_{t-1}-b_{t}\mathbf{z}_{T})^{2}}{\delta_{t}^{2}}+\frac{(\mathbf{z}_{t-1}-m_{t-1}\mathbf{z}_{0}-n_{t-1}\mathbf{z}_{T}^{2}}{\sigma_{t-1}^{2}}-\frac{(\mathbf{z}_{t}-m_{t}\mathbf{z}_{0}-n_{t}\mathbf{z}_{T}^{2}}{\sigma_{t}^{2}}\bigg{)}\bigg{]}$
+$\propto$ $\exp\bigg[-\frac{1}{2}\bigg(\frac{(\mathbf{z}_{t}-a_{t}\mathbf{z}_{t-1}-b_{t}\mathbf{z}_{T})^{2}}{\delta_{t}^{2}}+\frac{(\mathbf{z}_{t-1}-m_{t-1}\mathbf{z}_{0}-n_{t-1}\mathbf{z}_{T})^{2}}{\sigma_{t-1}^{2}}-\frac{(\mathbf{z}_{t}-m_{t}\mathbf{z}_{0}-n_{t}\mathbf{z}_{T})^{2}}{\sigma_{t}^{2}}\bigg)\bigg]$
 $=$ $\exp\bigg[-\frac{1}{2}\bigg(\bigg(\frac{a_{t}^{2}}{\delta_{t}^{2}}+\frac{1}{\sigma_{t-1}^{2}}\bigg)\mathbf{z}_{t-1}^{2}-2\bigg(\frac{a_{t}\mathbf{z}_{t}-a_{t}b_{t}\mathbf{z}_{T}}{\delta_{t}^{2}}+\frac{m_{t-1}\mathbf{z}_{0}+n_{t-1}\mathbf{z}_{T}}{\sigma_{t-1}^{2}}\bigg)\mathbf{z}_{t-1}+C(\mathbf{z}_{t},\mathbf{z}_{0},\mathbf{z}_{T})\bigg)\bigg].$
@@
 $=\frac{\sigma_{t-1}^{2}}{\sigma_{t}^{2}}a_{t}\mathbf{z}_{t}+\frac{\delta_{t}^{2}}{\sigma_{t}^{2}}m_{t-1}\mathbf{z}_{0}+\bigg(\frac{\delta_{t}^{2}}{\sigma_{t}^{2}}n_{t-1}-\frac{\sigma_{t-1}^{2}}{\sigma_{t}^{2}}a_{t}b_{t}\bigg)\mathbf{z}_{T}$ (18)
-whose results are equivalent to Equation 6 and Equation 7.
+whose results are equivalent to Equation˜6 and Equation˜7.
 Accordingly, maximizing the Evidence Lower Bound (ELBO) is equivalent to minimizing the KL divergence between ground truth and predicted posterior:
@@
 $=\mathbb{E}_{t,\mathbf{z}_{0},\mathbf{z}_{t},\mathbf{z}_{T}}\left[\left\|\boldsymbol{\epsilon}-\boldsymbol{\epsilon}_{\theta}(\mathbf{z}_{t},\mathbf{z}_{T},t)\right\|^{2}\right]$
-which is the same as in Equation 8. This concludes our derivation.
+which is the same as in Equation˜8. This concludes our derivation.
 Appendix B Training and Inference Pipeline
@@
 0: Encoder $\mathcal{E}$; Decoder $\mathcal{D}$; Trained $\boldsymbol{\epsilon}_{\theta}(\cdot,\cdot,\cdot)$
-1:  $\mathbf{z}_{T}=\mathcal{E}(\mathbf{x})$
+[…]
 2:for $t=T,\dots,1$ do
@@
 7:end for
-8:  $\mathbf{y}=\mathcal{D}(\hat{\mathbf{z}}_{0})$
+[…]
 9:return $\mathbf{y}$
 B.2 Accelerate Inference
-Similar to DDIM [45] and DBIM [64], the inference process of DPBridge can be accelerated using non-Markovian sampling scheme that preserves the marginal distributions of intermediate states.
+Similar to DDIM [song2020denoising] and DBIM [zheng2024diffusion], the inference process of DPBridge can be accelerated using non-Markovian sampling scheme that preserves the marginal distributions of intermediate states.
 Specifically, we perform sampling over a reduced set of timesteps $\{t_{i}\}_{i=1}^{N}$, where $0=t_{0}<t_{1}<\cdots<t_{N-1}<t_{N}=T$, allowing the number of inference steps $N$ to be decoupled from the total number of training steps $T$.
@@
 where $g_{n}$ denotes variance parameter that governs the stochasticity of the reverse sampling process.
-Notably, when $g_{n}=\tilde{\sigma}_{t}^{2}$, as defined Equation 5, the process becomes a Markovian bridge.
+Notably, when $g_{n}=\tilde{\sigma}_{t}^{2}$, as defined Equation˜5, the process becomes a Markovian bridge.
 Appendix C Network Architecture
@@
 To address the dimension incompatibility with pretrained backbone, the input channels of the network are thus doubled to accommodate the concatenated input.
-In addition, as suggested by [28], we duplicate the weight tensor of these accommodated layers and divide the values by two to preserve the input magnitude to subsequent layers.
+In addition, as suggested by [ke2024repurposing], we duplicate the weight tensor of these accommodated layers and divide the values by two to preserve the input magnitude to subsequent layers.
 As DPBridge is an image-conditioned generative framework, we employ null-text embedding for all cross-attention layers to eliminate the impact of text prompts.
@@
 During training, the number of timesteps of is set to be 1000, and we use 50 sampling steps during inference with considerations of both quality and efficiency.
-Training our method takes 30K iterations using the Adam optimizer with a base learning rate of $3\times 10^{-5}$ and an exponential decay strategy after 100-step warm-up.
+Training our method takes 30K iterations using the Adam optimizer with a base learning rate of $3\times 10^{-5}$ and an exponential decay strategy after 100-steps warm-up.
 The batch size is 2, but with 8 steps gradient accumulation, the effective batch size is equivalent to 16.
@@
 E.1 Additional Quantitative Results
-In this section, we further evaluate the robustness of DPBridge system by perturbing the input image with different types and different level intensities of noise.
+In this section, we further evaluate the robustness of DPBridge system by perturbing the input image with different types and different levels of intensity of noise.
 In this experiment setting, we use Stable Diffusion 1.5 as pretrained backbone.
-The evaluation is conducted using under both depth estimation and surface normal prediction tasks.
-We use NYUv2 and KITTI to report the depth estimation results, as for surface normal prediction, we report the results on NYUv2 and iBims-1.
-Table 4 reveals that the model’s performance degrades as noise intensity increases, but the model performance remains acceptable under moderate-level-intensity noise.
+The evaluation is conducted under both depth estimation and surface normal prediction tasks.
+We use NYUv2 and KITTI to report the depth estimation results, and for surface normal prediction, we report the results on NYUv2 and iBims-1.
+Tab.˜4 reveals that the model’s performance degrades as noise intensity increases, but the model performance remains acceptable under moderate-level-intensity noise.
 Among noise types, uniform noise shows the mildest effect on degradation, while salt & pepper and high-intensity Gaussian noise cause the most significant performance decline.
@@
 E.2 Additional Qualitative Results
-To further evaluate the effectiveness and generalization capability of DPBridge, we extend the framework to additional dense prediction tasks, including semantic segmentation, optical flow estimation, and style transfer.
-Qualitative results are shown in Figure 7, Figure 8, Figure 9, and Figure 10.
-These extensions are implemented by framing each task as a continuous-valued image-to-image translation problem, and replacing the paired ground truth data accordingly.
+To further evaluate the effectiveness and generalization capability of DPBridge, we extend the framework to additional dense prediction tasks, including semantic segmentation, optical flow estimation, edge detection, and style transfer.
+Qualitative results are shown in Figure˜7, Figure˜9, Figure˜8, and Figure˜10.
+These extensions are implemented by framing each task as a continuous-valued image-to-image translation problem and replacing the paired ground truth data accordingly.
 For semantic segmentation, we convert discrete labels into color maps and train the bridge model to predict colors in RGB space rather than class indices.
@@
 We note that these results are qualitative; quantitative evaluation and task-specific architectural refinements are left for future work.
-Figure 7: Qualitative examples of applying DPBridge to semantic segmentation task on CityScape [12] dataset.
+Figure 7: Qualitative examples of applying DPBridge to semantic segmentation task on CityScape [cordts2016cityscapes] dataset.
 We transform discrete labels into color maps and us our bridge model to predict colors in RGB space instead of categorical labels.
-Figure 8: Qualitative examples of applying DPBridge to semantic segmentation task on CelebAMask-HQ [30] dataset.
-We transform discrete labels into color maps and us our bridge model to predict colors in RGB space instead of categorical labels.
-Figure 9: Qualitative examples of applying DPBridge to optical flow prediction task on MPI-Sintel [5] dataset.
+Figure 8: Qualitative examples of applying DPBridge to optical flow prediction task on MPI-Sintel [butler2012naturalistic] dataset.
 The input image is the overlayed RGB pair, and we transform the flow map into RGB space so that the diffusion bridge process can be constructed accordingly.
-Figure 10: Qualitative examples of applying DPBridge to style transfer task on Face-to-Comics [48] dataset.
-Appendix F Broader Impacts
-DPBridge aims to advance the field of dense prediction through a generative modeling perspective, integrating the strengths of diffusion bridge formulation with visual priors from pretrained backbones.
-Our proposed method is a unified framework that can be easily extend to different tasks, including depth prediction, surface normal estimation, semantic segmentation, optical flow, etc., which will benefit the whole dense prediction research community.
-In addition, improved generalization and precision in dense prediction can also benefit a wide range of real-world applications, such as autonomous driving, augmented reality, and digital content creation.
+Figure 9: Qualitative examples of applying DPBridge to edge detection task on MultiCue [mely2016systematic] dataset.
+Figure 10: Qualitative examples of applying DPBridge to style transfer task on Face-to-Comics [face2comics] dataset.